### v1

# Generalization of Gibbs and Langevin Monte Carlo Algorithms in the Interpolation Regime

Abstract
The paper provides data-dependent bounds on the test error of the Gibbs algorithm in the overparameterized interpolation regime, where low training errors are also obtained for impossible data, such as random labels in classification. The bounds are stable under approximation with Langevin Monte Carlo algorithms. Experiments on the MNIST and CIFAR-10 datasets verify that the bounds yield nontrivial predictions on true labeled data and correctly upper bound the test error for random labels. Our method indicates that generalization in the low-temperature, interpolation regime is already signaled by small training errors in the more classical high temperature regime.
1 Introduction
Modern learning algorithms can achieve very small training errors on arbitrary data if the underlying hypothesis space is large enough. For reasonable data originating from real-world problems, the chosen hypotheses also tend to have small test errors, a fortunate circumstance, which has given great technological and economic thrust to deep learning. Unfortunately, the same algorithms also achieve very small training errors for data specifically designed to produce very large test errors, such as random labels in classification. Consequently, the hypothesis space and the training error do not suffice to predict the test error. The key to generalization must be more deeply buried in the data.
While not so disquieting to practitioners, this mystery has troubled theoreticians for many years (Zhang et al., 2016; 2021), and it seems safe to say that the underlying mechanisms still have not been completely understood.
We are far from solving this riddle in generality, but for the Gibbs posterior we show how nontrivial bounds on the test error can be recovered from the training data. The Gibbs posterior assigns probabilities, which decrease exponentially with the training error of the hypotheses. The exponential decay parameter $\beta$ can be interpreted as an inverse temperature in an analogy to statistical physics. The Gibbs measure is a sufficient idealization to have tractable theoretical properties, but also the limiting distribution of several concrete stochastic algorithms, here summarized as Langevin Monte Carlo (LMC), including Stochastic Gradient Langevin Dynamics (SGLD), (Gelfand & Mitter, 1991; Welling & Teh, 2011), a popular modern learning algorithm.
When $\beta$ is large and the hypothesis space is rich, these algorithms can reproduce the dilemma described above by achieving very small training errors on data designed to have large test errors. Our paper addresses this interpolation regime of the Gibbs posterior and makes the following three contributions:
  • We give high-probability data-dependent bounds on the test error, both for a hypothesis drawn from the Gibbs posterior and for the posterior mean, assuming that we can freely draw samples from it. These bounds holds for the entire range of temperatures.
  • We weaken the above assumption by showing that the bounds are stable under approximations of the posterior in the total variation and $W_{2}$-Wasserstein metrics. Given sufficient computing resources, this yields bounds for LMC algorithms.
  • Using LMC on real-world classification data (MNIST and CIFAR-10), we obtain non-trivial upper bounds on the test error for true labels and upper bounds for random labels, despite the algorithms achieving very small training error on the random labels.
Our method is based on a combination of the PAC-Bayesian bounds (McAllester, 1999; Alquier, 2021; Rivasplata et al., 2020) with an integral representation of the log-partition function. This makes it possible to bound the logarithm of the density of the Gibbs posterior at a given temperature in terms of empirical averages at higher temperatures. A qualitative conclusion is that generalization in the under-regularized low-temperature regime ($\beta>n$) is already indicated by small training errors in the over-regularized high-temperature regime ($\beta<n$), where $n$ is the number of training examples.
1.1 Related literature
Many papers address the generalization of the Gibbs algorithm and Langevin Monte Carlo, with special focus on SGLD, which is the most popular algorithm. Most similar to this work is Raginsky et al. (2017), which bounds the distance to the Gibbs posterior and then its generalization error. Their bound, however, applies only to the high temperature regime $\beta<n$.
Several works concentrate on the optimization path of SGLD. Mou et al. (2018) gives both stability and PAC-Bayesian bounds. Pensia et al. (2018) applies the information theoretic generalization bounds of Xu & Raginsky (2017). These ideas are further developed by Negrea et al. (2019), where random subsets of the training data are used to define data-dependent priors. Farghly & Rebeschini (2021) gives time-independent bounds for SGLD, which are further improved by Futami & Fujisawa (2024). Most of the bounds in the above papers are in expectation. The very recent paper of Harel et al. (2025) gives a very elegant argument for Markov chain algorithms based on the second law of thermodynamics. If the invariant distribution is the Gibbs posterior, the bound along the entire optimization path is of order $\sqrt{\beta/n}$ but improvable to $\beta/n$.
Some papers give similar bounds for the Gibbs posterior, roughly of the form $\beta/n$ or $\sqrt{\beta/n}$ (Raginsky et al., 2017; Dziugaite & Roy, 2018; Kuzborskij et al., 2019; Rivasplata et al., 2020) or Maurer (2024) and Harel et al. (2025)). These bounds hold equally for random labels and are therefore vacuous for overparametrized hypothesis spaces in the low temperature regime $\beta>n$. To our knowledge, ours is the only bound for the Gibbs posterior, which is valid in this regime.
Other bounds have been developed for specific algorithms designed to optimize them. The milestone paper by Dziugaite & Roy (2017) is the most prominent example, and (Dziugaite & Roy, 2018) and (Pérez-Ortiz et al., 2021) are also in this category. Our bounds by contrast apply to the Gibbs posterior and LMC in their standard forms.
2 Preliminaries
The relative entropy of two Bernoulli variables with expectations $p$ and $q$ is denoted
$\kappa\left(p,q\right)=p\ln\frac{p}{q}+\left(1-p\right)\ln\frac{1-p}{1-q}.$ (1)
We also define the function $\kappa^{-1}:\left[0,1\right]\times\left[0,\infty\right)\rightarrow\left[0,1\right]$ by
$\kappa^{-1}\left(p,t\right)=\inf\left\{q:q\geq p,\kappa\left(p,q\right)\geq t\right\}.$
Throughout the following $\left(\mathcal{X},\Sigma\right)$ is a measurable space of data with probability measure $\mu$. The iid random vector $\mathbf{x}\sim\mu^{n}$ is the training sample.
We let $\left(\mathcal{H},\Omega\right)$ be a measurable space of hypotheses, and let $\ell:\mathcal{H\times X}\rightarrow\left[0,\infty\right)$ be a prescribed loss function. Members of $\mathcal{H}$ are denoted $h$ or $g$. We write $L\left(h\right):=\mathbb{E}_{x\sim\mu}\left[\ell\left(h,x\right)\right]$ and $\hat{L}\left(h,\mathbf{x}\right):=\left(1/n\right)\sum_{i}\ell\left(h,x_{i}\right)$ respectively for the true (expected) and empirical error of hypothesis $h\in\mathcal{H}$. The set of probability measures on $\left(\mathcal{H},\Omega\right)$ is denoted $\mathcal{P}\left(\mathcal{H}\right)$.
A stochastic algorithm is a function $\nu:\mathcal{X}^{n}\rightarrow\mathcal{P}\left(\mathcal{H}\right)$, which assigns to a training sample $\mathbf{x}$ a probability measure $\nu\left(\mathbf{x}\right)\in\mathcal{P}\left(\mathcal{H}\right)$. The $KL$-divergence between two probability measures is the function $KL:(\rho,\nu)\in\mathcal{P}(\mathcal{H})\times\mathcal{P}(\mathcal{H})\mapsto\mathbb{E}_{h\sim\rho}\big[\ln\frac{d\rho}{d\nu}\big]$ if $\rho$ absolutely continuous w.r.t. $\nu$, otherwise the value is $\infty$. The total variation distance is defined as $d_{TV}:(\rho,\nu)\in\mathcal{P}(\mathcal{H})\times\mathcal{P}(\mathcal{H})\mapsto\sup_{A\in\Omega}|\rho(A)-\nu(A)|$. The $W_{p}$-Wasserstein distance is $W_{p}(\rho,\nu)=(\inf_{W}\mathbb{E}_{(x,y)\sim W}[\|x-y\|^{p}])^{1/p}$ with the infimum over all probability measures on $\mathcal{P}(\mathcal{H})\times\mathcal{P}(\mathcal{H})$ with $\rho$ and $\nu$ as marginals.
There is an a-priori reference measure $\pi\in\mathcal{P}\left(\mathcal{H}\right)$, called the prior. With a fixed prior, the Gibbs algorithm at inverse temperature $\beta>0$ is the stochastic algorithm $G_{\beta}:\mathbf{x}\in\mathcal{X}^{n}\mapsto G_{\beta}\left(\mathbf{x}\right)\in\mathcal{P}\left(\mathcal{H}\right)$ defined by
$G_{\beta}\left(\mathbf{x}\right)\left(A\right)=\frac{1}{Z_{\beta}\left(\mathbf{x}\right)}\int_{A}e^{-\beta\hat{L}\left(h,\mathbf{x}\right)}d\pi\left(h\right)\text{ for }A\in\Omega\text{.}$
$G_{\beta}\left(\mathbf{x}\right)$ is called the Gibbs-posterior, the normalizing factor
$Z_{\beta}\left(\mathbf{x}\right):=\int_{\mathcal{H}}e^{-\beta\hat{L}\left(h,\mathbf{x}\right)}d\pi\left(h\right)$
is called the partition function. The motivation for the Gibbs posterior is that it puts larger weights on hypotheses with smaller empirical error.
Given a stochastic algorithm $\nu$ we define a probability measure $\rho_{\nu}$ on $\mathcal{H}\times\mathcal{X}^{n}$ by
$\rho_{\nu}\left(A\right)=\mathbb{E}_{\mathbf{x}\sim\mu^{n}}\mathbb{E}_{h\sim\nu\left(\mathbf{x}\right)}\left[1_{A}\left(h,\mathbf{x}\right)\right]\text{ for }A\in\Omega\otimes\Sigma^{\otimes n}.$ (2)
Then, $\mathbb{E}_{\left(h,\mathbf{x}\right)\sim\rho_{\nu}}\left[\phi\left(h,\mathbf{x}\right)\right]=\mathbb{E}_{\mathbf{x}}\mathbb{E}_{h\sim\nu\left(\mathbf{x}\right)}\left[\phi\left(h,\mathbf{x}\right)\right]$ for measurable $\phi:\mathcal{H}\times\mathcal{X}^{n}\rightarrow\mathbb{R}$. To draw the pair $\left(h,\mathbf{x}\right)$ from $\rho_{\nu}$ we first draw the training sample $\mathbf{x}$, and then sample $h$ from $\nu\left(\mathbf{x}\right)$. The main objective in learning is that the risk $\mathbb{E}_{x\sim\mu}\left[f\left(h,x\right)\right]$ is small with high probability in $\left(h,\mathbf{x}\right)\sim\rho_{\nu}$, where $f$ is some application-dependent loss function, possibly different from $\ell$. In the sequel we will give corresponding guarantees.
3 Bounds for the Gibbs posterior
In this section, we make the idealized assumption that we are free to sample from the Gibbs posterior at any finite $\beta\geq 0$.
3.1 An integral representation of the free energy
Lemma 3.1.
Let $0=\beta_{0}<\beta_{1}<\cdots<\beta_{K}=\beta$. Then
$\ln Z_{\beta}(\mathbf{x})=\int_{0}^{\beta}\mathbb{E}_{h\sim G_{\gamma}(\mathbf{x})}\big[\hat{L}(h,\mathbf{x})\big]d\gamma\leq\sum_{k=1}^{K}(\beta_{k}-\beta_{k-1})\mathbb{E}_{g\sim G_{\beta_{k-1}}(\mathbf{x})}\big[\hat{L}(g,\mathbf{x})\big].$
Proof.
Let $A(\beta)=-\ln Z_{\beta}(\mathbf{x})$. One verifies the identities
$A(0)$ $=$ $0,$
$A^{\prime}(\beta)$ $=$ $\frac{1}{Z_{\beta,\pi}(\mathbf{x})}\int_{\mathcal{H}}\hat{L}(h,\mathbf{x})e^{-\beta\hat{L}(h,\mathbf{x})}d\pi(h)=\mathbb{E}_{h\sim G_{\beta}(\mathbf{x}}\big[\hat{L}(h,\mathbf{x})\big],$
$A^{\prime\prime}(\beta)$ $=$ $-\Big(\mathbb{E}_{h\sim G_{\beta}(\mathbf{x})}\big[\hat{L}(h,\mathbf{x})^{2}\big]-\big(\mathbb{E}_{h\sim G_{\beta}(\mathbf{x})}\big[\hat{L}(h,\mathbf{x})\big]\big)^{2}\Big)\leq 0\text{.}$
The equality in the lemma then follows from the first two identities above and the fundamental theorem of calculus, and the inequality follows from the last identity, which shows that $\mathbb{E}_{g\sim G_{\beta_{k-1},\pi}(\mathbf{x})}\big[\hat{L}(g,\mathbf{x})\big]$ is non-increasing in $\beta$. ∎
In statistical physics there is a formal analogy, where the function $h\mapsto\hat{L}(h,\mathbf{x})$ is the ($\mathbf{x}$-dependent) energy of the system in the state $h$, and $\beta$ is the inverse temperature. The Gibbs posterior then becomes the "canonical ensemble" (Gibbs, 1902), describing the probability of states in equilibrium with a heat bath at temperature $\beta^{-1}$. The function $\beta\mapsto A(\beta)$ plays an important role: $\beta^{-1}A(\beta)$ is the Helmholz free energy, $A^{\prime}(\beta)=\mathbb{E}_{h\sim G_{\beta}(\mathbf{x})}\big[\hat{L}(h,\mathbf{x})\big]$ is the thermal average of the energy, $\beta A^{\prime}(\beta)+A(\beta)$ is the entropy and $A^{\prime\prime}(\beta)$ is proportional to the heat capacity at temperature $\beta^{-1}$ (see e.g. Huang, 2008).
For $h\in\mathcal{H}$, $\mathbf{x\in}\mathcal{X}^{n}$ and an increasing sequence $\bm{\beta}=\left(\beta_{1}<\cdots<\beta_{K}\right)$ of positive numbers, we denote
$\Gamma(h,\mathbf{x},\bm{\beta})=-\beta_{K}\hat{L}(h,\mathbf{x})+\sum_{k=1}^{K}(\beta_{k}-\beta_{k-1})\mathbb{E}_{g\sim G_{\beta_{k-1}}(\mathbf{x})}\big[\hat{L}(g,\mathbf{x})\big].$ (3)
So Lemma 3.1 states that, when $\beta_{0}=0$, then,
$\beta_{K}\hat{L}\left(h,\mathbf{x}\right)-\ln Z_{\beta_{K}}\left(\mathbf{x}\right)\leq\Gamma\left(h,\mathbf{x},\bm{\beta}\right).$ (4)
Note that $\Gamma\left(h,\mathbf{x},\bm{\beta}\right)$ depends only on the training data $\mathbf{x}$, the sequence $\bm{\beta}$ and the hypothesis $h$.
3.2 Bounds
The function $F$ in the following is a placeholder for a random variable related to the generalization gap, which we would like to bound with high probability.
Theorem 3.2.
Let $F:\mathcal{H}\times\mathcal{X}^{n}\rightarrow\mathbb{R}$ be some measurable function, $\beta>0$ and $\bm{\beta}=\left(\beta_{1}<\cdots<\beta_{K}\right)$ as above with $\beta_{0}=0$ and $\beta_{K}=\beta$. Then,
(i) for $\delta>0$ with probability at least $1-\delta$ in $\mathbf{x}\sim\mu^{n}$ and $h\sim G_{\beta}\left(\mathbf{x}\right)$
$F\left(h,\mathbf{x}\right)\leq\Gamma\left(h,\mathbf{x},\bm{\beta}\right)+\ln\mathbb{E}_{\mathbf{x}}\mathbb{E}_{g\sim\pi}\left[e^{F\left(g,\mathbf{x}\right)}\right]+\ln\left(1/\delta\right),$
(ii) for $\delta>0$ with probability at least $1-\delta$ in $\mathbf{x}\sim\mu^{n}$
$\mathbb{E}_{h\sim G_{\beta}\left(\mathbf{x}\right)}\left[F\left(h,\mathbf{x}\right)\right]\leq\mathbb{E}_{h\sim G_{\beta}\left(\mathbf{x}\right)}\left[\Gamma\left(h,\mathbf{x},\bm{\beta}\right)\right]+\ln\mathbb{E}_{\mathbf{x}}\mathbb{E}_{g\sim\pi}\left[e^{F\left(g,\mathbf{x}\right)}\right]+\ln\left(1/\delta\right).$
Proof.
By Markov’s inequality, for any real random variable $Y$
$\Pr\left\{Y>\ln\mathbb{E}\left[e^{Y}\right]+\ln\left(1/\delta\right)\right\}=\Pr\left\{e^{Y}>\mathbb{E}\left[e^{Y}\right]/\delta\right\}\leq\delta.$
To prove (i), we apply this to the random variable $Y=F\left(h,\mathbf{x}\right)+\beta\hat{L}\left(h,\mathbf{x}\right)+\ln Z_{\beta}\left(\mathbf{x}\right)$ on the probability space $\left(\mathcal{H\times X}^{n},\Omega\otimes\Sigma^{\otimes n},\rho_{G_{\beta}}\right)$ as defined in (2). Together with the definition of the Gibbs posterior, this gives, with probability at least $1-\delta$ in $\left(h,\mathbf{x}\right)\sim\rho_{G_{\beta}}$ (equivalent to saying $\mathbf{x}\sim\mu^{n}$ and $h\sim G_{\beta}\left(\mathbf{x}\right)$),
$F\left(h,\mathbf{x}\right)+\beta\hat{L}\left(h,\mathbf{x}\right)+\ln Z_{\beta}\left(\mathbf{x}\right)$
$\leq\ln\mathbb{E}_{\mathbf{x}}\mathbb{E}_{g\sim G_{\beta}\left(\mathbf{x}\right)}\left[e^{F\left(g,\mathbf{x}\right)+\beta\hat{L}\left(g,\mathbf{x}\right)+\ln Z_{\beta}\left(\mathbf{x}\right)}\right]+\ln\left(1/\delta\right)$
$=\ln\mathbb{E}_{\mathbf{x}}\mathbb{E}_{g\sim\pi}\left[e^{F\left(g,\mathbf{x}\right)+\beta\hat{L}\left(g,\mathbf{x}\right)+\ln Z_{\beta}\left(\mathbf{x}\right)-\beta\hat{L}\left(g,\mathbf{x}\right)-\ln Z_{\beta}\left(\mathbf{x}\right)}\right]+\ln\left(1/\delta\right)$
$=\ln\mathbb{E}_{\mathbf{x}}\mathbb{E}_{g\sim\pi}\left[e^{F\left(g,\mathbf{x}\right)}\right]+\ln\left(1/\delta\right).$
Subtract $\beta\hat{L}\left(h,\mathbf{x}\right)+\ln Z_{\beta}\left(\mathbf{x}\right)$ and use (4). For (ii) apply Markov’s inequality to $\mathbb{E}_{h\sim G_{\beta}\left(\mathbf{x}\right)}\left[F\left(h,\mathbf{x}\right)+\beta\hat{L}\left(h,\mathbf{x}\right)+\ln Z_{\beta}\left(\mathbf{x}\right)\right]$ instead. By Jensen’s inequality
$e^{\mathbb{E}_{h\sim G_{\beta}\left(\mathbf{x}\right)}\left[F\left(h,\mathbf{x}\right)+\beta\hat{L}\left(h,\mathbf{x}\right)+\ln Z_{\beta}\left(\mathbf{x}\right)\right]}\leq\mathbb{E}_{g\sim G_{\beta}\left(\mathbf{x}\right)}\left[e^{F\left(g,\mathbf{x}\right)+\beta\hat{L}\left(g,\mathbf{x}\right)+\ln Z_{\beta}\left(\mathbf{x}\right)}\right]$
and proceed as above using (4). ∎
Up to the application of (4), the above proof of (i) just gives the single-draw version of the PAC-Bayesian bound as in Rivasplata et al. (2020) applied to the Gibbs posterior, while (ii) is the standard PAC-Bayesian bound applied to the Gibbs posterior until (4) is invoked.
3.3 Loss functions and secondary loss functions
To apply Theorem 3.2, we need to control the exponential moment $\mathbb{E}_{\mathbf{x}}\mathbb{E}_{g\sim\pi}\left[e^{F\left(g,\mathbf{x}\right)}\right]$, but otherwise we have free choice of the function $F$. This gives the method some flexibility. If $\exp\left(F\right)$ has sufficient integrability properties, we can exchange the two expectations and often there is a bound on $\mathbb{E}_{\mathbf{x}}\left[e^{F\left(g,\mathbf{x}\right)}\right]$ uniform in $g$, which then carries over to $\mathbb{E}_{g\sim\pi}\mathbb{E}_{\mathbf{x}}\left[e^{F\left(g,\mathbf{x}\right)}\right]$, since $\pi$ is a probability measure. In this way bounds for sub-Gaussian or sub-exponential losses can be obtained, but also for U-statistics or even non-iid data, sampled from the trajectories of time-homogeneous, ergodic Markov chains. In Section B.1 in the appendix, we derive a bound for sub-Gaussian losses from Theorem 3.2; other examples are planned for a longer version of the paper.
The function $F$ may be defined in terms of other, application-dependent loss functions, which are different from the loss $\ell$, which defines the Gibbs posterior and the functional $\Gamma$. To illustrate this point, let $f:\mathcal{H}\times\mathcal{X}\rightarrow[0,1]$ be measurable and set $F(h,\mathbf{x})=n~\kappa\left(\frac{1}{n}\sum_{i}f(h,x_{i}),\mathbb{E}_{x}\left[f(h,x)\right]\right)$, with $\kappa$ the relative entropy as in (1). Then, Theorem 1 of Maurer (2004) gives $\mathbb{E}_{\mathbf{x}}\left[e^{F(h,\mathbf{x})}\right]\leq 2\sqrt{n}$ for $n\geq 8$. Substitution in Theorem 3.2 and division by $n$ then give the following corollary.
Corollary 3.3.
Let $f:\mathcal{H}\times\mathcal{X}\rightarrow[0,1]$ be measurable, $\delta>0$ and $n\geq 8$. Then, with probability at least $1-\delta$ in $\mathbf{x}\sim\mu^{n}$ and $h\sim G_{\beta}(\mathbf{x})$
$\kappa\left(\frac{1}{n}\sum_{i}f(h,x_{i}),\mathbb{E}_{x}[f(h,x)]\right)\leq\frac{1}{n}\left(\Gamma(h,\mathbf{x},\bm{\beta})+\ln\left(\frac{2\sqrt{n}}{\delta}\right)\right)$
and with probability at least $1-\delta$ in $\mathbf{x}\sim\mu^{n}$
$\kappa\left(\frac{1}{n}\sum_{i}\mathbb{E}_{h\sim G_{\beta}\left(\mathbf{x}\right)}\left[f\left(h,x_{i}\right)\right],\mathbb{E}_{h\sim G_{\beta}\left(\mathbf{x}\right)}\mathbb{E}_{x}\left[f\left(h,x\right)\right]\right)\leq\frac{1}{n}\left(\Gamma\left(h,\mathbf{x},\bm{\beta}\right)+\ln\left(\frac{2\sqrt{n}}{\delta}\right)\right).$
For the second part, we used the joint convexity of $\kappa$. Under the conditions of this corollary, the second inequality becomes
$\mathbb{E}_{h\sim G_{\beta}\left(\mathbf{x}\right)}\mathbb{E}_{x}\left[f\left(h,x\right)\right]\leq\kappa^{-1}\left(\frac{1}{n}\sum_{i}\mathbb{E}_{h\sim G_{\beta}\left(\mathbf{x}\right)}\left[f\left(h,x_{i}\right)\right],\frac{1}{n}\left(\Gamma\left(h,\mathbf{x},\mathbf{\beta}\right)+\ln\left(\frac{2\sqrt{n}}{\delta}\right)\right)\right)$
with an analogous version for the single-draw case. This is how we compute bounds in our experiments. For an illustration, please refer to C.2.1.
Here $f$ plays the role of a secondary loss function, typically different from the loss function $\ell$, which defines the Gibbs posterior. In applications one would define the Gibbs posterior in terms of a differentiable loss function $\ell$ and approximate it with a suitable Monte Carlo method. For classification, however, one is interested in bounding the 0-1 loss obtained by comparing $\ell$ to some threshold $\theta$. Then (momentarily changing notation by replacing $x\in\mathcal{X}$ by $\left(x,y\right)$, where $y$ is the label corresponding to $x$) one can define $f\left(h,\left(x,y\right)\right)=1_{\left(-\infty,\theta\right)}\left(y\ell\left(h,x\right)\right)$ to obtain a bound on the expected 0-1 loss.
4 Bounds for Langevin Monte Carlo
For this section, we assume $\mathcal{H}=\mathbb{R}^{d}$ and an isotropic Gaussian prior $\pi$ of width $\sigma$. We condition on the training data $\mathbf{x}$, reference to which we omit. The Gibbs posterior is an idealization, from which it is impossible to sample directly. Nevertheless a number of works (Raginsky et al., 2017; Dalalyan & Karagulyan, 2017; Brosse et al., 2018; Vempala & Wibisono, 2019; Dwivedi et al., 2019; Nemeth & Fearnhead, 2021; Balasubramanian et al., 2022) discuss algorithms (SGLD, ULA, MALA, etc, here summarized as Langevin Monte Carlo (LMC)), capable of approximating a probability measure $\nu$ on $\mathbb{R}^{d}$ of the form $\nu\propto\exp\left(-V\right)$ or some nearby limiting distribution. In the following, we discuss one of these algorithms.
4.1 ULA
We focus on the results of Vempala & Wibisono (2019), which do not require convexity of $V$ and instead assume that the measure $\nu$ satisfies a log-Sobolev inequality (LSI) in the sense that for all smooth $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$
$\mathbb{E}_{h\sim\nu}\left[f^{2}\left(h\right)\ln f^{2}\left(h\right)\right]-\mathbb{E}_{h\sim\nu}\left[f^{2}\left(h\right)\right]\ln\mathbb{E}_{h\sim\nu}\left[f^{2}\left(h\right)\right]\leq\frac{2}{\alpha}\mathbb{E}_{h\sim\nu}\left[\left\|\left(\nabla f\right)\left(h\right)\right\|^{2}\right]$ (5)
for some $\alpha>0$. An LSI is satisfied when $V$ is strongly convex, but, importantly, also for measures which are bounded perturbations of measures satisfying an LSI (Holley & Stroock (1986)). Vempala & Wibisono (2019) give further examples and a list of references for measures, which are not log-concave and satisfy an LSI. Raginsky et al. (2017) show, that under dissipativity conditions of the loss the Gibbs posterior $G_{\beta}\left(\mathbf{x}\right)$ satisfies an LSI with constant independent of $\mathbf{x}$.
Consider the iterative algorithm
$h_{t+1}=h_{t}-\epsilon\nabla V\left(h_{t}\right)+\sqrt{2\epsilon}\xi_{t},$ (6)
where $\epsilon$ is a step size, the $\xi_{t}\sim\mathcal{N}\left(0,I\right)$ are independent Gaussian vectors and $h_{0}$ is drawn from some initial distribution $\nu_{0}$. Some authors call this algorithm simply LMC, for Langevin Monte Carlo. We call it ULA, alongside Durmus & Moulines (2017),Dwivedi et al. (2019) and Vempala & Wibisono (2019), for Un-adjusted Langevin Algorithm, because it misses the Metropolis-type accept-reject step, which would guarantee that the invariant distribution is indeed the Gibbs posterior. A popular variant of ULA is Stochastic Gradient Langevin Dynamics (SGLD) (Welling & Teh, 2011; Raginsky et al., 2017) where the gradient is replaced by an unbiased estimate, typically realized with random minibatches. Here, we restrict ourselves to ULA with a constant step size, because it has the least number of parameters to adjust, but in experiments we also use the more computationally efficient SGLD.
As $\epsilon\,{\rightarrow}\,0$, ULA recovers the Continuous Langevin Dynamics (CLD) given by the stochastic differential equation
$dh_{t}=-V\left(h_{t}\right)dt+\sqrt{2}dB_{t},$
where $B_{t}$ is centered standard Brownian motion in $\mathbb{R}^{d}$. CLD converges exponentially to the Gibbs posterior (Chiang et al., 1987). For $\epsilon>0$, the distribution $\nu_{\epsilon,t}$ of ULA converges as $t\rightarrow\infty$ to a biased limiting distribution $\nu_{\epsilon}$ which is generally different from $\nu$, but expected to be closer to $\nu$ as $\epsilon$ becomes smaller. Vempala & Wibisono (2019) use the LSI assumption and coupling to control the difference between CLD and ULA along their path and prove the following result.
Theorem 4.1.
Assume that $\nu$ satisfies the log-Sobolev inequality (5) with $\alpha>0$, that the Hessian of $V$ satisfies $LI\preceq\nabla^{2}V\left(h\right)\preceq LI$ for all $h$ and some $L<\infty$, and that $0<\epsilon\leq\alpha/\left(4L^{2}\right)$. Then, for $t\geq 0$
$KL\left(\nu,\nu_{\epsilon,t}\right)\leq e^{-\alpha\epsilon t}KL\left(\nu,\nu_{0}\right)+\frac{8\epsilon dL^{2}}{\alpha}.$
The first exponential term is due to the mismatch of the initial distribution and $\nu$. The second term bounds the divergence between the limiting distribution $\nu_{\epsilon}\,\$and $\nu$. Similar results exist under different conditions on the potential $V$. Cheng et al. (2018) for example require $V$ to be strongly convex outside of a ball instead of the log-Sobolev inequality and gives bounds in terms of the $W_{1}$-Wasserstein metric. Raginsky et al. (2017) give bounds for $W_{2}$ under dissipativity assumptions. The next corollary adapts Theorem 4.1 to the situation studied in this paper.
Corollary 4.2.
For $\beta>0$ consider the Gibbs posterior $G_{\beta}$ corresponding to $\hat{L}\left(h\right)$, with centered Gaussian prior of width $\sigma$. Assume that it satisfies the log-Sobolev inequality (5) with $\alpha>0$, that the Hessian of $\hat{L}$ satisfies $RI\preceq\nabla^{2}\hat{L}\left(h\right)\preceq RI$ for all $h$ and some $R<\infty$, and that $0<\eta\leq\alpha/\left(4\left(\beta R+\frac{1}{\sigma^{2}}\right)^{2}\right)$. Consider the algorithm
$h_{t+1}=h_{t}-\eta\nabla_{h}\hat{L}\left(h_{t}\right)-\frac{\eta h_{t}}{\beta\sigma^{2}}+\sqrt{\frac{2\eta}{\beta}}\xi_{t},$ (7)
where $h_{0}\sim\nu_{0}$ and the $\xi_{t}\sim\mathcal{N}\left(0,I\right)$ are independent Gaussian random variables. Let $D\left(\beta\right)=KL\left(G_{\beta},\nu_{0}\right)$ and let $\nu_{\beta,\eta,t}$ be the distribution of $h_{t}$ after $t$ steps. Then,
(i) $KL\left(G_{\beta},\nu_{\beta,\eta,t}\right)\leq e^{-\alpha\eta t/\beta}D\left(\beta\right)+\frac{8\eta d}{\beta\alpha}\left(\beta R+\frac{1}{\sigma^{2}}\right)^{2}.$
(ii) $W_{2}\left(G_{\beta},\nu_{\beta,\eta,t}\right)\leq\frac{2}{\alpha}e^{-\alpha\eta t/\beta}D\left(\beta\right)+\frac{16\eta d}{\beta\alpha^{2}}\left(\beta R+\frac{1}{\sigma^{2}}\right)^{2}.$
(iii) $d_{TV}\left(G_{\beta},\nu_{\beta,\eta,t}\right)\leq e^{-\alpha\eta t/\left(2\beta\right)}\sqrt{D\left(\beta\right)}+2\sqrt{\frac{\eta d}{\beta\alpha}}\left(\beta R+\frac{1}{\sigma^{2}}\right).$
Proof.
We make the identifications $V\left(h\right)=\beta\hat{L}\left(h\right)+\left\|h\right\|^{2}/\left(2\sigma^{2}\right)$, $\epsilon=\eta/\beta$ and $L=\beta R+\frac{1}{\sigma^{2}}$. Then, $\nu=G_{\beta}$ with Gaussian prior of width $\sigma$, ULA becomes (7) and (i) follows directly from Theorem 4.1, (ii) follows from Otto & Villani (2000) and the LSI assumption, and (iii) follows from Pinsker’s inequality (see e.g. Boucheron et al. (2013)). ∎
4.2 Stability of the bounds
We now show the stability of our bounds for approximation in total variation and $W_{2}$-Wasserstein metrics, under boundedness or Lipschitz conditions. Together with Corollary 4.2 this implies bounds for the algorithm defined in (7).
We assume that there is a target approximation $\nu_{\beta}\left(\mathbf{x}\right)$ of $G_{\beta}\left(\mathbf{x}\right)$, for which we want to compute a high probability bound, either for the single draw version on $F\left(h,\mathbf{x}\right)$ as $\mathbf{x}\sim\mu^{n}$ and $h\sim\nu_{\beta}\left(\mathbf{x}\right)$, or, for the classical PAC-Bayesian version, on $\mathbb{E}_{h\sim\nu_{\beta}\left(\mathbf{x}\right)}\left[F\left(h,\mathbf{x}\right)\right]$ as $\mathbf{x}\sim\mu^{n}$. It is not surprising that the single-draw bound will require a much closer approximation of the Gibbs posterior.
Since the bounding functional $\Gamma\left(h,\mathbf{x},\bm{\beta}\right)$ depends on the Gibbs posteriors $G_{\beta_{k}}\left(\mathbf{x}\right)$ for $k\in\{1,\dots,K-1\}$, we require corresponding approximations $\nu_{\beta_{k}}$ of $G_{\beta_{k}}\left(\mathbf{x}\right)$ to compute the bound. To streamline notation, we define
$\Gamma_{\nu}\left(h,\mathbf{x},\bm{\beta}\right)=-\beta\hat{L}\left(h,\mathbf{x}\right)+\sum_{k=1}^{K}\left(\beta_{k}-\beta_{k-1}\right)\mathbb{E}_{g\sim\nu_{\beta_{k-1}}\left(\mathbf{x}\right)}\big[\hat{L}(g,\mathbf{x})\big]$
for $0=\beta_{0}<\beta_{1}<\cdots<\beta_{K}=\beta$ and $\left(\nu_{0}\left(\mathbf{x}\right),\nu_{\beta_{1}}\left(\mathbf{x}\right),\cdots,\nu_{\beta_{k-1}}\left(\mathbf{x}\right)\right)\in\mathcal{P}\left(\mathcal{H}\right)^{K}$. Like $\Gamma\left(h,\mathbf{x},\mathbf{\beta}\right)$, the functional $\Gamma_{\nu}\left(h,\mathbf{x},\mathbf{\beta}\right)$ depends on the training data, but it can also be computed by repeated execution of the algorithm (7). The next theorem states the obtained bound in terms of the approximation errors in total variation.
Theorem 4.3.
Suppose that $\mathcal{H}=\mathbb{R}^{d}$ and that there are numbers $m,M<\infty$ such that for every $\mathbf{x}$ in $\mathcal{X}^{n}$ and $h\in\mathcal{H}$ we have $\left|\ell\left(h,\mathbf{x}\right)\right|\leq m$ and $\left|F\left(h,\mathbf{x}\right)\right|\leq M$. Let $0=\beta_{0}<\beta_{1}<\dots<\beta_{K}=\beta$ and $\nu_{\beta_{k}}\left(\mathbf{x}\right)\in\mathcal{P}\left(\mathcal{H}\right)$ be such that $d_{TV}\left(\nu_{\beta_{k}}\left(\mathbf{x}\right),G_{\beta_{k}}\left(\mathbf{x}\right)\right)=\epsilon_{\beta_{k}}$. Then,
(i) with probability at least $1-\delta$ as $\mathbf{x}\sim\mu^{n}$ and $h\sim\nu_{\beta}\left(\mathbf{x}\right)$
$F\left(h,\mathbf{x}\right)\leq\Gamma_{\nu}\left(h,\mathbf{x},\mathbf{\beta}\right)+\ln\mathbb{E}_{\mathbf{x}}\mathbb{E}_{h\sim\pi}\left[e^{F\left(h,\mathbf{x}\right)}\right]+\ln\frac{1}{\delta}+\ln\left(2e^{M+\beta m}\epsilon_{\beta}\right)+\sum_{k=1}^{K}\left(\beta_{k}-\beta_{k-1}\right)m\epsilon_{\beta_{k-1}}.$
(ii) with probability at least $1-\delta$ as $\mathbf{x}\sim\mu^{n}$
$\mathbb{E}_{h\sim\nu_{\beta}\left(\mathbf{x}\right)}\left[F\left(h,\mathbf{x}\right)\right]$ $\leq$ $\mathbb{E}_{h\sim\nu_{\beta}\left(\mathbf{x}\right)}\left[\Gamma_{\nu}\left(h,\mathbf{x},\bm{\beta}\right)\right]+\ln\mathbb{E}_{\mathbf{x}}\mathbb{E}_{h\sim\pi}\left[e^{F\left(h,\mathbf{x}\right)}\right]+\ln\frac{1}{\delta}$
$+\left(M+\beta m\right)\epsilon_{\beta}+\sum_{k=1}^{K}\left(\beta_{k}-\beta_{k-1}\right)m\epsilon_{\beta_{k-1}}.$
The proof, given in Section B.2, is similar to that of Theorem 3.2 and applies Markov’s inequality with $\nu_{\beta}$ instead of $G_{\beta}$. It then uses the fact that, if $f$ is a bounded measurable function, then $\left|\mathbb{E}_{\nu_{1}}\left[f\right]-\mathbb{E}_{\nu_{2}}\left[f\right]\right|\leq\left\|f\right\|_{\infty}d_{TV}\left(\nu_{1},\nu_{2}\right)$. For the single-draw version (i) this is applied to $\left(\mathbb{E}_{\nu_{\beta}}-\mathbb{E}_{G_{\beta}}\right)\left[e^{F\left(h,\mathbf{x}\right)+\beta\hat{L}\left(h,\mathbf{x}\right)+\ln Z_{\beta}\left(\mathbf{x}\right)}\right]$, which causes the exponential dependence on $\beta m$ and $M$. For (ii) we can apply this in the exponent to $\left(\mathbb{E}_{\nu_{\beta}}-\mathbb{E}_{G_{\beta}}\right)\big[F(h,\mathbf{x})+\beta\hat{L}(h,\mathbf{x})\big],$ and the logarithm makes the dependence linear. The rest of the proof is mechanical.
The last terms in (i) and (ii) are the additional errors due to the approximations of the Gibbs posteriors. The worst term is clearly the first one in (i) due to the exponential dependence on the proxy function $F$, which is typically of order $n$ and on $\beta$, which is larger than $n$ in the regime in which we are interested. In practice, the requirement of such an approximation is prohibitive for the single-draw version. The bound in (ii) has more moderate approximation requirements. In this case, we can also give bounds in terms of the $W_{2}$-Wasserstein metric (as guaranteed by Corollary 4.2), if $F$ sand $\ell$ satisfy a Lipschitz condition instead of boundedness. We will use the following fact: Since $W_{1}\leq W_{2}$ it follows from the Kantorovich-Rubinstein Theorem (Villani, 2009), that for any real Lipschitz function $f$ on $\mathcal{H}$ and probability measures $\nu_{1},\nu_{2}\in\mathcal{P}\left(\mathcal{H}\right)$
$\left|\mathbb{E}_{h\sim\nu_{1}}\left[f\left(h\right)\right]-\mathbb{E}_{h\sim\nu_{2}}\left[f\left(h\right)\right]\right|\leq\left\|f\right\|_{\rm Lip}W_{1}\left(\nu_{1},\nu_{2}\right)\leq\left\|f\right\|_{\rm Lip}W_{2}\left(\nu_{1},\nu_{2}\right),$
where $\left\|.\right\|_{\rm Lip}$ is the Lipschitz-seminorm. The following result is then immediate, with proof exactly as in (ii) of Theorem 4.3.
Theorem 4.4.
Assume the conditions of Theorem 4.3, except that instead of $\left|\ell\left(h,\mathbf{x}\right)\right|\leq m$ and $\left|F\left(h,\mathbf{x}\right)\right|\leq M$ we have $\left\|\ell\left(.,\mathbf{x}\right)\right\|_{\rm Lip}\leq m$ and $\left\|F\left(.,\mathbf{x}\right)\right\|_{\rm Lip}\leq M$ and that $W_{2}\left(\nu_{\beta_{k}}\left(\mathbf{x}\right),G_{\beta}\left(\mathbf{x}\right)\right)=\epsilon$. Then, with probability at least $1-\delta$ as $\mathbf{x}\sim\mu^{n}$
$\mathbb{E}_{h\sim\nu_{\beta}\left(\mathbf{x}\right)}\left[F\left(h,\mathbf{x}\right)\right]$ $\leq$ $\mathbb{E}_{h\sim\nu_{\beta}\left(\mathbf{x}\right)}\left[\Gamma_{\nu}\left(h,\mathbf{x},\bm{\beta}\right)\right]+\ln\mathbb{E}_{\mathbf{x}}\mathbb{E}_{h\sim\pi}\left[e^{F\left(h,\mathbf{x}\right)}\right]+\ln\frac{1}{\delta}$
$+\left(M+\beta m\right)\epsilon_{\beta}+\sum_{k=1}^{K}\left(\beta_{k}-\beta_{k-1}\right)m\epsilon_{\beta_{k-1}}.$
5 Experiments
The purpose of our experiments is to show that our method gives nontrivial bounds on the test error for real-world data, while correctly bounding the test error on impossible data, where the same algorithm also achieves a small training error. For this demonstration, we use binary classification tasks. The real-world data are either the MNIST dataset, subdivided into the two classes of characters 0-4 and 5-9, or the CIFAR-10 dataset to distinguish between animals and vehicles. For impossible data, we randomize the labels of the training data. Our experiments are computationally heavy, so we generally use small sample sizes, from 2000 to 8000 examples. The hypothesis space is the set of weight vectors for a neural network with ReLU activation functions constrained by a Gaussian prior distribution with $\sigma=5$. Neural network architectures are described in Section C.1.1 of the appendix. To approximately sample the weight vectors in the vicinity of the Gibbs posterior, we use ULA as in (7) or SGLD (Welling & Teh, 2011) with constant step size $\eta$. To ensure reproducibility, we provide the code and experimental results in an anonymous repository at https://anonymous.4open.science/r/Gibbs-Generalization-45F1.
5.1 The loss function $\ell$
Unless $\sigma$ is very small, which would prevent near-zero training errors for random labels, an unbounded loss function like binary cross-entropy entropy (BCE) would cause $\mathbb{E}_{h\sim G_{0}}\big[\hat{L}(h)\big]=\mathbb{E}_{h\sim\pi}\big[\hat{L}(h)\big]$ and therefore the first term in the summation of $\Gamma$ in equation 3 to be very large. For this reason we use a bounded loss function $\ell$, either bounded binary cross-entropy as described in Appendix D of Dziugaite & Roy (2018) or the savage loss (Masnadi-Shirazi & Vasconcelos, 2008). We compute bounds for the 0-1 loss, using the method described in Section 3.3.
5.2 Approximating the ergodic mean
As we know of no sufficient criterion for convergence, we terminate iterations at time $T$, when a very slow running mean $\mathbb{M}_{\rm stop}$ of the loss trajectory $\big(\hat{L}(h_{\beta_{k}t},\mathbf{x})\big)_{t=0}^{T}$ stops decreasing. A second running mean $\mathbb{M}_{\rm erg}$ is used as an approximation of the ergodic mean and thus of expectations in the invariant distribution. We thus replace all expectations $\mathbb{E}_{h\sim G_{\beta_{k}}}\big[\hat{L}(h,\mathbf{x})\big]$ occurring in the bounds by $\mathbb{M}_{\rm erg}\big[\big(\hat{L}(h_{\beta_{k}t},\mathbf{x})\big)_{t=0}^{T}\big]$. Both running means $\mathbb{M}_{\rm stop}$ and $\mathbb{M}_{\rm erg}$ are implemented as first-order, recursive lowpass filters described in Section C.1.3 of the appendix.
5.3 Calibration
The theoretical bounds of Corollary 4.2 in combination with Theorem 4.3 can only serve as guidance for the computation of practical bounds, because the quantities $R$ and $\alpha$ are impossible to estimate in practice. But even if we assume these to be in the order of unity, the bounds are too coarse to distinguish between different temperatures with realistic stepsizes.
A simple calculation shows that $KL\left(G_{\beta},G_{2\beta}\right)\leq\beta\left(\mathbb{E}_{G_{\beta}}\left[\hat{L}\right]-\mathbb{E}_{G_{2\beta}}\left[\hat{L}\right]\right)\leq\beta\mathbb{E}_{G_{\beta}}\left[\hat{L}\right$] (Lemma B.5 in Appendix B.3). By Corollary 4.2 we should therefore have at least $8\eta dR^{2}/\alpha<\mathbb{E}_{G_{\beta}\left(\mathbf{x}\right)}\left[\hat{L}\right$] to distinguish between the expectations in the Gibbs posterior for $\beta$ and $2\beta$. The smallest neural network we use has $d=392,500$. If $\ell$ has values in $\left[0,1\right]$ then $\mathbb{E}_{G_{\beta}\left(\mathbf{x}\right)}\left[\hat{L}\right]\leq 1$, so even if $R$ and $\alpha$ are set to $1$, we would need stepsizes in the order of $10^{-7}$. Safe values of $\eta$, as suggested by the theoretical results in Section 4, are therefore impossible in practice, and the bound has to be adapted to a realistic choice of $\eta$.
The first possibility is to simply ignore the un-estimable error terms in Theorem 4.3, and to compute the bounds from $\Gamma_{\nu}\left(h,\mathbf{x},\mathbf{\beta}\right)$ in place of $\Gamma\left(h,\mathbf{x},\mathbf{\beta}\right)$ as in Theorem 3.2. We show some of these uncalibrated bounds, which already make nontrivial predictions, in Section C.2.6.
A second method uses a single calibration parameter, whose value is computed from the training data. We assume that the computed functional $\Gamma_{\nu}\left(h,\mathbf{x},\mathbf{\beta}\right)$ fails to estimate $\Gamma\left(h,\mathbf{x},\mathbf{\beta}\right)$ by a factor $r\left(\mathbf{x}\right)>0$, which we compute as
$r\left(\mathbf{x}\right)=\min\left\{r:\forall k\in\left[K\right],\text{ }\kappa^{-1}\left(\mathbb{E}_{h\sim\nu_{\beta_{k}}\left(\mathbf{x}\right)}\left[\hat{L}_{01}\left(h,\mathbf{\tilde{x}}\right)\right],\frac{1}{n}\left(r\Gamma_{\nu}\left(h,\mathbf{\tilde{x}},\mathbf{\beta}_{1}^{k}\right)+\ln\frac{2\sqrt{n}}{\delta}\right)\right)\geq\frac{1}{2}\right\},$
where $\mathbf{\tilde{x}}$ is the training set $\mathbf{x}$ with random labels and $\hat{L}_{01}$ the empirical 01-error. The calibration value $r$ is thus the smallest factor of $\Gamma_{\nu}$, for which we obtain a correct upper bound on the 01-error with random labels for all the $\beta_{k}$. By definition of $\Gamma_{\nu}$, replacing it by $r\Gamma_{\nu}$ is equivalent to a multiplicative shift of the temperature scale.
It is a purely experimental finding, that our choice of $r$ leads to correct and surprisingly tight upper bounds on the test error of correctly labeled data in all cases we tried. We emphasize that our calibration procedure depends only on the training data.
5.4 Results
Figure 1: SGLD on MNIST and CIFAR-10 with 8000 training examples, MNIST above and CIFAR-10 below, random labels on the left, correct labels on the right. Both random and true labels are trained with the same algorithm and parameters on a fully connected ReLU network with two hidden layers of 1000 and 1500 units, respectively. The calibration factor for MNIST is 0.19, for CIFAR-10 0.22. Train error, test error and our bound for the Gibbs posterior average of the 0-1 loss are plotted against $\beta$.
Several experiments confirm the validity of the proposed bounds. An example is shown in Figure 1, where a fully connected ReLU-network with two hidden layers of 1000 (respectively 1500) units each is trained with SGLD at inverse temperatures $\beta=$ 0, 500, 1000, 2000, 4000, 8000, 16000, 32000, and 64000. The train error for random labels is about 0.1 (or even less) at $\beta=64000$, where the bound is above 0.5. The test error for correct labels, however, is tightly bounded above.
Notice that for MNIST, which has the tightest bounds, the training error for the true labels is rapidly decreasing from 0.5 to 0.17 at $\beta=500$ and to 0.1 at $\beta=1000$. The more moderate initial decrease for CIFAR-10 corresponds to the tendency to overfit on this more difficult dataset. This confirms the intuition, that good generalization at low temperatures is already announced in the high temperature regime.
We generally found the uncalibrated bounds for ULA tighter than those for SGLD, consistent with the findings of Brosse et al. (2018). Experimental bounds for single draws from the posterior and various other experiments are reported in Section C.2.
6 Conclusion
Using the integral representation of the log-partition function, the Gibbs posterior admits the computation of upper bounds on the true error based on the training data and for any temperature. These bounds are stable under perturbation in the total-variation and Wasserstein metrics, and can be approximated by Langevin Monte Carlo (LMC) algorithms. However, for realistic experiments, the approximations obtained by these algorithms are coarse and require calibration, which leads to rather tight bounds in the interpolation regime of overparametrized neural networks.
The fact that the calibrated bounds are very tight is, at this point, a purely experimental finding, requiring more theoretical investigation in future work.
Acknowledgments
This work was supported in part from the PNRR MUR Project PE000013 CUP J53C22003010006 "Future Artificial Intelligence Research (FAIR)", funded by the European Union – NextGenerationEU and by European Lighthouse of AI for Sustainability, Grant N. 101120237.
Appendix
In this appendix, we summarize a glossary of notation, give additional theoretical results and missing proofs, and provide more information on the numerical experiments, as well as additional experimental results.
Appendix A Table of notation
Notation Brief description Section
$\mathcal{X}$ space of data 2
$\Sigma$ sigma algebra (events) on $\mathcal{X}$ 2
$\mu$ probability of data 2
$n$ sample size 1, 2
$\mathbf{x}$ generic member $\left(x_{1},...,x_{n}\right)\in\mathcal{X}^{n}$, training sample 2
$\mathcal{H}$ hypothesis space 2
$\Omega$ sigma algebra (events) on $\mathcal{H}$ 2
$\ell$ $\ell:\mathcal{H\times X\rightarrow}\left[0,\infty\right)$ loss function 2
$f$ secondary loss function 3.3
$\mathcal{P}\left(\mathcal{H}\right)$ probability measures on $\mathcal{H}$ 2
$\pi$ nonnegative a-priori measure on $\mathcal{H}$ 2
$\sigma$ width of Gaussian prior 4
$L\left(h\right)$ $L\left(h\right)=\mathbb{E}_{x\sim\mu}\left[h\left(x\right)\right]$, expected loss of $h\in\mathcal{H}$ 2
$\hat{L}\left(h,\mathbf{x}\right)$ $\hat{L}\left(h,\mathbf{x}\right)=\left(1/n\right)\sum_{i=1}^{n}\ell\left(h,x_{i}\right)$, empirical loss of $h\in\mathcal{H}$ 2
$\beta$ inverse temperature 1, 2
$Z_{\beta}\left(\mathbf{x}\right)$ partition function 2
$G_{\beta,\pi}\left(\mathbf{x}\right)$ Gibbs posterior with energy $\hat{L}$ and prior $\pi$ 2
$\mathbb{E}_{g\sim G_{\beta}\left(\mathbf{x}\right)}$ posterior expectation 2
$\bm{\beta}$ increasing sequence $\left(\beta_{1}<...<\beta_{K}\right)$ of positive reals 3.1
$\Gamma\left(h,\mathbf{x},\bm{\beta}\right)$ bounding functional 3.1
$F\left(h,\mathbf{x}\right)$ placeholder for generalization gap 3.2
$\kappa$ $kl\left(p,q\right)=p\ln\frac{p}{q}+\left(1-p\right)\ln\frac{1-p}{1-q}$, rel. entropy of Bernoulli variables 2
$KL\left(\rho,\nu\right)$ $\int\left(\ln\frac{d\rho}{d\nu}\right)d\rho$, KL-divergence of $\rho,\nu\in\mathcal{P}\left(\mathcal{H}\right)$ 2, 4.1, 4.2
$d_{TV}\left(\rho,\nu\right)$ total variation distance 2, 4.1, 4.2
$W_{p}\left(\rho,\nu\right)$ $p$-Wasserstein distance 2, 4.1, 4.2
$\Gamma_{\nu}\left(h,\mathbf{x},\bm{\beta}\right)$ LMC approximation of $\Gamma\left(h,\mathbf{x},\bm{\beta}\right)$ 4.2
$\eta$ step size or learning rate 4.1
$\nu_{\beta,\eta}$ invariant measure of LMC approximation of $G_{\beta}$ with step size $\eta$ 4.1
$\nu_{\beta,\eta,t}$ LMC approximation of $G_{\beta}$ with step size $\eta$ at iteration $t$ 4.1
$r({\mathbf{x}})$ calibration factor 5.3
$\mathbf{\tilde{x}}$ randomly labeled data 5.3
$\mathbb{M}_{\rm stop}$$\mathbb{M}_{\rm erg}$ filters for stopping and ergodic mean 5.2, C.1.3
Appendix B Additional results and proofs
B.1 Sub-gaussian losses
The freedom in the choice of $F$ allows a number of bounds to be derived from Theorem 3.2. A centered real random variable $Y$ is called $\sigma$-sub-Gaussian if $\ln\mathbb{E}e^{\lambda Y-\mathbb{E}Y}\leq\lambda^{2}\sigma^{2}/2$ for all $\lambda\in\mathbb{R}$. Now, suppose that for some real function $f$ all the $x\in\mathcal{X\mapsto}f\left(h,x\right)$ are $\sigma$-sub-Gaussian as $x\sim\mu$. Let $\hat{L}(h,\mathbf{x})=\frac{1}{n}\sum_{i=1}^{n}f(h,x_{i})$ and $L(h)=E_{x}[f(h,\mathbf{x})]$. Then, $\mathbf{x}\in\mathcal{X}^{n}\mapsto\hat{L}\left(h,\mathbf{x}\right)$ as $\mathbf{x}\sim\mu^{n}$ is $\sigma/\sqrt{n}$-sub-Gaussian. It is tempting to set $F\left(h,\mathbf{x}\right)=\lambda\left(L\left(h\right)-\hat{L}\left(h,\mathbf{x}\right)\right)$ in Theorem 3.2, divide by $\lambda$ and then optimize over $\lambda$. Unfortunately, the last step is impossible, since the optimal $\lambda$ is data-dependent in its dependence on $\Gamma$ and ruins the exponential moment bound on $F$. A more careful argument establishes the following.
Corollary B.1.
Suppose that for all $h\in\mathcal{H}$ the random variables $x\in\mathcal{X\mapsto}f\left(h,x\right)$ as $x\sim\mu$ are $\sigma$-sub-Gaussian. For $\delta>0$ with probability at least $1-\delta$ as $\mathbf{x}\sim\mu^{n}$ and $h\sim G_{\beta,\pi}\left(\mathbf{x}\right)$ if $\Gamma\left(h,\mathbf{x},\bm{\beta}\right)\geq 1$, then,
$E_{x}\left[f\left(h,x\right)\right]-\frac{1}{n}\sum_{i}f\left(h,x_{i}\right)\leq\sigma\sqrt{\frac{2[\Gamma\left(h,\mathbf{x},\mathbf{\beta}\right)\left(1+1/n\right)+\ln\left(\Gamma\left(h,\mathbf{x},\bm{\beta}\right)\left(n+1\right)/\delta\right)]}{n}}.$
For the proof, we use the following auxiliary result.
Lemma B.2.
(Anthony & Bartlett, 1999, Lemma 15.6) Suppose $\Pr$ is a probability distribution and
$\left\{E\left(\alpha_{1},\alpha_{2},\delta\right):0<\alpha_{1},\alpha_{2},\delta\leq 1\right\}$
is a set of events, such that
(i) For all $0<\alpha\leq 1$ and $0<\delta\leq 1$,
$\Pr\left\{E\left(\alpha,\alpha,\delta\right)\right\}\leq\delta.$
(ii) For all $0<\alpha_{1}\leq\alpha\leq\alpha_{2}\leq 1$ and $0<\delta_{1}\leq\delta\leq 1$
$E\left(\alpha_{1},\alpha_{2},\delta_{1}\right)\subseteq E\left(\alpha,\alpha,\delta\right).$
Then for $0<a,\delta<1$,
$\Pr\bigcup_{\alpha\in\left(0,1\right]}E\left(\alpha a,\alpha,\delta\alpha\left(1-a\right)\right)\leq\delta.$
We put this lemma in a more convenient form.
Lemma B.3.
Let $Y$ and $X\geq 0$ be real random variables, $\psi:\mathbb{R}\times\left(0,1\right)\rightarrow\mathbb{R}$ be increasing in the 1st argument and $\forall C>1$, $\delta\in\left(0,1\right)$,
$\Pr\left\{X\leq C\wedge Y>\psi\left(C,\delta\right)\right\}<\delta.$
Then, for every $\epsilon>0$
$\Pr\left\{X\geq 1\wedge Y>\psi\left(X\left(1+\epsilon\right),\frac{\delta\epsilon}{X\left(1+\epsilon\right)}\right)\right\}.$
Proof.
This follows from Lemma B.2 using the events
$E\left(\alpha_{1},\alpha_{2},\delta\right)=\left\{X\leq\alpha_{2}^{-1}\wedge Y>f\left(\alpha_{1}^{-1},\delta\right)\right\}$
and $a=1/\left(1+\epsilon\right)$. ∎
Proof of Corollary B.1.
Take $F=\lambda\left(L\left(h\right)-\hat{L}\left(h,\mathbf{x}\right)\right)$, so $\ln\mathbb{E}_{\mathbf{x}}\mathbb{E}_{g\sim\pi}\left[e^{F\left(g,\mathbf{x}\right)}\right]\leq\lambda^{2}\sigma^{2}/\left(2n\right)$. From Theorem 3.2 and the properties of sub-Gaussian variables we get with $\lambda=\sigma^{-1}\sqrt{2n\left(C+\ln\left(1/\delta\right)\right)}$ that
$\Pr\left\{\Gamma\left(h,\mathbf{x},\bm{\beta}\right)\leq C\wedge L\left(h\right)-\hat{L}\left(h,\mathbf{x}\right)>\sigma\sqrt{\frac{2\left(C+\ln\left(1/\delta\right)\right)}{n}}\right\}$
$=$ $\Pr\left\{\Gamma\left(h,\mathbf{x},\bm{\beta}\right)\leq C\wedge L\left(h\right)-\hat{L}\left(h,\mathbf{x}\right)>\frac{C+\ln\left(1/\delta\right)}{\lambda}+\frac{\lambda\sigma^{2}}{2n}\right\}$
$=$ $\Pr\left\{\Gamma\left(h,\mathbf{x},\bm{\beta}\right)\leq C\wedge\lambda\left(L\left(h\right)-\hat{L}\left(h,\mathbf{x}\right)\right)>C+\lambda^{2}\sigma^{2}/\left(2n\right)+\ln\left(1/\delta\right)\right\}\leq\delta.$
Substitution in Lemma B.3 with $\psi\left(C,\delta\right)=\sigma\sqrt{2\left(C+\ln\left(1/\delta\right)\right)/n}$, $X=\Gamma\left(h,\mathbf{x},\bm{\beta}\right)$, $Y=L\left(h\right)-\hat{L}\left(h,\mathbf{x}\right)$ and $\epsilon=1/n$ gives Corollary B.1. ∎
B.2 Proofs for Section 4.2
Restatement of Theorem 4.3:
Theorem B.4.
Suppose that $\mathcal{H}=\mathbb{R}^{d}$ and that there are numbers $m,M<\infty$ such that for every $\mathbf{x}$ in $\mathcal{X}^{n}$ and $h\in\mathcal{H}$ we have $\left|\ell\left(h,\mathbf{x}\right)\right|\leq m$ and $\left|F\left(h,\mathbf{x}\right)\right|\leq M$. Let $0=\beta_{0}<\beta_{1}<\cdots<\beta_{K}=\beta$ and $\nu_{\beta_{k}}\left(\mathbf{x}\right)\in\mathcal{P}\left(\mathcal{H}\right)$ be such that $d_{TV}\left(\nu_{\beta_{k}}\left(\mathbf{x}\right),G_{\beta_{k}}\left(\mathbf{x}\right)\right)=\epsilon_{\beta_{k}}$. Then,
(i) with probability at least $1-\delta$ as $\mathbf{x}\sim\mu^{n}$ and $h\sim\nu_{\beta}\left(\mathbf{x}\right)$
$F\left(h,\mathbf{x}\right)$ $\leq$ $\Gamma_{\nu}\left(h,\mathbf{x},\mathbf{\beta}\right)+\ln\mathbb{E}_{\mathbf{x}}\mathbb{E}_{h\sim\pi}\left[e^{F\left(h,\mathbf{x}\right)}\right]+\ln\frac{1}{\delta}$
$+\ln\left(2e^{M+\beta m}\epsilon_{\beta}\right)+\sum_{k=1}^{K}\left(\beta_{k}-\beta_{k-1}\right)m\epsilon_{\beta_{k-1}}.$
(ii) with probability at least $1-\delta$ as $\mathbf{x}\sim\mu^{n}$
$\mathbb{E}_{h\sim\nu_{\beta}\left(\mathbf{x}\right)}\left[F\left(h,\mathbf{x}\right)\right]$ $\leq$ $\mathbb{E}_{h\sim\nu_{\beta}\left(\mathbf{x}\right)}\left[\Gamma_{\nu}\left(h,\mathbf{x},\bm{\beta}\right)\right]+\ln\mathbb{E}_{\mathbf{x}}\mathbb{E}_{h\sim\pi}\left[e^{F\left(h,\mathbf{x}\right)}\right]+\ln\frac{1}{\delta}$
$+\left(M+\beta m\right)\epsilon_{\beta}+\sum_{k=1}^{K}\left(\beta_{k}-\beta_{k-1}\right)m\epsilon_{\beta_{k-1}}.$
Proof.
By the bound on $\ell$ we have
$\Gamma\left(h,\mathbf{x},\bm{\beta}\right)\leq\Gamma_{\nu}\left(h,\mathbf{x},\bm{\beta}\right)+\sum_{k=1}^{K}\left(\beta_{k}-\beta_{k-1}\right)m\epsilon_{\beta_{k-1}}\text{.}$ (8)
(i) From Markov’s inequality we have (in analogy to the proof of Theorem 3.2) with probability at least $1-\delta$ as $x\sim\mu^{n}$ and $h\sim\nu_{\beta}\left(\mathbf{x}\right)$, that
$F\left(h,\mathbf{x}\right)+\beta\hat{L}\left(h,\mathbf{x}\right)+\ln Z_{\beta}\left(\mathbf{x}\right)$
$\leq$ $\ln\mathbb{E}_{\mathbf{x}}\mathbb{E}_{h\sim\nu_{\beta}\left(\mathbf{x}\right)}\left[e^{F\left(h,\mathbf{x}\right)+\beta\hat{L}\left(h,\mathbf{x}\right)+\ln Z_{\beta}\left(\mathbf{x}\right)}\right]+\ln\left(1/\delta\right)$
$\leq$ $\ln\left(\mathbb{E}_{\mathbf{x}}\mathbb{E}_{h\sim G_{\beta}\left(\mathbf{x}\right)}\left[e^{F\left(h,\mathbf{x}\right)+\beta\hat{L}\left(h,\mathbf{x}\right)+\ln Z_{\beta}\left(\mathbf{x}\right)}\right]+e^{M+\beta m}\epsilon_{\beta}\right)+\ln\left(1/\delta\right)$
$=$ $\ln\left(\mathbb{E}_{\mathbf{x}}\mathbb{E}_{h\sim\pi}\left[e^{F\left(h,\mathbf{x}\right)}\right]+e^{M+\beta m}\epsilon_{\beta}\right)+\ln\left(1/\delta\right)$
$\leq$ $\ln\mathbb{E}_{\mathbf{x}}\mathbb{E}_{h\sim\pi}\left[e^{F\left(h,\mathbf{x}\right)}\right]+\ln\left(2e^{M+\beta m}\epsilon_{\beta}\right)+\ln\left(1/\delta\right).$
In the second inequality we used $\ln Z_{\beta}\left(\mathbf{x}\right)\leq 0$ and in the last line we used for $a,b\geq 1$ that $\ln\left(a+b\right)\leq\ln\max\left\{a,b\right\}+\ln 2\leq\ln a+\ln b+\ln 2=\ln a+\ln 2b$. Subtract $\beta\hat{L}\left(h,\mathbf{x}\right)+\ln Z_{\beta}\left(\mathbf{x}\right)$, use (4) and (8).
(ii) Again with Markov’s inequality, with probability at least $1-\delta$ as $x\sim\mu^{n}$
$\mathbb{E}_{h\sim\nu_{\beta}\left(\mathbf{x}\right)}\left[F\left(h,\mathbf{x}\right)+\beta\hat{L}\left(h,\mathbf{x}\right)+\ln Z_{\beta}\left(\mathbf{x}\right)\right]$
$\leq$ $\ln\mathbb{E}_{\mathbf{x}\sim\mu^{n}}\left[e^{\mathbb{E}_{h\sim\nu_{\beta}\left(\mathbf{x}\right)}\left[F\left(h,\mathbf{x}\right)+\beta\hat{L}\left(h,\mathbf{x}\right)+\ln Z_{\beta}\left(\mathbf{x}\right)\right]}\right]$
$\leq$ $\ln\mathbb{E}_{\mathbf{x}\sim\mu^{n}}\left[e^{\mathbb{E}_{h\sim G_{\beta}\left(\mathbf{x}\right)}\left[F\left(h,\mathbf{x}\right)+\beta\hat{L}\left(h,\mathbf{x}\right)+\ln Z_{\beta}\left(\mathbf{x}\right)\right]+\left(M+\beta m\right)\epsilon_{\beta}}\right].$
In the second inequality we used the fact that $\ln Z_{\beta}\left(\mathbf{x}\right)\leq 0$ and the bounds on $F$ and $\ell$. Then, Jensen’s inequality bounds the last line as
$\ln\mathbb{E}_{\mathbf{x}\sim\mu^{n}}\mathbb{E}_{h\sim G_{\beta}\left(\mathbf{x}\right)}\left[e^{F\left(h,\mathbf{x}\right)+\beta\hat{L}\left(h,\mathbf{x}\right)+\ln Z_{\beta}\left(\mathbf{x}\right)}\right]e^{\left(M+\beta m\right)\epsilon_{\beta}}$
$=$ $\ln\mathbb{E}_{\mathbf{x}\sim\mu^{n}}\mathbb{E}_{h\sim\pi}\left[e^{F\left(h,\mathbf{x}\right)}\right]+\left(M+\beta m\right)\epsilon_{\beta}.$
Again subtract $\beta\hat{L}\left(h,\mathbf{x}\right)+\ln Z_{\beta}\left(\mathbf{x}\right)$, use (4) and (8). ∎
B.3 Miscellaneous lemmata
Lemma B.5.
For $0<\beta<\infty$
$\max\left\{KL\left(G_{\beta},G_{2\beta}\right),KL\left(G_{2\beta},G_{2}\right)\right\}\leq\beta\left(\mathbb{E}_{h\sim G_{\beta}}\left[\hat{L}\left(h\right)\right]-\mathbb{E}_{h\sim G_{2\beta}}\left[\hat{L}\left(h\right)\right]\right)$
Proof.
Using Lemma 3.1
$KL\left(G_{\beta},G_{2\beta}\right)$ $=$ $\mathbb{E}_{h\sim G_{\beta}}\left[-\beta\hat{L}\left(h\right)-\ln Z_{\beta}+2\beta\hat{L}\left(h\right)+\ln Z_{2\beta}\right]$
$=$ $\mathbb{E}_{h\sim G_{\beta}}\left[\beta\hat{L}\left(h\right)\right]-\int_{\beta}^{2\beta}\mathbb{E}_{h\sim G_{\gamma}}\left[\hat{L}\left(h\right)\right]d\gamma$
$\leq$ $\beta\left(\mathbb{E}_{h\sim G_{\beta}}\left[\hat{L}\left(h\right)\right]-\mathbb{E}_{h\sim G_{2\beta}}\left[\hat{L}\left(h\right)\right]\right).$
Similarly
$KL\left(G_{2\beta},G_{2}\right)$ $=$ $\mathbb{E}_{h\sim G_{2\beta}}\left[-2\beta\hat{L}\left(h\right)-\ln Z_{2\beta}+\beta\hat{L}\left(h\right)+\ln Z_{\beta}\right]$
$=$ $-\mathbb{E}_{h\sim G_{\beta}}\left[\beta\hat{L}\left(h\right)\right]+\int_{\beta}^{2\beta}\mathbb{E}_{h\sim G_{\gamma}}\left[\hat{L}\left(h\right)\right]d\gamma$
$\leq$ $\beta\left(\mathbb{E}_{h\sim G_{\beta}}\left[\hat{L}\left(h\right)\right]-\mathbb{E}_{h\sim G_{2\beta}}\left[\hat{L}\left(h\right)\right]\right).$
Appendix C Experimental details and additional results
C.1 Experimental details
All the codes to reproduce the results are provided through this https://anonymous.4open.science/r/Gibbs-Generalization-45F1. For all the experiments we use the isotropic Gaussian prior with $\mu=0$ and $\sigma=5$. This induces an L2-regularization term in the energy function that is stated in the proof of Corollary 4.2. The confidence parameter $\delta$ appearing in our bounds is set to $0.01$ for all experiments
We use either standard SGLD or ULA with a constant step size and without additional correction terms. When ULA has been used, we use a step size of 0.01 for both datasets. However with SGLD, we set the step size to 0.01 for MNIST and 0.005 for CIFAR-10. For both datasets, MNIST and CIFAR-10, we use neural networks with ReLU activation functions.
C.1.1 Network architecture
The fully connected networks consist of one, two, three hidden layers, each containing a constant number of units. Besides that we are using LeNet-5 architecture for MNIST and VGG16 architecture for CIFAR-10 to achieve low test error. For loss function $\ell$, we are using either bounded binary cross-entropy (BBCE) as described in Appendix D of Dziugaite & Roy (2018) or the savage loss (Masnadi-Shirazi & Vasconcelos, 2008).
The LeNet-5 network follows a systematic pattern of alternating convolutional and pooling layers, followed by fully connected layers (LeCun et al., 2002). It begins with an input layer that accepts $32\times 32$ grayscale images. Thus, we pad our images to fit. The first convolutional layer (C1) applies 6 filters of size $5\times 5$ to extract low-level features, followed by a $2\times 2$ average pooling layer (S2) for spatial downsampling. The second convolutional layer (C3) uses 16 filters of size $5\times 5$ to capture more complex feature combinations, followed again by a $2\times 2$ average pooling layer (S4). A third convolutional layer (C5) with 120 filters of size $5\times 5$ acts as a feature extractor, producing 120 feature maps, each of size $1\times 1$. The architecture concludes with two fully connected layers: F6 with 84 neurons and a final output layer with 10 neurons for the original digit classification task. However, for our binary classification task, we modify F6 to have 420 neurons and use a single-neuron output layer. Throughout the network, ReLU activation functions replace the original tanh activations, which improves gradient flow and training performance in modern implementations.
VGG-16 is a widely used deep convolutional neural network architecture known for its simplicity and strong performance in image classification tasks (Simonyan & Zisserman, 2014). The architecture follows a consistent design using only $3\times 3$ convolutional filters and $2\times 2$ max pooling operations throughout the network. In our implementation, VGG-16 is adapted to handle CIFAR-10’s smaller $32\times 32$ RGB images. The network consists of 13 convolutional layers organized into five blocks: the first two blocks contain two convolutional layers each with 64 and 128 filters, respectively, while the last three blocks contain three convolutional layers each with 256, 512, and 512 filters, respectively. Each block is followed by a $2\times 2$ max pooling layer for spatial downsampling. All convolutional layers employ $3\times 3$ kernels with padding to preserve spatial dimensions, and ReLU activation functions introduce non-linearity.The convolutional feature extractor is followed by a classifier head consisting of three fully connected layers: two hidden layers with 1024 neurons each, using ReLU activation, and a final output layer with 1 neurons for binary classification. We also removed dropout to ensure that SGLD minimizes the defined energy function without any additional terms.
For MNIST, the input is a 784-dimensional vector, and the output is a scalar since we perform binary classification between digits 0–4 and 5–9. For CIFAR-10, the input dimension is 3072, and the output is again scalar, corresponding to binary classification between vehicles and animals. For evaluating our models, we are using all 10,000 test examples for both datasets.
C.1.2 Minibatches
When using SGLD, we adopt minibatches of size proportional to $\sqrt{n}$. Thus, for $n=2000$ the mini-batch size is 50, and for $n=8000$ it is 100.
C.1.3 Moving average filters
As we explained in Section 5.2, we are using a running mean $\mathbb{M}(x_{1},\cdots,x_{t})$ of $\hat{L}(h_{j},\mathbf{x})$ from $j=1,\cdots,t$ both as a criterion to stop the experiment and an estimation for $\mathbb{E}_{h\sim G_{\beta_{k}}}\left[\hat{L}\left(h,\mathbf{x}\right)\right]$. We define the running mean recursively in one of two ways:
$\mathbb{M}_{t}$ $=\tfrac{\alpha}{2}\hat{L}(h_{t},\mathbf{x})+\tfrac{\alpha}{2}\hat{L}(h_{t-1},\mathbf{x})+(1-\alpha)\,\mathbb{M}_{t-1},$
$\mathbb{M}_{t}$ $=\alpha\hat{L}(h_{t},\mathbf{x})+(1-\alpha)\,\mathbb{M}_{t-1},$
with $\mathbb{M}_{0}=0$ and small $\alpha$. We use the first (symmetric) form in the experiments with ULA, and the second (standard exponential moving average) form with SGLD for convenience. We set different values of $\alpha$ for the two roles: $\alpha=0.0025$ for the stopping criterion ($\mathbb{M}_{\mathrm{stop}}$) and $\alpha=0.01$ for approximating the ergodic mean ($\mathbb{M}_{\mathrm{erg}}$). The stopping rule is triggered when
$\mathbb{M}_{t}-\mathbb{M}_{t-1}\geq\epsilon,$
with $\epsilon=10^{-7}$. To avoid premature termination, we impose a minimum of 4000 steps before applying this criterion. As $\alpha\to 0$ and $t\to\infty$, the quantity $\mathbb{M}_{t}$ converges to the ergodic mean.
C.2 Experimental results
C.2.1 Illustration of Bound Computation
In this section, we demonstrate again the figure in the main body in more details. The figure 2 illustrates how our bounds are computed. The sequence of mean training losses in $\ell$ is used to compute for each $\beta$ the functional $\Gamma$ and the "KL-Bound", which corresponds to the right hand side of the inequalities in Corollary 3.3. Our bound on the test loss is then computed by applying the function $\kappa^{-1}$ to the empirical 0-1 error and to this kl-bound. The graph of "KL(Train, Test)" corresponds to the left hand side in Corollary 3.3.
It is remarkable that the close fit of the upper bound on the random labels is achieved by the adjustment of a single calibration parameter.
Figure 2: A more detailed version of Figure 1 to illustrate how the bounds are computed.
C.2.2 Single-draws
For the setting described in Section 5.4, we also present the bounds for the single-draw case in Figure 3. It is noteworthy that, although the theoretical guarantees for this scenario are rather weak, the empirical bounds behave well. However, as visible in the plots, the results exhibit fluctuations and irregularities caused by stochastic effects, which make them less reliable.
Figure 3: SGLD on MNIST and CIFAR-10 with 2000 training examples using BBCE loss function. The first row corresponds to MNIST and the second row to CIFAR-10. Random labels are shown on the left, correct labels on the right. Both random and true labels are trained with exactly the same algorithm and parameters on a fully connected ReLU network with two hidden layers of 1000 (respectively 1500) units. The calibration factor for MNIST is 0.19, for CIFAR-10 0.22. Train error, test error and our bound for a single-draw of the 0-1 loss are plotted against $\beta$.
C.2.3 Different Architectures
In this section, we evaluate the performance of different models and architectures on both MNIST and CIFAR-10, demonstrating that our bound can be used to guide model selection. In addition to the two-hidden-layer neural networks described in Section 1, we consider fully connected neural networks with three hidden layers, containing 500 and 1000 units for MNIST and CIFAR-10, respectively. Furthermore, we employ the LeNet-5 architecture for MNIST and VGG-16 for CIFAR-10 to achieve high test accuracy. Detailed descriptions of these architectures are provided in Section C.1.1.
Figure 4 demonstrates the robustness of our bound across different models. We observe that the bounds can be very tight even when the test error is small. For convolutional neural networks, especially on the MNIST dataset, we observe strong performance with the true labels, but relatively poor performance with random labels, despite having more parameters than training examples. This can be explained by the fact that convolutional architectures are still far from being highly overparameterized. For the MNIST dataset, we use fully connected neural networks with two or three hidden layers, containing 1000 or 500 units per layer, respectively. This corresponds to a total of approximately 1,787,000 and 893,000 parameters, resulting in a parameter-to-training-example ratio of roughly 200 and 100, respectively. In contrast, LeNet-5 has around 100,000 parameters, yielding a ratio of approximately 12.5.
Figure 4: SGLD on MNIST and CIFAR-10 with 8000 training examples using BBCE loss function. The first two rows correspond to MNIST, and the remaining rows to CIFAR-10. Random labels are shown on the left, and correct labels on the right. Both random and true labels are trained using the same algorithm and hyperparameters on a fully connected ReLU network with three hidden layers of 500 (MNIST) or 1000 (CIFAR-10) units, followed by LeNet-5 (MNIST) or VGG-16 (CIFAR-10) shown in the subsequent row. The calibration factors for MNIST are 0.26 and 0.08, for CIFAR-10 0.24 and 0.18. The training error, test error, and our bound for the Gibbs posterior average of the 0–1 loss are plotted against $\beta$.
The empirical test bounds can serve as a selection criterion among different models. Table 1 show that test bounds at low temperature are useful for model selection, and that bounds at high temperature can also predict the behavior of the model at low temperature.
2HL (W=1000) 3HL (W=500) LeNet-5
Test Bound at $\beta=1k$ 0.1766 0.2347 0.0887
Test Error at $\beta=64k$ 0.0498 0.0549 0.0317
Test Bound at $\beta=64k$ 0.0860 0.1314 0.0375
(a) MNIST, 8k training examples (true labels).
2HL (W=1500) 3HL (W=1000) VGG-16
Test Bound at $\beta=1k$ 0.2905 0.3635 0.2330
Test Error at $\beta=64k$ 0.1719 0.1782 0.0903
Test Bound at $\beta=64k$ 0.2266 0.2807 0.2030
(b) CIFAR-10, 8k training examples (true labels).
Table 1: Test bounds and test errors for different neural network architectures on MNIST and CIFAR-10. The bounds at both low and high temperatures reliably reflect test error performance at low temperature.
C.2.4 ULA
We have also conducted experiments using ULA for both datasets. The main difference from SGLD is that we use all the information to compute the gradient at each step. The results are shown in Figure 5.
Figure 5: ULA on MNIST and CIFAR-10 with 2000 training examples using BBCE loss function. The first row corresponds to MNIST and the second row to CIFAR-10. Random labels are shown on the left, correct labels on the right. Both random and true labels are trained with the same algorithm and parameters on a fully connected ReLU network with one (respectively two) hidden layers of 500 (respectively 1000) units. The calibration factor for MNIST is 0.49, for CIFAR-10 0.46. Train error, test error and our bound for the Gibbs posterior average of the 0-1 loss are plotted against $\beta$.
C.2.5 Savage Loss function
We additionally performed experiments using the Savage loss to verify the robustness of our results across different loss functions. Following the same setup as in the previous section, the outcomes are reported in Figure 7.
Figure 6: ULA on MNIST and CIFAR-10 with 2000 training examples using savage loss function. The first row corresponds to MNIST and the second row to CIFAR-10. Random labels are shown on the left, correct labels on the right. Both random and true labels are trained with the same algorithm and parameters on a fully connected ReLU network with one (respectively two) hidden layers of 500 (respectively 1000) units. The calibration factor for MNIST is 0.49, for CIFAR-10 0.59. Train error, test error and our bound for the Gibbs posterior average of the 0-1 loss are plotted against $\beta$.
C.2.6 Uncalibrated Bounds
We show uncalibrated bounds for the MNIST dataset with the BBCE and Savage loss functions under the experimental conditions of Sections C.2.4 and C.2.5. The bounds are somewhat looser than the calibrated ones, but still far from trivial. As in all other scenarios the test errors are upper bounded correctly.
Figure 7: ULA on MNIST with 2000 training examples using BBCE and savage loss functions. The first row corresponds to MNIST with BBCE and the second row to savage. Random labels are shown on the left, correct labels on the right. Both random and true labels are trained with the same algorithm and parameters on a fully connected ReLU network with one hidden layers of 500 units. Train error, test error and our bound for the Gibbs posterior average of the 0-1 loss are plotted against $\beta$ without any calibration.
C.2.7 Real-world use cases
We further evaluated Stochastic Gradient Descent (SGD) to examine the practical relevance of our bounds in real-world interpolation regimes.
Based on our observations, we suggest the following procedure for practitioners who wish to train overparameterized neural networks with standard SGD while also obtaining generalization guarantees. First, randomly permute the labels, train the network at different temperatures, and compute the bound together with the calibration factor. Then, repeat the same procedure using the true labels. At very low temperatures, this approach provides generalization guarantees that may transfer to SGD. The corresponding results are presented in Table 2.
2HL (W=1000) 3HL (W=500) LeNet-5
Test Error with SGD 0.0364 0.0363 0.0308
Test Error at $\beta=64k$ 0.0498 0.0549 0.0317
Test Bound at $\beta=64k$ 0.0860 0.1314 0.0375
(a) MNIST, 8k training examples (true labels).
2HL (W=1500) 3HL (W=1000) VGG-16
Test Error with SGD 0.1423 0.1415 0.0933
Test Error at $\beta=64k$ 0.1719 0.1782 0.0903
Test Bound at $\beta=64k$ 0.2266 0.2807 0.2030
(b) CIFAR-10, 8k training examples (true labels).
Table 2: Comparing SGD test error with SGLD test errors and bounds for different neural network architectures on MNIST and CIFAR-10.

### v2

@@
 Abstract
-The paper provides data-dependent bounds on the test error of the Gibbs algorithm in the overparameterized interpolation regime, where low training errors are also obtained for impossible data, such as random labels in classification.
+This paper provides data-dependent bounds on the expected error of the Gibbs algorithm in the overparameterized interpolation regime, where low training errors are also obtained for impossible data, such as random labels in classification.
+The results show that generalization in the low-temperature regime is already signaled by small training errors in the noisier high-temperature regime.
 The bounds are stable under approximation with Langevin Monte Carlo algorithms.
-Experiments on the MNIST and CIFAR-10 datasets verify that the bounds yield nontrivial predictions on true labeled data and correctly upper bound the test error for random labels.
-Our method indicates that generalization in the low-temperature, interpolation regime is already signaled by small training errors in the more classical high temperature regime.
+The analysis motivates the design of an algorithm to compute bounds, which on the MNIST and CIFAR-10 datasets yield nontrivial, close predictions on the test error for true labeled data, while maintaining a correct upper bound on the test error for random labels.
 1 Introduction
 Modern learning algorithms can achieve very small training errors on arbitrary data if the underlying hypothesis space is large enough.
-For reasonable data originating from real-world problems, the chosen hypotheses also tend to have small test errors, a fortunate circumstance, which has given great technological and economic thrust to deep learning.
+For meaningful data, the chosen hypotheses also tend to have small test errors, a fortunate circumstance, which has given great technological and economic thrust to deep learning.
 Unfortunately, the same algorithms also achieve very small training errors for data specifically designed to produce very large test errors, such as random labels in classification.
-Consequently, the hypothesis space and the training error do not suffice to predict the test error.
+In such a situation, which we will loosely refer to as the interpolation regime, the hypothesis space and the training error do not suffice to predict the test error.
 The key to generalization must be more deeply buried in the data.
-While not so disquieting to practitioners, this mystery has troubled theoreticians for many years (Zhang et al., 2016; 2021), and it seems safe to say that the underlying mechanisms still have not been completely understood.
-We are far from solving this riddle in generality, but for the Gibbs posterior we show how nontrivial bounds on the test error can be recovered from the training data.
-The Gibbs posterior assigns probabilities, which decrease exponentially with the training error of the hypotheses.
+While not so disquieting to practitioners, this mystery has troubled theoreticians for many years (Zhang et al., 2016, 2021), and it seems safe to say that the underlying mechanisms still have not been completely understood.
+We are far from solving this riddle in generality, but for sufficiently close approximations of the Gibbs posterior we show how nontrivial bounds on the test error can be recovered from the training data.
+The Gibbs posterior assigns probabilities that decrease exponentially with the training error of the hypotheses.
 The exponential decay parameter $\beta$ can be interpreted as an inverse temperature in an analogy to statistical physics.
-The Gibbs measure is a sufficient idealization to have tractable theoretical properties, but also the limiting distribution of several concrete stochastic algorithms, here summarized as Langevin Monte Carlo (LMC), including Stochastic Gradient Langevin Dynamics (SGLD), (Gelfand & Mitter, 1991; Welling & Teh, 2011), a popular modern learning algorithm.
+The Gibbs measure is a sufficient idealization to have tractable theoretical properties, but it is also the limiting distribution of several concrete stochastic algorithms, here summarized as Langevin Monte Carlo (LMC), including Stochastic Gradient Langevin Dynamics (SGLD), (Gelfand and Mitter, 1991; Welling and Teh, 2011), a popular modern learning algorithm.
 When $\beta$ is large and the hypothesis space is rich, these algorithms can reproduce the dilemma described above by achieving very small training errors on data designed to have large test errors.
-Our paper addresses this interpolation regime of the Gibbs posterior and makes the following three contributions:
-  • We give high-probability data-dependent bounds on the test error, both for a hypothesis drawn from the Gibbs posterior and for the posterior mean, assuming that we can freely draw samples from it.
-These bounds holds for the entire range of temperatures.
-  • We weaken the above assumption by showing that the bounds are stable under approximations of the posterior in the total variation and $W_{2}$-Wasserstein metrics.
-Given sufficient computing resources, this yields bounds for LMC algorithms.
-  • Using LMC on real-world classification data (MNIST and CIFAR-10), we obtain non-trivial upper bounds on the test error for true labels and upper bounds for random labels, despite the algorithms achieving very small training error on the random labels.
-Our method is based on a combination of the PAC-Bayesian bounds (McAllester, 1999; Alquier, 2021; Rivasplata et al., 2020) with an integral representation of the log-partition function.
-This makes it possible to bound the logarithm of the density of the Gibbs posterior at a given temperature in terms of empirical averages at higher temperatures.
-A qualitative conclusion is that generalization in the under-regularized low-temperature regime ($\beta>n$) is already indicated by small training errors in the over-regularized high-temperature regime ($\beta<n$), where $n$ is the number of training examples.
+Our paper addresses this regime of the Gibbs posterior and makes the following three contributions:
+  • We give high-probability data-dependent bounds on the true error, both for a hypothesis drawn from the Gibbs posterior and for the posterior mean, assuming that we can freely draw samples from it.
+These bounds hold for the entire range of temperatures.
+  • We show that these bounds are stable under approximations of the Gibbs posterior in relative entropy.
+Given enough computing resources, this yields bounds for LMC algorithms, based on known results for non-convex sampling.
+  • Existing convergence guarantees for LMC are insufficient for both a practical and rigorous computation of these bounds on real-world problems.
+A heuristic calibration step, based exclusively on the training data, leads to very close upper bounds on the test error for various neural networks trained with LMC on the MNIST and CIFAR-10 datasets.
+The idea underlying our bound is the following.
+The PAC-Bayesian theorem or its single draw variant (McAllester, 1999; Alquier, 2024; Rivasplata et al., 2020) bounds the generalization error roughly proportional to the logarithm of the posterior density or its posterior expectation (the relative entropy to the prior).
+The log-density of the Gibbs posterior at inverse temperature $\beta$ has an explicit expression in terms of an integral from 0 to $\beta$ of mean training errors, a fact which seems to have been overlooked in the PAC-Bayesian analysis of generalization (Lemma 3.1 below).
+Substitution of this integral in the PAC-Bayesian theorem then gives a bound on the generalization error.
+As an illustrative example: if the loss $\ell$ has values in $\left[0,1\right]$, and we happen to draw from the Gibbs posterior at $\beta$ a hypothesis $h$ with training error $\hat{L}\left(h,\mathbf{x}\right)=0$, then we have the following bound on the expected error of this hypothesis.
+$E_{x}\left[\ell\left(h,x\right)\right]\leq\frac{2\left(A+\ln\left(2\sqrt{n}/\delta\right)\right)}{n},$ (1)
+where $A$ is the area in Figure 1, $n$ is the sample size and $\delta$ is the confidence parameter.
+Figure 1: The mean training error of the Gibbs posterior is plotted against the inverse temperature.
+If the training error $\hat{L}\left(h,\mathbf{x}\right)$ is zero, then the log density, $\ln\frac{dG_{\beta}\left(\mathbf{x}\right)}{d\pi}\left(h\right)$, is equal to the area AA.
+For the Gibbs posterior this simple, but novel reasoning resolves the dilemma of the interpolation regime.
+The training error of sufficiently overparametrized systems at a large value of $\beta$ (low temperature) is typically near zero and does not distinguish between easy and hard (e.g.
+random-label) data, but the mean training losses at small values of $\beta$ (high temperatures) will be quite different, leading to different predictions also at large $\beta$ (low temperature).
+Paraphrased:
+Better generalization in the low-temperature regime is already indicated by smaller training errors in the high-temperature regime.
+Figure 2: Mean training error of a fully-connected neural network with 400,000 parameters trained with SGLD on 2,000 examples of MNIST, both with true and random labels.
+This principle appears to hold also for the distributions generated by practical LMC algorithms trained on real-world data, as is witnessed by Figure 2.
+In Section 5.4 we use the area ratio $A/\bar{A}$ and the fact that the true classification error for random labels is $1/2$, to develop a practical method to compute bounds in realistic environments.
+After a brief survey of related literature, we review the PAC-Bayesian theorem, introduce the Gibbs posterior and present our bounds, followed by a stability analysis.
+We conclude with a section describing our experiments.
+The appendix contains a glossary of notation with section references.
 1.1 Related Literature
@@
 Most similar to this work is Raginsky et al.
-(2017), which bounds the distance to the Gibbs posterior and then its generalization error.
-Their bound, however, applies only to the high temperature regime $\beta<n$.
+(2017), which bounds the distance of the distribution generated by SGLD to the Gibbs posterior and then the latter’s generalization error.
+Their bound applies only to the high-temperature regime $\beta<n$, but their convergence guarantees can be combined with our method to give bounds for SGLD in the entire temperature range.
 Several works concentrate on the optimization path of SGLD.
 Mou et al.
-(2018) gives both stability and PAC-Bayesian bounds.
+(2018) give both stability and PAC-Bayesian bounds.
 Pensia et al.
-(2018) applies the information theoretic generalization bounds of Xu & Raginsky (2017).
+(2018) apply the information-theoretic generalization bounds of Xu and Raginsky (2017).
 These ideas are further developed by Negrea et al.
 (2019), where random subsets of the training data are used to define data-dependent priors.
-Farghly & Rebeschini (2021) gives time-independent bounds for SGLD, which are further improved by Futami & Fujisawa (2024).
+Farghly and Rebeschini (2021) give time-independent bounds for SGLD, which are further improved by Futami and Fujisawa (2024).
 Most of the bounds in the above papers are in expectation.
@@
 If the invariant distribution is the Gibbs posterior, the bound along the entire optimization path is of order $\sqrt{\beta/n}$ but improvable to $\beta/n$.
-Some papers give similar bounds for the Gibbs posterior, roughly of the form $\beta/n$ or $\sqrt{\beta/n}$ (Raginsky et al., 2017; Dziugaite & Roy, 2018; Kuzborskij et al., 2019; Rivasplata et al., 2020) or Maurer (2024) and Harel et al.
-(2025)).
+Some papers give similar bounds for the Gibbs posterior, roughly of the form $\beta/n$ or $\sqrt{\beta/n}$ (Raginsky et al., 2017; Dziugaite and Roy, 2018; Kuzborskij et al., 2019; Rivasplata et al., 2020; Maurer, 2024; Harel et al., 2025)).
 These bounds hold equally for random labels and are therefore vacuous for overparametrized hypothesis spaces in the low temperature regime $\beta>n$.
-To our knowledge, ours is the only bound for the Gibbs posterior, which is valid in this regime.
+To our knowledge, ours is the only bound for the Gibbs posterior which is valid in this regime.
 Other bounds have been developed for specific algorithms designed to optimize them.
-The milestone paper by Dziugaite & Roy (2017) is the most prominent example, and (Dziugaite & Roy, 2018) and (Pérez-Ortiz et al., 2021) are also in this category.
-Our bounds by contrast apply to the Gibbs posterior and LMC in their standard forms.
-2 Preliminaries
-The relative entropy of two Bernoulli variables with expectations $p$ and $q$ is denoted
-$\kappa\left(p,q\right)=p\ln\frac{p}{q}+\left(1-p\right)\ln\frac{1-p}{1-q}.$ (1)
-We also define the function $\kappa^{-1}:\left[0,1\right]\times\left[0,\infty\right)\rightarrow\left[0,1\right]$ by
-$\kappa^{-1}\left(p,t\right)=\inf\left\{q:q\geq p,\kappa\left(p,q\right)\geq t\right\}.$
-Throughout the following $\left(\mathcal{X},\Sigma\right)$ is a measurable space of data with probability measure $\mu$. The iid random vector $\mathbf{x}\sim\mu^{n}$ is the training sample.
-We let $\left(\mathcal{H},\Omega\right)$ be a measurable space of hypotheses, and let $\ell:\mathcal{H\times X}\rightarrow\left[0,\infty\right)$ be a prescribed loss function.
+The milestone paper by Dziugaite and Roy (2017) is the most prominent example; Zhou et al.
+(2018), Dziugaite and Roy (2018) and Pérez-Ortiz et al.
+(2021) are also in this category.
+Our bounds, by contrast, apply to the Gibbs posterior and LMC in their standard forms.
+2 The PAC-Bayesian Bound
+Throughout the following $\left(\mathcal{X},\Sigma\right)$ is a measurable space of data with probability measure $\mu$. The i.i.d. random vector $\mathbf{x}\sim\mu^{n}$ is the training sample.
+$\left(\mathcal{H},\Omega\right)$ is a measurable space of hypotheses, and $\ell:\mathcal{H\times X}\rightarrow\left[0,\infty\right)$ is a prescribed loss function.
 Members of $\mathcal{H}$ are denoted $h$ or $g$.
+For a function $f:\mathcal{H}\to\mathbb{R}$, the sup-norm is denoted $f||_{\infty}$.
 We write $L\left(h\right):=\mathbb{E}_{x\sim\mu}\left[\ell\left(h,x\right)\right]$ and $\hat{L}\left(h,\mathbf{x}\right):=\left(1/n\right)\sum_{i}\ell\left(h,x_{i}\right)$ respectively for the true (expected) and empirical error of hypothesis $h\in\mathcal{H}$.
 The set of probability measures on $\left(\mathcal{H},\Omega\right)$ is denoted $\mathcal{P}\left(\mathcal{H}\right)$.
+The relative entropy or Kullback-Leibler-divergence between two probability measures is the function $\text{KL}:(\rho,\nu)\in\mathcal{P}(\mathcal{H})\times\mathcal{P}(\mathcal{H})\mapsto\mathbb{E}_{h\sim\rho}\big[\ln\frac{d\rho}{d\nu}(h)\big]$ if $\rho$ is absolutely continuous w.r.t.
+$\nu$, otherwise the value is $\infty$.
+The Rényi-infinity divergence (Rényi, 1961) is $R_{\infty}(\rho,\nu)=\sup_{h\in\mathcal{H}}\ln\frac{d\rho}{d\nu}(h)$ for $\nu,\rho\in\mathcal{P}(\mathcal{H})$.
+There is an a-priori reference measure $\pi\in\mathcal{P}\left(\mathcal{H}\right)$, called the prior.
 A stochastic algorithm is a function $\nu:\mathcal{X}^{n}\rightarrow\mathcal{P}\left(\mathcal{H}\right)$, which assigns to a training sample $\mathbf{x}$ a probability measure $\nu\left(\mathbf{x}\right)\in\mathcal{P}\left(\mathcal{H}\right)$.
-The $KL$-divergence between two probability measures is the function $KL:(\rho,\nu)\in\mathcal{P}(\mathcal{H})\times\mathcal{P}(\mathcal{H})\mapsto\mathbb{E}_{h\sim\rho}\big[\ln\frac{d\rho}{d\nu}\big]$ if $\rho$ absolutely continuous w.r.t.
-$\nu$, otherwise the value is $\infty$.
-The total variation distance is defined as $d_{TV}:(\rho,\nu)\in\mathcal{P}(\mathcal{H})\times\mathcal{P}(\mathcal{H})\mapsto\sup_{A\in\Omega}|\rho(A)-\nu(A)|$.
-The $W_{p}$-Wasserstein distance is $W_{p}(\rho,\nu)=(\inf_{W}\mathbb{E}_{(x,y)\sim W}[\|x-y\|^{p}])^{1/p}$ with the infimum over all probability measures on $\mathcal{P}(\mathcal{H})\times\mathcal{P}(\mathcal{H})$ with $\rho$ and $\nu$ as marginals.
-There is an a-priori reference measure $\pi\in\mathcal{P}\left(\mathcal{H}\right)$, called the prior.
+The following general version of the PAC-Bayesian theorem appears in this form for the first time in Rivasplata et al.
+(2020).
+It gives a bound for single hypotheses drawn from the posterior (i) as well as for posterior averages (ii).
+Appendix B.1 gives an easy proof for the reader’s benefit.
+Theorem 2.1.
+Let $F:\mathcal{H}\times\mathcal{X}^{n}\rightarrow\mathbb{R}$ be some measurable function, and let $\nu$ be a stochastic algorithm such that $\nu\left(\mathbf{x}\right)$ is absolutely continuous w.r.t.
+$\pi$ for all $\mathbf{x}\in\mathcal{X}^{n}$.
+Then
+(i) for $\delta>0$ with probability at least $1-\delta$ in $\mathbf{x}\sim\mu^{n}$ and $h\sim\nu\left(\mathbf{x}\right)$
+$F\left(h,\mathbf{x}\right)\leq\ln\frac{d\nu\left(\mathbf{x}\right)}{d\pi}\left(h\right)+\ln\mathbb{E}_{\mathbf{x}}\mathbb{E}_{g\sim\pi}\left[e^{F\left(g,\mathbf{x}\right)}\right]+\ln\left(\frac{1}{\delta}\right)$
+(ii) for $\delta>0$ with probability at least $1-\delta$ in $\mathbf{x}\sim\mu^{n}$
+$\mathbb{E}_{h\sim\nu\left(\mathbf{x}\right)}\left[F\left(h,\mathbf{x}\right)\right]\leq\text{KL}\left(\nu\left(\mathbf{x}\right),\pi\right)+\ln\mathbb{E}_{\mathbf{x}}\mathbb{E}_{g\sim\pi}\left[e^{F\left(g,\mathbf{x}\right)}\right]+\frac{1}{\delta}$
+Here $F$ is a placeholder for a random variable related to the generalization gap, which we want to bound.
+Suppose $\ell$ has values in $\left[0,1\right]$.
+With a suitable choice of $F$, we can use (i) above to derive, with probability at least $1-\delta$ as $\mathbf{x}\sim\mu^{n}$ and $h\sim\nu\left(\mathbf{x}\right),$ that
+$L\left(h\right)$ $\!\!\!\!\!\leq\!\!\!\!\!$ $\hat{L}\left(h,\mathbf{x}\right)+\sqrt{\frac{2\hat{L}\left(h,\mathbf{x}\right)}{n}\left(\ln\frac{d\nu\left(\mathbf{x}\right)}{d\pi}+\ln\frac{2\sqrt{n}}{\delta}\right)}+\frac{2}{n}\left(\ln\frac{d\nu\left(\mathbf{x}\right)}{d\pi}+\ln\frac{2\sqrt{n}}{\delta}\right).$ (2)
+[…]
+[…]
+3 Bounds for the Gibbs Algorithm
 With a fixed prior, the Gibbs algorithm at inverse temperature $\beta>0$ is the stochastic algorithm $G_{\beta}:\mathbf{x}\in\mathcal{X}^{n}\mapsto G_{\beta}\left(\mathbf{x}\right)\in\mathcal{P}\left(\mathcal{H}\right)$ defined by
 $G_{\beta}\left(\mathbf{x}\right)\left(A\right)=\frac{1}{Z_{\beta}\left(\mathbf{x}\right)}\int_{A}e^{-\beta\hat{L}\left(h,\mathbf{x}\right)}d\pi\left(h\right)\text{ for }A\in\Omega\text{.}$
-$G_{\beta}\left(\mathbf{x}\right)$ is called the Gibbs-posterior, the normalizing factor
+$G_{\beta}\left(\mathbf{x}\right)$ is called the Gibbs posterior, the normalizing factor
 $Z_{\beta}\left(\mathbf{x}\right):=\int_{\mathcal{H}}e^{-\beta\hat{L}\left(h,\mathbf{x}\right)}d\pi\left(h\right)$
-is called the partition function. The motivation for the Gibbs posterior is that it puts larger weights on hypotheses with smaller empirical error.
+is called the partition function.
+The Gibbs posterior provides a principled, albeit idealized, way to put larger weights on hypotheses with smaller empirical error.
+It appears as the minimizer of the PAC-Bayesian bound in various forms (see McAllester (1999); Guedj (2019); Alquier (2024)).
+As $\beta\rightarrow\infty$, the Gibbs posterior concentrates on the set of empirical risk minimizers (Athreya and Hwang, 2010), so the low-temperature regime is particularly interesting.
+Evidently $\ln\left(dG_{\beta}\left(\mathbf{x}\right)/d\pi\right)\left(h\right)=-\beta\hat{L}\left(h,\mathbf{x}\right)-\ln Z_{\beta}\left(\mathbf{x}\right)$.
+This function has an important integral representation, which is well known from statistical mechanics (see, e.g.
+Huang (2008)).
+Despite its simplicity, it seems to have been overlooked in the literature on generalization.
+Lemma 3.1.
+For all $\beta\geq 0$, $\mathbf{x}\in\mathcal{X}^{n}$ and $h\in\mathcal{H}$
+$\!\!\!-\ln Z_{\beta}\left(\mathbf{x}\right)$ $=$ $\int_{0}^{\beta}\mathbb{E}_{g\sim G_{\gamma}\left(\mathbf{x}\right)}\left[\hat{L}\left(g,\mathbf{x}\right)\right]d\gamma$ (3)
+$\ln\frac{dG_{\beta}\left(\mathbf{x}\right)}{d\pi}\left(h\right)$ $=$ $\int_{0}^{\beta}\left(\mathbb{E}_{g\sim G_{\gamma}\left(\mathbf{x}\right)}\left[\hat{L}\left(g,\mathbf{x}\right)\right]-\hat{L}\left(h,\mathbf{x}\right)\right)d\gamma$ (4)
+$\text{KL}\left(G_{\beta}\left(\mathbf{x}\right),\pi\right)$ $=$ $\int_{0}^{\beta}\left(\mathbb{E}_{g\sim G_{\gamma}\left(\mathbf{x}\right)}\left[\hat{L}\left(g,\mathbf{x}\right)\right]-\mathbb{E}_{h\sim G_{\beta}\left(\mathbf{x}\right)}\left[\hat{L}\left(h,\mathbf{x}\right)\right]\right)d\gamma.$ (5)
+Also the function $\beta\mapsto\mathbb{E}_{g\sim G_{\beta}(\mathbf{x})}\big[\hat{L}(g,\mathbf{x})\big]$, if not constant for all $\beta$, is strictly decreasing in $\beta$.
+Proof.
+Let $A(\beta)=-\ln Z_{\beta}(\mathbf{x})$. One verifies the identities
+$A(0)$ $=$ $0,$
+$A^{\prime}(\beta)$ $=$ $\frac{1}{Z_{\beta,\pi}(\mathbf{x})}\int_{\mathcal{H}}\hat{L}(h,\mathbf{x})e^{-\beta\hat{L}(h,\mathbf{x})}d\pi(h)=\mathbb{E}_{h\sim G_{\beta}(\mathbf{x})}\big[\hat{L}(h,\mathbf{x})\big],$
+$A^{\prime\prime}(\beta)$ $=$ $-\text{Var}_{h\sim G_{\beta}(\mathbf{x})}\big[\hat{L}(h,\mathbf{x})\big]\leq 0,$
+where Var denotes variance.
+(3) then follows from the first two of these identities and the fundamental theorem of calculus, and the last assertion in the lemma follows from the last identity, since zero variance at any $\beta$ implies that $\hat{L}\left(h,\mathbf{x}\right)$ is almost surely constant with respect to $\pi$.
+Since the logarithm of the density of the Gibbs posterior is $\beta\hat{L}\left(.,\mathbf{x}\right)-\ln Z_{\beta}\left(\mathbf{x}\right)$ we get (4).
+Then (5) follows from taking the expectation of (4) in $G_{\beta}\left(\mathbf{x}\right)$.
+∎
+Figure 3: The mean training error of the Gibbs posterior is plotted against the inverse temperature.
+The relative entropy $KL\left(G_{\beta}\left(\mathbf{x}\right),\pi\right)$ is equal to the shaded area AA.
+Figure 3 provides a simple geometrical interpretation for conclusion (5).
+For conclusion (4), if the shaded area were equal to $\ln(dG_{\beta}\left(\mathbf{x}\right)/d\pi)\left(h\right)$, its lower boundary would fluctuate with a variance equal to the negative slope at $\beta$, depending on the draw of $h$ (compare the proof of Lemma 3.1).
+Equations (4) and (5) of Lemma 3.1 can now be substituted into Theorem 2.1 to give our bounds for the Gibbs algorithm.
+4 Bounds for Langevin Monte Carlo
+The Gibbs posterior is an idealization, from which it is impossible to sample directly.
+Nevertheless, a number of works describe Markov processes, here summarized as Langevin Monte Carlo (LMC), capable of approximating a probability measure $\nu$ on $\mathbb{R}^{d}$ of the form $\nu\propto\exp\left(-V\right)$, or some nearby limiting distribution.
+The classical prototype is Continuous Langevin Dynamics (CLD), a Markov process in continuous time describing thermalization in statistical physics and originating in the study of Brownian motion (Langevin and others, 1908).
+To turn the continuous process into an iterative algorithm several discretized versions have been proposed.
+There is the Metropolis Adjusted Langevin Algorithm (MALA) (Roberts and Tweedie, 1996), which is the Euler-discretization of CLD with an additional Metropolis-style accept-reject step to ensure that the invariant distribution is indeed the desired Gibbs distribution.
+Hamiltonian Monte Carlo Markov Chain (HMCMC) (Duane et al., 1987; Betancourt, 2017) is a refinement of MALA.
+The Unadjusted Langevin Algorithm (ULA) (Parisi, 1981) is the discretization of CLD without the accept-reject step and converges to a slightly different distribution.
+Stochastic Gradient Langevin Dynamics (SGLD) (Gelfand and Mitter, 1991; Welling and Teh, 2011) is an accelerated version of ULA replacing the true gradient of $V$ by an unbiased estimate realized with minibatches.
+All these processes are of theoretical interest as models for Stochastic Gradient Descent (SGD), but at least SGLD is also used as a learning algorithm in practice.
+In Appendix C.1, we give more recent references containing convergence guarantees and discuss CLD and ULA in some detail.
+In the next section we show how the PAC-Bayesian bound and our integral representation can be applied to time-homogeneous Markov processes.
+Throughout this section, we assume $\mathcal{H}=\mathbb{R}^{d}$ and an isotropic Gaussian prior $\pi$ of width $\sigma$.
+We condition on the training data $\mathbf{x}$, reference to which we often omit.
+4.1 Bounds for Markov processes
+In this section, we both apply and improve the method in the recent work of Harel et al.
+(2025) on Markov processes.
+We take $\left\{h_{t}\left(\mathbf{x}\right)\right\}_{t\in I}$ to be a time-homogeneous Markov-process in real or discrete time, $I=\left[0,\infty\right)$ or $I=\mathbb{N}_{0}$, with values in $\mathcal{H}$.
+This is our model of a training process, such as CLD or the practically implementable MALA, ULA or SGLD.
+The distribution of $h_{t}\left(\mathbf{x}\right)$ will be denoted $\nu_{t}\left(\mathbf{x}\right)$.
+We will use the following result, sometimes referred to as the second law of thermodynamics.
+A proof is given in Section C.3.
+Lemma 4.1.
+(Second law of thermodynamics) If $\nu$ is a stationary distribution of $\left\{h_{t}\right\}_{t\in I}$ and $s<t$ then $KL\left(\nu_{t},\nu\right)\leq KL\left(\nu_{s},\nu\right)$ and $R_{\infty}\left(\nu_{t},\nu\right)\leq R_{\infty}\left(\nu_{s},\nu\right)$, with equality in either case if and only if $\nu_{s}=\nu$.
+Now for any $t\in I$ and any $\beta>0$
+$KL\left(\nu_{t},\pi\right)$ $=$ $\mathbb{E}_{h\sim\nu_{t}}\left[\ln\frac{d\nu_{t}}{dG_{\beta}}\right]+\mathbb{E}_{h\sim\nu_{t}}\left[\ln\frac{dG_{\beta}}{d\pi}\right]$ (6)
+$=$ $KL\left(\nu_{t},G_{\beta}\right)-\beta~\mathbb{E}_{h\sim\nu_{t}}\left[\hat{L}\left(h\right)\right]-\ln Z_{\beta}.$
+There is an analogous identity for the single draw, omitting the expectations in $\nu_{t}$ and replacing $KL$ by $R_{\infty}$.
+Now assume that $G_{\beta}$ is a stationary distribution of the process, and that $\nu_{0}=\pi$, so the process is started from the prior.
+Then, Harel et al.
+(2025) use Lemma 4.1 above, to get $KL\left(\nu_{t},G_{\beta}\right)\leq KL\left(\pi,G_{\beta}\right)=\beta\ \mathbb{E}_{h\sim\pi}\left[\hat{L}\left(h\right)\right]+\ln Z_{\beta}$, and they substitute this bound in the above identity.
+Since the partition functions cancel each other, and since $\beta\mathbb{E}_{h\sim\nu_{t}}\left[\hat{L}\left(h\right)\right]\geq 0$, they end up with $KL\left(\nu_{t},\pi\right)\leq\beta\ \mathbb{E}_{h\sim\pi}\left[\hat{L}\left(h\right)\right]$, to be substituted in the PAC-Bayesian bound.
+There is a similar bound for the single draw in terms of the relative Rényi entropy $R_{\infty}\left(\nu_{\beta,t},G_{\beta}\right)$.
+The resulting generalization bounds of Harel et al.
+(2025) remain valid along the entire training trajectory.
+They are, however, largely independent of distribution and data, and vacuous for $\beta>n$, since typically $\mathbb{E}_{h\sim\pi}\left[\hat{L}\left(h\right)\right]$and $\left\|\hat{L}\right\|_{\infty}$are on the order of unity or larger (here, we omitted several substantial refinements in Harel et al.
+(2025)).
+The use of the second law is elegant, but in this form it forgoes the potential benefit of a process converging to $G_{\beta}$ (as CLD, MALA, or HMCMC) or to some nearby distribution (as with ULA or SGLD).
+The following proposition takes advantage of convergence as well as of close approximation.
+Proposition 4.2.
+Assume $\nu_{0}=\pi$. (i) Let $\lambda_{t}=R_{\infty}\left(\nu_{t},G_{\beta}\right)/R_{\infty}\left(\pi,G_{\beta}\right)$. Then
+$R_{\infty}\left(\nu_{t},\pi\right)\leq\lambda_{t}\beta~\left\|\hat{L}\right\|_{\infty}+\left(1-\lambda_{t}\right)\int_{0}^{\beta}\mathbb{E}_{g\sim G_{\gamma}}\left[\hat{L}\left(g\right)\right]d\gamma.$
+(ii) If instead $\lambda_{t}=KL\left(\nu_{t},G_{\beta}\right)/KL\left(\pi,G_{\beta}\right)$, then
+$KL\left(\nu_{t},\pi\right)\leq\lambda_{t}\beta~\mathbb{E}_{g\sim\pi}\left[\hat{L}\left(g\right)\right]+\left(1-\lambda_{t}\right)\int_{0}^{\beta}\mathbb{E}_{g\sim G_{\gamma}}\left[\hat{L}\left(g\right)\right]d\gamma.$
+Proof.
+We only prove (i), the proof of (ii) being analogous.
+$R_{\infty}\left(\nu_{t},\pi\right)$ $=$ $\sup_{h\in\mathcal{H}}\left(\ln\frac{d\nu_{t}}{dG_{\beta}}\left(h\right)+\ln\frac{dG_{\beta}}{d\pi}\left(h\right)\right)$
+$\leq$ $\lambda_{t}R_{\infty}\left(\pi,G_{\beta}\right)+\sup_{h\in\mathcal{H}}\left(-\beta\hat{L}\left(h\right)-\ln Z_{\beta}\right)$
+$\leq$ $\lambda_{t}\sup_{h\in\mathcal{H}}\left(\beta\hat{L}\left(h\right)+\ln Z_{\beta}\right)-\ln Z_{\beta}$
+$=$ $\lambda_{t}\beta~\left\|\hat{L}\right\|_{\infty}+\left(1-\lambda_{t}\right)\int_{0}^{\beta}\mathbb{E}_{h\sim G_{\gamma}}\left[\hat{L}\left(h\right)\right]d\gamma,$
+where we used Lemma 3.1 in the last step. ∎
+Remarks: 1.
+Without any additional assumptions these bounds remain valid for all $\beta$ and along the entire training trajectory.
+For all $t$, where $\lambda_{t}\leq 1$, the bounding expression is a convex combination.
+By the last assertion of Lemma 3.1 the integral is generically smaller than $\beta~\left\|\hat{L}\right\|_{\infty}$ or $\beta~\mathbb{E}_{h\sim\pi}\left[\hat{L}\left(h\right)\right]$ respectively.
+So whenever $\lambda_{t}<1$ our bounds are also smaller than those in Section 3 of Harel et al.
+(2025).
+For ULA and SGLD we expect that $\lambda_{t}$ converges to some value $\lambda_{\infty}\ll 1$.
+2.
+If $G_{\beta}$ is indeed stationary, then by the second law (Lemma 4.1) $\lambda_{t}$ is non-increasing in $t$ and $\lambda_{t}<1$ for $t>0$ in all non-trivial cases.
+The bounds then move in convex interpolation towards the integral.
+3.
+If the process converges to $G_{\beta}$ in relative entropy, meaning that $KL\left(\nu_{t},G_{\beta}\right)\rightarrow 0$, then $\lambda_{t}\rightarrow 0$ as $t\rightarrow\infty$.
+In this case the bounds asymptotically approach those of the Gibbs posterior in Lemma 3.1, with the modification that the negative terms are omitted.
+For large $t$ they exhibit the same sensitivity of generalization to mean training errors of the Gibbs posterior at higher temperatures.
+This is the case for CLD and all algorithms containing a Metropolis-style accept-reject mechanism, such as MALA or HMCMC.
+To illustrate this, assume that $\mathcal{H}=\mathbb{R}^{d}$ with Gaussian prior $\pi$ of width $\sigma$.
+When adapting the stochastic differential equation for CLD to temperature and prior, it becomes
+$dh_{\beta,t}=-\left(\nabla\hat{L}\left(h_{\beta,t}\right)+\frac{h_{\beta,t}}{\beta\sigma^{2}}\right)dt+\sqrt{\frac{2}{\beta}}dB_{t},$
+where $B_{t}$ is standard centered Brownian motion in $\mathbb{R}^{d}$.
+Let $\nu_{\beta,t}$ be the distribution of $h_{\beta,t}$ at time $t$.
+In Lemma C.3 in Appendix C.2 we show, that, if $G_{\beta}$ satisfies a logarithmic Sobolev inequality with constant $\alpha$ (see Section C.1), then $KL\left(\nu_{\beta,t},G_{\beta}\right)\leq e^{-2\alpha t/\beta}KL\left(\nu_{\beta,0},G_{\beta}\right)$.
+So if CLD starts from the prior, we can use Proposition 4.2 (ii) with $\lambda_{t}=e^{-2\alpha t/\beta}$ and obtain
+$KL\left(\nu_{\beta,t},\pi\right)\leq e^{-2\alpha t/\beta}~\beta~\mathbb{E}_{h\sim\pi}\left[\hat{L}\left(h\right)\right]+\left(1-e^{-2\alpha t/\beta}\right)\int_{0}^{\beta}\mathbb{E}_{G_{\gamma}}\left[\hat{L}\left(g\right)\right]d\gamma.$
+4.2 Stability of the Bounds
+If $\nu:\mathcal{X}^{n}\to[0,\infty)$ is the stochastic algorithm for which we want the bound, then the results of the previous section show, that with sufficient approximation of $\nu\left(\mathbf{x}\right)$ by $G_{\beta}\left(\mathbf{x}\right)$ in relative entropy we can largely recover the bounds for the Gibbs posterior.
+But these bounds, though data-dependent, are still inaccessible because of the continuous nature of the temperature integral and the impossibility to sample directly from the Gibbs posterior.
+We now study the following question of stability: given some algorithm to approximate $G_{\beta}$ for any $\beta$ to arbitrary precision in relative entropy, can we also approximate our bounds to arbitrary precision?
+To this end we discretize the temperature scale of the integral and approximate each $G_{\beta_{k}}$ by some distribution $\nu_{k}$.
+The error incurred on the corresponding expectations of $\hat{L}$, can then be controlled under either boundedness or Lipschitz conditions on the loss $\ell$.
+Definition 4.3.
+For $\mathbf{x\in}\mathcal{X}^{n}$, an increasing sequence $\beta_{0}^{K}=\left(0=\beta_{0}<\beta_{1}<\cdots<\beta_{K}=\beta\right)$ of positive numbers, and a corresponding vector of data-dependent distributions $\nu_{0}^{K-1}(\mathbf{x})=\left(\nu_{0}\left(\mathbf{x}\right),\nu_{1}\left(\mathbf{x}\right),\cdots,\nu_{K-1}\left(\mathbf{x}\right)\right)\in\mathcal{P}\left(\mathcal{H}\right)^{K}$ we denote
+$\Gamma(\nu_{0}^{K-1},\mathbf{x},\beta_{0}^{K})=\sum_{k=1}^{K}(\beta_{k}-\beta_{k-1})\mathbb{E}_{g\sim\nu_{k-1}(\mathbf{x})}\big[\hat{L}(g,\mathbf{x})\big].$
+The next lemma bounds the estimation error relative to the temperature integral in terms of the relative entropies.
+Lemma 4.4.
+With $\beta_{0}^{K}$ and $\nu_{0}^{K-1}$ as in Definition 4.3 denote
+$\Delta:=\int_{0}^{\beta}\mathbb{E}_{h\sim G_{\gamma}\left(\mathbf{x}\right)}\left[\hat{L}\left(h,\mathbf{x}\right)\right]d\gamma-\Gamma(\nu_{0}^{K-1},\mathbf{x},\beta_{0}^{K})$
+(i) If $\mathbb{E}_{h\sim G_{\beta_{k}}\left(\mathbf{x}\right)}\left[\hat{L}\left(h,\mathbf{x}\right)\right]\leq\mathbb{E}_{h\sim\nu_{k}\left(\mathbf{x}\right)}\left[\hat{L}\left(h,\mathbf{x}\right)\right]$ for all $k$ and $\mathbf{x}$, then $\Delta\leq 0$.
+(ii) If $\ell\left(h,\mathbf{x}\right)$ is bounded in $h$ for all $\mathbf{x}$, $\left\|\ell\left(h,\mathbf{x}\right)\right\|\leq m$ then
+$\Delta\leq m\sum_{k=1}^{K}(\beta_{k}-\beta_{k-1})\sqrt{KL\left(\nu_{k-1}\left(\mathbf{x}\right),G_{\beta_{k-1}}\left(\mathbf{x}\right)\right)/2}.$
+(iii) If instead $\ell\left(h,\mathbf{x}\right)$ is $m$-Lipschitz in $h$ for all $\mathbf{x}$, $\ell\left(h,\mathbf{x}\right)-\ell\left(g,\mathbf{x}\right)\leq m\left\|h-g\right\|$ and $G_{\beta_{k}}\left(\mathbf{x}\right)$ satisfies an LSI with constant $\alpha$ for all $k$ and $\mathbf{x}$, then
+$\Delta\leq\frac{2m}{\alpha}\sum_{k=1}^{K}(\beta_{k}-\beta_{k-1}){KL\left(\nu_{k-1}\left(\mathbf{x}\right),G_{\beta_{k-1}}\left(\mathbf{x}\right)\right)}.$
+By the last conclusion of Lemma 3.1 part (i) is immediate.
+Proofs of (ii) and (iii) are given in Appendix C.4.
+The assumption in case (i) is not implausible, if we start LMC from a non-informative prior, and in our experiments we always observed decreasing losses along the LMC path.
+The next theorem gives our final bound in terms of arbitrary distributions and their relative entropies to Gibbs distributions.
+Theorem 4.5.
+Let $F:\mathcal{H}\times\mathcal{X}^{n}\rightarrow\mathbb{R}$ be some measurable function and $\beta_{0}^{K}$ and $\nu_{0}^{K-1}$ as in Definition 4.3.
+Let $\nu(\mathbf{x})$ be any data-dependent distribution on $\mathcal{H}$.
+Let $\Delta$ be bounded as in Lemma 4.4, depending on which of the conditions there is fulfilled by $\ell$.
+Then
+(i) with probability at least $1-\delta$ as $\mathbf{x}\sim\mu^{n}$ and $h\sim\nu(\mathbf{x})$
+$F\left(h,\mathbf{x}\right)$ $\leq$ $-\beta\hat{L}\left(h,\mathbf{x}\right)+\Gamma(\nu_{0}^{K-1},\mathbf{x},\beta_{0}^{K})+\ln\mathbb{E}_{\mathbf{x}}\mathbb{E}_{h\sim\pi}\left[e^{F\left(h,\mathbf{x}\right)}\right]+\ln\frac{1}{\delta}$
+$+R_{\infty}\left(\nu\left(\mathbf{x}\right),G_{\beta}\left(\mathbf{x}\right)\right)+\Delta.$
+If $F$ and $\ell$ are bounded, then $R_{\infty}\left(\nu\left(\mathbf{x}\right),G_{\beta}\left(\mathbf{x}\right)\right)$ can be replaced by
+$\max\left\{0,\beta\left\|\ell\right\|_{\infty}+\left\|F\right\|_{\infty}+\ln\sqrt{2KL\left(\nu\left(\mathbf{x}\right),G_{\beta}\left(\mathbf{x}\right)\right)}\right\}.$
+(ii) with probability at least $1-\delta$ as $\mathbf{x}\sim\mu^{n}$
+$\mathbb{E}_{\nu\left(\mathbf{x}\right)}\left[F(h,\mathbf{x})\right]$ $\leq$ $-\beta\ \mathbb{E}_{\nu\left(\mathbf{x}\right)}\big[\hat{L}\left(h,\mathbf{x}\right)\big]+\Gamma(\nu_{0}^{K-1},\mathbf{x},\beta_{0}^{K})+\ln\mathbb{E}_{\mathbf{x}}\mathbb{E}_{h\sim\pi}\left[e^{F\left(h,\mathbf{x}\right)}\right]+\ln\frac{1}{\delta}$
+$+KL\left(\nu\left(\mathbf{x}\right),G_{\beta}\left(\mathbf{x}\right)\right)+\Delta.$
+The left-hand side in both inequalities is the random variable that we wish to bound, depending on the distribution $\nu$.
+The right-hand side of the top line is the bound, as computed from the distributions $\nu(\mathbf{x})$ and $\nu_{0}^{K-1}(\mathbf{x})$ and includes the dependence on the exponential moment of $F$ and the confidence parameter $\delta$.
+The second line gives the error incurred by the fact that none of the distributions is really the right Gibbs distribution.
+The first term there gives the error for the target distribution $\nu(\mathbf{x})$, and is different in the single-draw and classical PAC-Bayesian cases.
+The term $\Delta$ results from approximating the temperature integral by the expectations in a finite number of distributions, as described in Lemma 4.4.
+The amendment to (i) is necessary, since we know of no process with useful bounds on $R_{\infty}$.
+The replacement indeed converges to zero with relative entropy, but, since $\beta\left\|\ell\right\|_{\infty}+\left\|F\right\|_{\infty}$ is typically of order $n$, it requires the relative entropy to be exponentially small in $n$.
+Nevertheless, if the bounds in Corollary C.2 are substituted in (ii) or in the amendment to (i), they guarantee, with sufficient computational budget and appropriate choices of $t$ and $\eta$, the convergence of LMC to our bounds for the Gibbs posterior.
+A detailed proof of Theorem 4.5 is given in Appendix C.4.
+Part (ii) and (i) without amendment follow more or less directly from the PAC-Bayesian theorem, Lemma 4.4 and equation (6) and a reasoning using $R_{\infty}$ analogous to (6).
+The amendment to (i) requires a special adapted proof of the PAC-Bayesian theorem.
+5 Experiments
+The purpose of our experiments is twofold.
+For one, we want to verify that the theoretical dependence of the generalization performance of the Gibbs posterior at low temperatures on its training errors at high temperatures carries over to practical temperature-regularized algorithms like SGLD in real world settings, with overparametrized neural networks.
+Indeed, in all our experiments the temperature plots of the mean training errors, computed as described below, verify the qualitative prediction that both the failure of generalization for random labels and the success for true labels are related to the areas under the curves at higher temperatures (see for example Figure 2).
+Second, we would like to use the training data, and only the training data, to make realistic quantitative predictions of test errors in such settings.
+This is more difficult, since the guarantees of Corollary C.2 in combination with Theorem 4.5 are inadequate in high-dimensional situations.
+We achieve this goal with a principled calibration scheme described below.
+5.1 Experimental Environment and Algorithms
+The real-world data are either the MNIST dataset, subdivided into the two classes of characters 0–4 and 5–9, or the CIFAR-10 dataset to distinguish between animals and vehicles.
+For impossible data, we randomize the labels of the training data.
+Our experiments are computationally heavy, so we generally use small sample sizes, from 2000 to 8000 examples.
+The hypothesis space is the set of weight vectors for a neural network with ReLU activation functions constrained by a Gaussian prior distribution with $\sigma=5$.
+Neural network architectures are described in Section D.1.1 of the appendix.
+To approximately sample the weight vectors in the vicinity of the Gibbs posterior, we use ULA as in (15) or SGLD (Welling and Teh, 2011) with constant step size $\eta$.
+5.2 The Loss Function $\ell$
+Most experiments were done with bounded loss functions $\ell$, either bounded binary cross-entropy as described in Appendix D of Dziugaite and Roy (2018) or the Savage loss (Masnadi-Shirazi and Vasconcelos, 2008).
+As unbounded loss function we tried binary cross-entropy (BCE) (Section D.2.6), but with a smaller value of $\sigma$, so as to avoid excessive training errors for small values of $\beta$.
+We compute bounds for the 0-1 loss, using the method described in Section B.2.
+5.3 Approximating the Ergodic Mean
+As we know of no sufficient criterion for convergence, we terminate iterations at time $T$, when a very slow running mean $\mathbb{M}_{\rm stop}$ of the loss trajectory $\big(\hat{L}(h_{\beta_{k},t},\mathbf{x})\big)_{t=0}^{T}$ stops decreasing.
+A second running mean $\mathbb{M}_{\rm erg}$ is used as an approximation of the ergodic mean and thus of expectations in the invariant distribution.
+We thus replace all expectations $\mathbb{E}_{h\sim G_{\beta_{k}}}\big[\hat{L}(h,\mathbf{x})\big]$ occurring in the bounds by $\mathbb{M}_{\rm erg}\big[\big(\hat{L}(h_{\beta_{k},t},\mathbf{x})\big)_{t=0}^{T}\big]$.
+Both running means $\mathbb{M}_{\rm stop}$ and $\mathbb{M}_{\rm erg}$ are implemented as first-order, recursive lowpass filters described in Section D.1.3 of the appendix.
+5.4 Computation of the Bounds and Calibration
+For the 01-error (cf. Section B.2) we compute our bounds from the PAC-Bayesian theorem in the form
+$\mathbb{E}_{h\sim\nu_{\beta}\left(\mathbf{x}\right)}\left[L_{\text{0}1}\left(h\right)\right]\leq\kappa^{-1}\left(\mathbb{E}_{h\sim\nu_{\beta}\left(\mathbf{x}\right)}\left[\hat{L}_{\text{0}1}\left(h,\mathbf{x}\right)\right],\frac{\mathcal{Q}+\ln\left(\frac{2\sqrt{n}}{\delta}\right)}{n}\right),$ (7)
+where $\nu_{\beta}\left(\mathbf{x}\right)$ is an LMC approximation to $G_{\beta}\left(\mathbf{x}\right)$ and $\mathcal{Q}$ is a proxy for $KL\left(\nu_{\beta}\left(\mathbf{x}\right),\pi\right)$, the computation of which is described below.
+Using $\kappa^{-1}$ is somewhat more accurate than the analog of (2).
+The function $\kappa$ is the relative entropy of two Bernoulli variables, the derivation of the above bound and the inverse function $\kappa^{-1}$ are explained in section B.2.
+Our approximation to $KL\left(\nu_{\beta}\left(\mathbf{x}\right),\pi\right)$ is
+$A=-\beta\ \mathbb{E}_{h\sim\nu_{\beta}}\left[\hat{L}\left(h,\mathbf{x}\right)\right]+\Gamma(\nu_{0}^{K-1},\mathbf{x},\beta_{0}^{K}).$
+For a rigorous bound following Theorem 4.5 we would have to set $\mathcal{Q}=A$ plus all the terms bounding the approximation errors as in Corollary C.2.
+Unfortunately, the quantities $R$ and $\alpha$ are impossible to estimate in practice.
+But even if we assume these to be in the order of unity, the bounds are too coarse to distinguish between different temperatures with realistic step sizes.
+A simple calculation (see Lemma C.7 in Appendix C.5) shows that
+$\text{KL}\left(G_{\beta},G_{2\beta}\right)\leq\beta\left(\mathbb{E}_{G_{\beta}}\left[\hat{L}\right]-\mathbb{E}_{G_{2\beta}}\left[\hat{L}\right]\right)\leq\beta\mathbb{E}_{G_{\beta}}\left[\hat{L}\right$].
+By Corollary C.2 we should therefore have at least $8\eta dR^{2}/\alpha<\mathbb{E}_{G_{\beta}\left(\mathbf{x}\right)}\left[\hat{L}\right$] to distinguish between the expectations in the Gibbs posterior for $\beta$ and $2\beta$.
+The smallest neural network we use has $d=392,500$.
+If $\ell$ has values in $\left[0,1\right]$ then $\mathbb{E}_{G_{\beta}\left(\mathbf{x}\right)}\left[\hat{L}\right]\leq 1$, so even if $R$ and $\alpha$ are set to $1$, we would need step sizes in the order of $10^{-7}$.
+Safe values of $\eta$, as suggested by the theoretical results in Section 4, are therefore impossible in practice, and the bound has to be adapted to a realistic choice of $\eta$.
+But the data for the true labels $\mathbf{x}$ and the data $\mathbf{\bar{x}}$ for the random labels are strongly related.
+Dimension and input marginals are equal.
+This suggests that the inaccuracy of the LMC approximations affects both in a similar way, and we make the following calibration assumption:
+$\frac{KL\left(\nu_{\beta}\left(\mathbf{x}\right),\pi\right)}{KL\left(\nu_{\beta}\left(\mathbf{\bar{x}}\right),\pi\right)}=\frac{A}{\bar{A}},$
+where $\bar{A}$ is given by the expression analogous to $A$ for the random labels, compare Figure 2.
+The expected error of the random labels in binary classification is $1/2.$ Now let $r$ be the smallest positive factor such that $r\bar{A}$, when substituted for $\mathcal{Q}$ in (7) yields a value greater or equal $1/2$.
+If the PAC-Bayesian theorem is tight, then this means that $r\bar{A}\geq KL\left(\nu_{\beta}\left(\mathbf{\bar{x}}\right),\pi\right)$.
+Our calibration assumption above then also implies that
+$rA\geq KL\left(\nu_{\beta}\left(\mathbf{x}\right),\pi\right),$
+and substitution of $rA$ for $\mathcal{Q}$ in (7) should result in a correct upper bound for the true labels. For a precise definition of $r$ see Section C.6.
+Clearly there is no rigorous justification of this argument.
+It is an experimental finding that our choice of $r$ leads to correct and surprisingly tight upper bounds on the test error of correctly labeled data in all cases we tried.
+We emphasize that our calibration procedure depends only on the training data.
+5.5 Results
+Figure 4: SGLD on MNIST and CIFAR-10 with 8000 training examples, MNIST above and CIFAR-10 below, random labels on the left, correct labels on the right.
+Both random and true labels are trained with the same algorithm and parameters on a fully connected ReLU network with two hidden layers of 1000 and 1500 units, respectively.
+The calibration factor for MNIST is 0.77, for CIFAR-10 0.89.
+Train error, test error and our bound for the Gibbs posterior average of the 0-1 loss are plotted against $\beta$.
+Several experiments confirm the validity of the proposed bounds.
+An example is shown in Figure 4, where a fully connected ReLU network with two hidden layers of 1000 (respectively 1500) units each is trained with SGLD at inverse temperatures $\beta=$ 0, 500, 1000, 2000, 4000, 8000, 16000, 32000, and 64000.
+The train error for random labels is about 0.1 (or even less) at $\beta=64000$, where the bound is above 0.5.
+The test error for correct labels, however, is tightly bounded above.
+Notice that for MNIST, which has the tightest bounds, the training error for the true labels is rapidly decreasing from 0.5 to 0.17 at $\beta=500$ and to 0.1 at $\beta=1000$.
+The more moderate initial decrease for CIFAR-10 corresponds to the tendency to overfit on this more difficult dataset.
+This confirms the intuition, that good generalization at low temperatures is already announced in the high-temperature regime.
+Experimental bounds for single posterior draws, along with additional experiments including applications to model selection, are presented in Section D.2.
+6 Conclusion
+Using the integral representation of the log-partition function, the Gibbs posterior admits the computation of upper bounds on the true error based on the training data and for any temperature.
+These bounds are stable under perturbation in relative entropy and can be approximated by Langevin Monte Carlo (LMC) algorithms.
+However, for realistic experiments, the approximations obtained by these algorithms are coarse and require calibration, which leads to rather tight bounds in the interpolation regime of overparametrized neural networks.
+The fact that the calibrated bounds are very tight is, at this point, a purely experimental finding, requiring more theoretical investigation in future work.
+Appendix
+In this appendix, we provide a glossary of notation, give additional theoretical results and missing proofs, and provide more information on the numerical experiments, as well as additional experimental results.
+Appendix A Table of notation
+Notation Brief description Section
+$\mathcal{X}$ space of data 2, 3, 4
+$\Sigma$ sigma algebra (events) on $\mathcal{X}$ 2
+$\mu$ probability of data 2, 4.2
+$n$ sample size 1, 2, 3, 4
+$\delta$ confidence parameter in high probability bounds 1, 2, 4.2
+$\mathbf{x}$ generic member $\left(x_{1},...,x_{n}\right)\in\mathcal{X}^{n}$, training sample 1, 2, 3, 4
+$\mathcal{H}$ hypothesis space 2, 3, 4.2
+$\Omega$ sigma algebra (events) on $\mathcal{H}$ 2, 3
+$\ell$ $\ell:\mathcal{H\times X\rightarrow}\left[0,\infty\right)$ loss function 1, 2, 4.2, 5
+$\mathcal{P}\left(\mathcal{H}\right)$ probability measures on $\mathcal{H}$ 2, 3
+$\pi$ nonnegative a-priori measure on $\mathcal{H}$ 2, 3, 4
+$\|f\|_{\infty}$ The sup-norm is defined: $\|f\|_{\infty}=\sup\{|f(s)|:s\in S\}$ 2, 4
+$\sigma$ width of Gaussian prior 4.1, 5
+$L\left(h\right)$ $L\left(h\right)=\mathbb{E}_{x\sim\mu}\left[\ell(h,x)\right]$, expected loss of $h\in\mathcal{H}$ 2
+$\hat{L}\left(h,\mathbf{x}\right)$ $\hat{L}\left(h,\mathbf{x}\right)=\left(1/n\right)\sum_{i=1}^{n}\ell\left(h,x_{i}\right)$, empirical loss of $h\in\mathcal{H}$ 2, 3, 4, 5
+$\beta$ inverse temperature 1, 2, 3, 4, 5
+$Z_{\beta}\left(\mathbf{x}\right)$ partition function 3, 4.1
+$G_{\beta,\pi}\left(\mathbf{x}\right)$ Gibbs posterior with energy $\hat{L}$ and prior $\pi$ 1, 3, 4, 5
+$\mathbb{E}_{g\sim G_{\beta}\left(\mathbf{x}\right)}$ posterior expectation 3
+$\beta_{0}^{K}$ increasing sequence $\left(0=\beta_{0}<\beta_{1}<\cdots<\beta_{K}=\beta\right)$ of positive reals 4.2, 5
+$\Gamma(\nu_{0}^{K-1},\mathbf{x},\beta_{0}^{K})$ bounding functional 4.2, 5
+$F\left(h,\mathbf{x}\right)$ placeholder for generalization gap 2, 4.2
+$\kappa(p,q)$ $kl\left(p,q\right)=p\ln\frac{p}{q}+\left(1-p\right)\ln\frac{1-p}{1-q}$, rel. entropy of Bernoulli variables 5
+$\kappa^{-1}(p,t)$ $\inf\left\{q:q\geq p,\kappa\left(p,q\right)\geq t\right\}$ 5.4
+$\text{KL}\left(\rho,\nu\right)$ $\int\left(\ln\frac{d\rho}{d\nu}\right)d\rho$, KL-divergence of $\rho,\nu\in\mathcal{P}\left(\mathcal{H}\right)$ 2, 3, 4, 5
+$R_{\infty}\left(\rho,\nu\right)$ $\sup_{h\in\mathcal{H}}\ln\frac{d\rho}{d\nu}(h)$, Rényi-infinity divergence of $\rho,\nu\in\mathcal{P}\left(\mathcal{H}\right)$ 2, 4.2
+$d_{TV}\left(\rho,\nu\right)$ total variation distance C.4
+$W_{p}\left(\rho,\nu\right)$ $p$-Wasserstein distance C.4
+$\eta$ step size or learning rate C.1
+$\nu_{\beta,\eta}$ invariant measure of LMC approximation of $G_{\beta}$ with step size $\eta$ C.1
+$\nu_{\beta,\eta,t}$ LMC approximation of $G_{\beta}$ with step size $\eta$ at iteration $t$ C.1
+$\nu_{\beta,\epsilon}$ invariant measure of CLD process with step size $\epsilon$ C.1
+$\nu_{\beta,\epsilon,t}$ CLD with step size $\epsilon$ at iteration $t$ C.1
+$r$ calibration factor 5.4
+$\mathbf{\bar{x}}$ randomly labeled data 5.4
+$\mathbb{M}_{\rm stop}$$\mathbb{M}_{\rm erg}$ filters for stopping and ergodic mean 5.3
+$L_{01}\left(h\right)$ $L_{01}\left(h\right)=\mathbb{E}_{x\sim\mu}\left[\ell_{01}(h,x)\right]$, expected 01-loss of $h\in\mathcal{H}$ 5.4
+$\hat{L}_{01}\left(h,\mathbf{x}\right)$ $\hat{L}\left(h,\mathbf{x}\right)=\left(1/n\right)\sum_{i=1}^{n}\ell_{01}\left(h,x_{i}\right)$, empirical 01-loss of $h\in\mathcal{H}$ 5.4
+Appendix B PAC-Bayes
+We review PAC-Bayesian theory. There is no claim to novelty.
+B.1 Proof of the PAC-Bayesian Theorem
+Definition B.1.
 Given a stochastic algorithm $\nu$ we define a probability measure $\rho_{\nu}$ on $\mathcal{H}\times\mathcal{X}^{n}$ by
@@
 To draw the pair $\left(h,\mathbf{x}\right)$ from $\rho_{\nu}$ we first draw the training sample $\mathbf{x}$, and then sample $h$ from $\nu\left(\mathbf{x}\right)$.
-The main objective in learning is that the risk $\mathbb{E}_{x\sim\mu}\left[f\left(h,x\right)\right]$ is small with high probability in $\left(h,\mathbf{x}\right)\sim\rho_{\nu}$, where $f$ is some application-dependent loss function, possibly different from $\ell$.
-In the sequel we will give corresponding guarantees.
-3 Bounds for the Gibbs posterior
-In this section, we make the idealized assumption that we are free to sample from the Gibbs posterior at any finite $\beta\geq 0$.
-3.1 An integral representation of the free energy
-Lemma 3.1.
-Let $0=\beta_{0}<\beta_{1}<\cdots<\beta_{K}=\beta$. Then
-$\ln Z_{\beta}(\mathbf{x})=\int_{0}^{\beta}\mathbb{E}_{h\sim G_{\gamma}(\mathbf{x})}\big[\hat{L}(h,\mathbf{x})\big]d\gamma\leq\sum_{k=1}^{K}(\beta_{k}-\beta_{k-1})\mathbb{E}_{g\sim G_{\beta_{k-1}}(\mathbf{x})}\big[\hat{L}(g,\mathbf{x})\big].$
-Proof.
-Let $A(\beta)=-\ln Z_{\beta}(\mathbf{x})$. One verifies the identities
-$A(0)$ $=$ $0,$
-$A^{\prime}(\beta)$ $=$ $\frac{1}{Z_{\beta,\pi}(\mathbf{x})}\int_{\mathcal{H}}\hat{L}(h,\mathbf{x})e^{-\beta\hat{L}(h,\mathbf{x})}d\pi(h)=\mathbb{E}_{h\sim G_{\beta}(\mathbf{x}}\big[\hat{L}(h,\mathbf{x})\big],$
-$A^{\prime\prime}(\beta)$ $=$ $-\Big(\mathbb{E}_{h\sim G_{\beta}(\mathbf{x})}\big[\hat{L}(h,\mathbf{x})^{2}\big]-\big(\mathbb{E}_{h\sim G_{\beta}(\mathbf{x})}\big[\hat{L}(h,\mathbf{x})\big]\big)^{2}\Big)\leq 0\text{.}$
-The equality in the lemma then follows from the first two identities above and the fundamental theorem of calculus, and the inequality follows from the last identity, which shows that $\mathbb{E}_{g\sim G_{\beta_{k-1},\pi}(\mathbf{x})}\big[\hat{L}(g,\mathbf{x})\big]$ is non-increasing in $\beta$.
-∎
-In statistical physics there is a formal analogy, where the function $h\mapsto\hat{L}(h,\mathbf{x})$ is the ($\mathbf{x}$-dependent) energy of the system in the state $h$, and $\beta$ is the inverse temperature.
-The Gibbs posterior then becomes the "canonical ensemble" (Gibbs, 1902), describing the probability of states in equilibrium with a heat bath at temperature $\beta^{-1}$.
-The function $\beta\mapsto A(\beta)$ plays an important role: $\beta^{-1}A(\beta)$ is the Helmholz free energy, $A^{\prime}(\beta)=\mathbb{E}_{h\sim G_{\beta}(\mathbf{x})}\big[\hat{L}(h,\mathbf{x})\big]$ is the thermal average of the energy, $\beta A^{\prime}(\beta)+A(\beta)$ is the entropy and $A^{\prime\prime}(\beta)$ is proportional to the heat capacity at temperature $\beta^{-1}$ (see e.g.
-Huang, 2008).
-For $h\in\mathcal{H}$, $\mathbf{x\in}\mathcal{X}^{n}$ and an increasing sequence $\bm{\beta}=\left(\beta_{1}<\cdots<\beta_{K}\right)$ of positive numbers, we denote
-$\Gamma(h,\mathbf{x},\bm{\beta})=-\beta_{K}\hat{L}(h,\mathbf{x})+\sum_{k=1}^{K}(\beta_{k}-\beta_{k-1})\mathbb{E}_{g\sim G_{\beta_{k-1}}(\mathbf{x})}\big[\hat{L}(g,\mathbf{x})\big].$ (3)
-So Lemma 3.1 states that, when $\beta_{0}=0$, then,
-$\beta_{K}\hat{L}\left(h,\mathbf{x}\right)-\ln Z_{\beta_{K}}\left(\mathbf{x}\right)\leq\Gamma\left(h,\mathbf{x},\bm{\beta}\right).$ (4)
-Note that $\Gamma\left(h,\mathbf{x},\bm{\beta}\right)$ depends only on the training data $\mathbf{x}$, the sequence $\bm{\beta}$ and the hypothesis $h$.
-3.2 Bounds
-The function $F$ in the following is a placeholder for a random variable related to the generalization gap, which we would like to bound with high probability.
-Theorem 3.2.
-Let $F:\mathcal{H}\times\mathcal{X}^{n}\rightarrow\mathbb{R}$ be some measurable function, $\beta>0$ and $\bm{\beta}=\left(\beta_{1}<\cdots<\beta_{K}\right)$ as above with $\beta_{0}=0$ and $\beta_{K}=\beta$.
-Then,
-(i) for $\delta>0$ with probability at least $1-\delta$ in $\mathbf{x}\sim\mu^{n}$ and $h\sim G_{\beta}\left(\mathbf{x}\right)$
-$F\left(h,\mathbf{x}\right)\leq\Gamma\left(h,\mathbf{x},\bm{\beta}\right)+\ln\mathbb{E}_{\mathbf{x}}\mathbb{E}_{g\sim\pi}\left[e^{F\left(g,\mathbf{x}\right)}\right]+\ln\left(1/\delta\right),$
+Restatement of Theorem 2.1
+Theorem B.2.
+[…]
+(i) for $\delta>0$ with probability at least $1-\delta$ in $\mathbf{x}\sim\mu^{n}$ and $h\sim\nu\left(\mathbf{x}\right)$
+$F\left(h,\mathbf{x}\right)\leq\ln\frac{d\nu\left(\mathbf{x}\right)}{d\pi}\left(h\right)+\ln\mathbb{E}_{\mathbf{x}}\mathbb{E}_{g\sim\pi}\left[e^{F\left(g,\mathbf{x}\right)}\right]+\ln\left(\frac{1}{\delta}\right)$
 (ii) for $\delta>0$ with probability at least $1-\delta$ in $\mathbf{x}\sim\mu^{n}$
-$\mathbb{E}_{h\sim G_{\beta}\left(\mathbf{x}\right)}\left[F\left(h,\mathbf{x}\right)\right]\leq\mathbb{E}_{h\sim G_{\beta}\left(\mathbf{x}\right)}\left[\Gamma\left(h,\mathbf{x},\bm{\beta}\right)\right]+\ln\mathbb{E}_{\mathbf{x}}\mathbb{E}_{g\sim\pi}\left[e^{F\left(g,\mathbf{x}\right)}\right]+\ln\left(1/\delta\right).$
+$\mathbb{E}_{h\sim\nu\left(\mathbf{x}\right)}\left[F\left(h,\mathbf{x}\right)\right]\leq\text{KL}\left(\nu\left(\mathbf{x}\right),\pi\right)+\ln\frac{\mathbb{E}_{\mathbf{x}}\mathbb{E}_{g\sim\pi}\left[e^{F\left(g,\mathbf{x}\right)}\right]}{\delta}$
 Proof.
@@
 $\Pr\left\{Y>\ln\mathbb{E}\left[e^{Y}\right]+\ln\left(1/\delta\right)\right\}=\Pr\left\{e^{Y}>\mathbb{E}\left[e^{Y}\right]/\delta\right\}\leq\delta.$
-To prove (i), we apply this to the random variable $Y=F\left(h,\mathbf{x}\right)+\beta\hat{L}\left(h,\mathbf{x}\right)+\ln Z_{\beta}\left(\mathbf{x}\right)$ on the probability space $\left(\mathcal{H\times X}^{n},\Omega\otimes\Sigma^{\otimes n},\rho_{G_{\beta}}\right)$ as defined in (2).
-Together with the definition of the Gibbs posterior, this gives, with probability at least $1-\delta$ in $\left(h,\mathbf{x}\right)\sim\rho_{G_{\beta}}$ (equivalent to saying $\mathbf{x}\sim\mu^{n}$ and $h\sim G_{\beta}\left(\mathbf{x}\right)$),
-$F\left(h,\mathbf{x}\right)+\beta\hat{L}\left(h,\mathbf{x}\right)+\ln Z_{\beta}\left(\mathbf{x}\right)$
-$\leq\ln\mathbb{E}_{\mathbf{x}}\mathbb{E}_{g\sim G_{\beta}\left(\mathbf{x}\right)}\left[e^{F\left(g,\mathbf{x}\right)+\beta\hat{L}\left(g,\mathbf{x}\right)+\ln Z_{\beta}\left(\mathbf{x}\right)}\right]+\ln\left(1/\delta\right)$
-$=\ln\mathbb{E}_{\mathbf{x}}\mathbb{E}_{g\sim\pi}\left[e^{F\left(g,\mathbf{x}\right)+\beta\hat{L}\left(g,\mathbf{x}\right)+\ln Z_{\beta}\left(\mathbf{x}\right)-\beta\hat{L}\left(g,\mathbf{x}\right)-\ln Z_{\beta}\left(\mathbf{x}\right)}\right]+\ln\left(1/\delta\right)$
+To prove (i), we apply this to the random variable $Y=F\left(h,\mathbf{x}\right)-\ln\left(d\left(\nu\left(\mathbf{x}\right)/d\pi\right)\right)\left(h\right)$ on the probability space $\left(\mathcal{H\times X}^{n},\Omega\otimes\Sigma^{\otimes n},\rho_{\nu}\right)$.
+This gives, with probability at least $1-\delta$ as $\mathbf{x}\sim\mu^{n}$ and $h\sim\nu\left(\mathbf{x}\right)$),
+$F\left(h,\mathbf{x}\right)-\ln\frac{d\nu\left(\mathbf{x}\right)}{d\pi}\left(h\right)$
+$\leq\ln\mathbb{E}_{\mathbf{x}}\mathbb{E}_{g\sim\nu\left(\mathbf{x}\right)}\left[e^{F\left(g,\mathbf{x}\right)-\ln\left(d\left(\nu\left(\mathbf{x}\right)/d\pi\right)\right)\left(g\right)}\right]+\ln\left(1/\delta\right)$
+$=\ln\mathbb{E}_{\mathbf{x}}\mathbb{E}_{g\sim\pi}\left[e^{F\left(g,\mathbf{x}\right)-\ln\left(d\left(\nu\left(\mathbf{x}\right)/d\pi\right)\right)\left(g\right)+\ln\left(d\left(\nu\left(\mathbf{x}\right)/d\pi\right)\right)\left(g\right)}\right]+\ln\left(1/\delta\right)$
 $=\ln\mathbb{E}_{\mathbf{x}}\mathbb{E}_{g\sim\pi}\left[e^{F\left(g,\mathbf{x}\right)}\right]+\ln\left(1/\delta\right).$
-Subtract $\beta\hat{L}\left(h,\mathbf{x}\right)+\ln Z_{\beta}\left(\mathbf{x}\right)$ and use (4).
-For (ii) apply Markov’s inequality to $\mathbb{E}_{h\sim G_{\beta}\left(\mathbf{x}\right)}\left[F\left(h,\mathbf{x}\right)+\beta\hat{L}\left(h,\mathbf{x}\right)+\ln Z_{\beta}\left(\mathbf{x}\right)\right]$ instead.
+For (ii) apply Markov’s inequality on the probability space $\left(\mathcal{X}^{n},\Sigma^{\otimes n},\mu^{n}\right)$ to the random variable $Y=\mathbb{E}_{g\sim\nu\left(\mathbf{x}\right)}\left[F\left(g,\mathbf{x}\right)\right]-KL\left(\nu\left(\mathbf{x}\right),\pi\right)=\mathbb{E}_{g\sim\nu\left(\mathbf{x}\right)}\left[F\left(g,\mathbf{x}\right)-\ln\left(d\left(\nu\left(\mathbf{x}\right)/d\pi\right)\right)\left(g\right)\right]$ instead.
 By Jensen’s inequality
-$e^{\mathbb{E}_{h\sim G_{\beta}\left(\mathbf{x}\right)}\left[F\left(h,\mathbf{x}\right)+\beta\hat{L}\left(h,\mathbf{x}\right)+\ln Z_{\beta}\left(\mathbf{x}\right)\right]}\leq\mathbb{E}_{g\sim G_{\beta}\left(\mathbf{x}\right)}\left[e^{F\left(g,\mathbf{x}\right)+\beta\hat{L}\left(g,\mathbf{x}\right)+\ln Z_{\beta}\left(\mathbf{x}\right)}\right]$
-and proceed as above using (4). ∎
-Up to the application of (4), the above proof of (i) just gives the single-draw version of the PAC-Bayesian bound as in Rivasplata et al.
-(2020) applied to the Gibbs posterior, while (ii) is the standard PAC-Bayesian bound applied to the Gibbs posterior until (4) is invoked.
-3.3 Loss functions and secondary loss functions
-To apply Theorem 3.2, we need to control the exponential moment $\mathbb{E}_{\mathbf{x}}\mathbb{E}_{g\sim\pi}\left[e^{F\left(g,\mathbf{x}\right)}\right]$, but otherwise we have free choice of the function $F$.
-This gives the method some flexibility.
-If $\exp\left(F\right)$ has sufficient integrability properties, we can exchange the two expectations and often there is a bound on $\mathbb{E}_{\mathbf{x}}\left[e^{F\left(g,\mathbf{x}\right)}\right]$ uniform in $g$, which then carries over to $\mathbb{E}_{g\sim\pi}\mathbb{E}_{\mathbf{x}}\left[e^{F\left(g,\mathbf{x}\right)}\right]$, since $\pi$ is a probability measure.
-In this way bounds for sub-Gaussian or sub-exponential losses can be obtained, but also for U-statistics or even non-iid data, sampled from the trajectories of time-homogeneous, ergodic Markov chains.
-In Section B.1 in the appendix, we derive a bound for sub-Gaussian losses from Theorem 3.2; other examples are planned for a longer version of the paper.
-The function $F$ may be defined in terms of other, application-dependent loss functions, which are different from the loss $\ell$, which defines the Gibbs posterior and the functional $\Gamma$.
-To illustrate this point, let $f:\mathcal{H}\times\mathcal{X}\rightarrow[0,1]$ be measurable and set $F(h,\mathbf{x})=n~\kappa\left(\frac{1}{n}\sum_{i}f(h,x_{i}),\mathbb{E}_{x}\left[f(h,x)\right]\right)$, with $\kappa$ the relative entropy as in (1).
-Then, Theorem 1 of Maurer (2004) gives $\mathbb{E}_{\mathbf{x}}\left[e^{F(h,\mathbf{x})}\right]\leq 2\sqrt{n}$ for $n\geq 8$.
-Substitution in Theorem 3.2 and division by $n$ then give the following corollary.
-Corollary 3.3.
-Let $f:\mathcal{H}\times\mathcal{X}\rightarrow[0,1]$ be measurable, $\delta>0$ and $n\geq 8$. Then, with probability at least $1-\delta$ in $\mathbf{x}\sim\mu^{n}$ and $h\sim G_{\beta}(\mathbf{x})$
-$\kappa\left(\frac{1}{n}\sum_{i}f(h,x_{i}),\mathbb{E}_{x}[f(h,x)]\right)\leq\frac{1}{n}\left(\Gamma(h,\mathbf{x},\bm{\beta})+\ln\left(\frac{2\sqrt{n}}{\delta}\right)\right)$
-and with probability at least $1-\delta$ in $\mathbf{x}\sim\mu^{n}$
-$\kappa\left(\frac{1}{n}\sum_{i}\mathbb{E}_{h\sim G_{\beta}\left(\mathbf{x}\right)}\left[f\left(h,x_{i}\right)\right],\mathbb{E}_{h\sim G_{\beta}\left(\mathbf{x}\right)}\mathbb{E}_{x}\left[f\left(h,x\right)\right]\right)\leq\frac{1}{n}\left(\Gamma\left(h,\mathbf{x},\bm{\beta}\right)+\ln\left(\frac{2\sqrt{n}}{\delta}\right)\right).$
-For the second part, we used the joint convexity of $\kappa$. Under the conditions of this corollary, the second inequality becomes
-$\mathbb{E}_{h\sim G_{\beta}\left(\mathbf{x}\right)}\mathbb{E}_{x}\left[f\left(h,x\right)\right]\leq\kappa^{-1}\left(\frac{1}{n}\sum_{i}\mathbb{E}_{h\sim G_{\beta}\left(\mathbf{x}\right)}\left[f\left(h,x_{i}\right)\right],\frac{1}{n}\left(\Gamma\left(h,\mathbf{x},\mathbf{\beta}\right)+\ln\left(\frac{2\sqrt{n}}{\delta}\right)\right)\right)$
-with an analogous version for the single-draw case. This is how we compute bounds in our experiments. For an illustration, please refer to C.2.1.
-Here $f$ plays the role of a secondary loss function, typically different from the loss function $\ell$, which defines the Gibbs posterior.
-In applications one would define the Gibbs posterior in terms of a differentiable loss function $\ell$ and approximate it with a suitable Monte Carlo method.
-For classification, however, one is interested in bounding the 0-1 loss obtained by comparing $\ell$ to some threshold $\theta$.
-Then (momentarily changing notation by replacing $x\in\mathcal{X}$ by $\left(x,y\right)$, where $y$ is the label corresponding to $x$) one can define $f\left(h,\left(x,y\right)\right)=1_{\left(-\infty,\theta\right)}\left(y\ell\left(h,x\right)\right)$ to obtain a bound on the expected 0-1 loss.
-4 Bounds for Langevin Monte Carlo
-For this section, we assume $\mathcal{H}=\mathbb{R}^{d}$ and an isotropic Gaussian prior $\pi$ of width $\sigma$.
-We condition on the training data $\mathbf{x}$, reference to which we omit.
-The Gibbs posterior is an idealization, from which it is impossible to sample directly.
-Nevertheless a number of works (Raginsky et al., 2017; Dalalyan & Karagulyan, 2017; Brosse et al., 2018; Vempala & Wibisono, 2019; Dwivedi et al., 2019; Nemeth & Fearnhead, 2021; Balasubramanian et al., 2022) discuss algorithms (SGLD, ULA, MALA, etc, here summarized as Langevin Monte Carlo (LMC)), capable of approximating a probability measure $\nu$ on $\mathbb{R}^{d}$ of the form $\nu\propto\exp\left(-V\right)$ or some nearby limiting distribution.
-In the following, we discuss one of these algorithms.
-4.1 ULA
-We focus on the results of Vempala & Wibisono (2019), which do not require convexity of $V$ and instead assume that the measure $\nu$ satisfies a log-Sobolev inequality (LSI) in the sense that for all smooth $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$
-$\mathbb{E}_{h\sim\nu}\left[f^{2}\left(h\right)\ln f^{2}\left(h\right)\right]-\mathbb{E}_{h\sim\nu}\left[f^{2}\left(h\right)\right]\ln\mathbb{E}_{h\sim\nu}\left[f^{2}\left(h\right)\right]\leq\frac{2}{\alpha}\mathbb{E}_{h\sim\nu}\left[\left\|\left(\nabla f\right)\left(h\right)\right\|^{2}\right]$ (5)
+$e^{\mathbb{E}_{g\sim\nu\left(\mathbf{x}\right)}\left[F\left(g,\mathbf{x}\right)-\ln\left(d\left(\nu\left(\mathbf{x}\right)/d\pi\right)\right)\left(g\right)\right]}\leq\mathbb{E}_{g\sim\nu\left(\mathbf{x}\right)}\left[e^{F\left(g,\mathbf{x}\right)-\ln\left(d\left(\nu\left(\mathbf{x}\right)/d\pi\right)\right)\left(g\right)}\right].$
+Then proceed as before. ∎
+B.2 Concrete PAC-Bayesian Bounds and the 0-1 Loss
+Assume that $\ell$ has values in $[0,1]$.
+To derive (2) from Theorem 2.1 let $F\left(h,\mathbf{x}\right)=n~\kappa\left(\hat{L}\left(h,\mathbf{x}\right),L\left(h\right)\right)$, where $\kappa$ is the relative entropy of two Bernoulli variables with expectations $p$ and $q$
+$\kappa\left(p,q\right)=p\ln\frac{p}{q}+\left(1-p\right)\ln\frac{1-p}{1-q}.$ (9)
+for which Tolstikhin and Seldin (2013) give the inversion rule $\kappa\left(p,q\right)\leq B$ $\implies$ $q-p\leq\sqrt{2pB}+2B$.
+Tonelli’s theorem and Theorem 1 of Maurer (2004) then give $\mathbb{E}_{\mathbf{x}}\mathbb{E}_{g\sim\pi}\left[e^{F\left(g,\mathbf{x}\right)}\right]=\mathbb{E}_{g\sim\pi}\mathbb{E}_{\mathbf{x}}\left[e^{F\left(g,\mathbf{x}\right)}\right]\leq 2\sqrt{n}$ for $n\geq 8$.
+Substitution in (i) of Theorem 2.1 and division by $n$ then gives with high probability as $\mathbf{x}\sim\mu^{n}$ and $h\sim\nu\left(\mathbf{x}\right)$
+$\kappa\left(\hat{L}\left(h,\mathbf{x}\right),L\left(h\right)\right)\leq\frac{1}{n}\left(\ln\frac{d\nu\left(\mathbf{x}\right)}{d\pi}+\ln\left(\frac{2\sqrt{n}}{\delta}\right)\right),$ (10)
+and the inversion rule of Tolstikhin and Seldin (2013) gives (2).
+The derivation of the bound for $\mathbb{E}_{h\sim\nu\left(\mathbf{x}\right)}\left[L\left(h\right)\right]$ from Theorem 2.1 (ii) is analogous, we just have to use $\kappa\left(\mathbb{E}_{h\sim\nu\left(\mathbf{x}\right)}\left[\hat{L}\left(h,\mathbf{x}\right)\right],\mathbb{E}_{h\sim\nu\left(\mathbf{x}\right)}\left[L\left(h\right)\right]\right)\leq\mathbb{E}_{h\sim\nu\left(\mathbf{x}\right)}\kappa\left(\hat{L}\left(h,\mathbf{x}\right),L\left(h\right)\right)$ by the joint convexity of $\kappa$ (see Cover (1999)).
+This gives
+$\kappa\left(\mathbb{E}_{h\sim\nu\left(\mathbf{x}\right)}\left[\hat{L}\left(h,\mathbf{x}\right)\right],\mathbb{E}_{h\sim\nu\left(\mathbf{x}\right)}\left[L\left(h\right)\right]\right)\leq\frac{1}{n}\left(KL\left(\nu\left(\mathbf{x}\right),\pi\right)+\ln\left(\frac{2\sqrt{n}}{\delta}\right)\right)$ (11)
+and the corresponding inequality obtained from the inversion rule.
+The inversion rule produces directly interpretable bounds, but stronger is the direct inversion using the function $\kappa^{1}:\left[0,1\right]\times\left[0,\infty\right)$ by
+$\kappa^{-1}\left(p,t\right)=\inf\left\{q:q\geq p,\kappa\left(p,q\right)\geq t\right\}.$
+In the definition of $F$, using the function $\kappa$, we can use other loss functions, possibly different from the loss function $\ell$ which define the Gibbs posterior.
+If these loss functions satisfy the conditions of Theorem 1 in Maurer (2004), we obtain analogous bounds.
+In particular for binary classification we can use the 01 loss.
+Momentarily changing notation by replacing $x\in\mathcal{X}$ by $\left(x,y\right)$, where $y$ is the label corresponding to $x$, the 01 loss is defined as
+$\hat{L}_{\textbf{01}}(h,\mathbf{x})=\frac{1}{n}\sum_{i=1}^{n}1_{\left(-\infty,0\right)}\left(y_{i}\ell\left(h,x_{i}\right)\right)$
+and $L_{\textbf{01}}(h)=\mathbb{E}_{x\sim\mu^{n}}\left[\hat{L}_{\textbf{01}}(h,\mathbf{x})\right]$. We then obtain the bounds
+$L_{\textbf{01}}\left(h\right)$ $\leq$ $\kappa^{-1}\left(\hat{L}_{\textbf{01}}(h,\mathbf{x}),\frac{1}{n}\left(\ln\frac{d\nu\left(\mathbf{x}\right)}{d\pi}+\ln\left(\frac{2\sqrt{n}}{\delta}\right)\right)\right)$
+$\mathbb{E}_{h\sim\nu\left(\mathbf{x}\right)}\left[L_{\textbf{01}}\left(h\right)\right]$ $\leq$ $\kappa^{-1}\left(\mathbb{E}_{h\sim\nu\left(\mathbf{x}\right)}\left[\hat{L}_{\textbf{01}}(h,\mathbf{x})\right],\frac{1}{n}\left(KL\left(\nu\left(\mathbf{x}\right),\pi\right)+\ln\left(\frac{2\sqrt{n}}{\delta}\right)\right)\right).$ (12)
+This is the bound used in our experiments.
+Appendix C Supplementary material for Section 4
+C.1 Examples of LMC: CLD and ULA
+A number of recent works give convergence guarantees for LMC algorithms and processes (Raginsky et al., 2017; Dalalyan and Karagulyan, 2017; Brosse et al., 2018; Vempala and Wibisono, 2019; Dwivedi et al., 2019; Nemeth and Fearnhead, 2021; Balasubramanian et al., 2022; Chen et al., 2022).
+Here we focus on the results of Vempala and Wibisono (2019), which do not require convexity of $V$ and instead assume that the measure $\nu$ satisfies a log-Sobolev inequality (LSI) in the sense that for all smooth $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$
+$\begin{split}\mathbb{E}_{h\sim\nu}\big[f^{2}(h)\ln f^{2}(h)\big]-\mathbb{E}_{h\sim\nu}\big[f^{2}(h)\big]\ln\mathbb{E}_{h\sim\nu}\big[f^{2}(h)\big]\leq\frac{2}{\alpha}\mathbb{E}_{h\sim\nu}\big[\|\nabla f(h)\|^{2}\big]\end{split}$ (13)
 for some $\alpha>0$.
-An LSI is satisfied when $V$ is strongly convex, but, importantly, also for measures which are bounded perturbations of measures satisfying an LSI (Holley & Stroock (1986)).
-Vempala & Wibisono (2019) give further examples and a list of references for measures, which are not log-concave and satisfy an LSI.
+An LSI is satisfied when $V$ is strongly convex, but, importantly, also for measures which are bounded perturbations of measures satisfying an LSI (Holley and Stroock, 1986).
+In our applications this will be ensured for bounded losses, because of the Gaussian prior, but $\alpha$ will deteriorate as $\beta$ increases.
+Vempala and Wibisono (2019) give further examples and a list of references for measures which are not log-concave and satisfy an LSI.
 Raginsky et al.
-(2017) show, that under dissipativity conditions of the loss the Gibbs posterior $G_{\beta}\left(\mathbf{x}\right)$ satisfies an LSI with constant independent of $\mathbf{x}$.
-Consider the iterative algorithm
+(2017) show that under dissipativity conditions of the loss the Gibbs posterior $G_{\beta}\left(\mathbf{x}\right)$ satisfies an LSI with constant independent of $\mathbf{x}$.
+There are a number of more recent works on this topic, in particular proximal methods (see, e.g., Chen et al.
+(2022)), but the work of Vempala and Wibisono (2019) is convenient for our purposes, because they guarantee convergence in relative entropy, and our emphasis is not on sampling.
+Continuous Langevin Dynamics (CLD) is specified by the stochastic differential equation
+$dh_{t}=-\nabla V\left(h_{t}\right)dt+\sqrt{2}dB_{t},$
+where $B_{t}$ is centered standard Brownian motion in $\mathbb{R}^{d}$.
+The distribution of CLD converges exponentially to the Gibbs posterior under mild conditions (Chiang et al., 1987).
+In Section C.2, we give a convergence result for CLD adapted to temperature dependence and prior, on the condition of a log-Sobolev inequality, with convergence in relative entropy.
+The Euler-discretization of CLD is the iterative algorithm
 $h_{t+1}=h_{t}-\epsilon\nabla V\left(h_{t}\right)+\sqrt{2\epsilon}\xi_{t},$ (14)
-where $\epsilon$ is a step size, the $\xi_{t}\sim\mathcal{N}\left(0,I\right)$ are independent Gaussian vectors and $h_{0}$ is drawn from some initial distribution $\nu_{0}$.
+where $\epsilon>0$ is a step size, the $\xi_{t}\sim\mathcal{N}\left(0,I\right)$ are independent Gaussian vectors and $h_{0}$ is drawn from some initial distribution $\nu_{0}$.
 Some authors call this algorithm simply LMC, for Langevin Monte Carlo.
-We call it ULA, alongside Durmus & Moulines (2017),Dwivedi et al.
-(2019) and Vempala & Wibisono (2019), for Un-adjusted Langevin Algorithm, because it misses the Metropolis-type accept-reject step, which would guarantee that the invariant distribution is indeed the Gibbs posterior.
-A popular variant of ULA is Stochastic Gradient Langevin Dynamics (SGLD) (Welling & Teh, 2011; Raginsky et al., 2017) where the gradient is replaced by an unbiased estimate, typically realized with random minibatches.
-Here, we restrict ourselves to ULA with a constant step size, because it has the least number of parameters to adjust, but in experiments we also use the more computationally efficient SGLD.
-As $\epsilon\,{\rightarrow}\,0$, ULA recovers the Continuous Langevin Dynamics (CLD) given by the stochastic differential equation
-$dh_{t}=-V\left(h_{t}\right)dt+\sqrt{2}dB_{t},$
-where $B_{t}$ is centered standard Brownian motion in $\mathbb{R}^{d}$.
-CLD converges exponentially to the Gibbs posterior (Chiang et al., 1987).
-For $\epsilon>0$, the distribution $\nu_{\epsilon,t}$ of ULA converges as $t\rightarrow\infty$ to a biased limiting distribution $\nu_{\epsilon}$ which is generally different from $\nu$, but expected to be closer to $\nu$ as $\epsilon$ becomes smaller.
-Vempala & Wibisono (2019) use the LSI assumption and coupling to control the difference between CLD and ULA along their path and prove the following result.
-Theorem 4.1.
+We call it ULA, alongside Durmus and Moulines (2017); Dwivedi et al.
+(2019); Vempala and Wibisono (2019), for Unadjusted Langevin Algorithm.
+A popular variant of ULA is Stochastic Gradient Langevin Dynamics (SGLD) (Welling and Teh, 2011; Raginsky et al., 2017) where the gradient is replaced by an unbiased estimate, typically realized with random minibatches.
+Here, we restrict ourselves to ULA with a constant step size, because it has the fewest parameters to adjust, but in experiments we also use the computationally more efficient SGLD.
+The distribution $\nu_{\epsilon,t}$ of ULA converges as $t\rightarrow\infty$ to a biased limiting distribution $\nu_{\epsilon}$, which is generally different from $\nu$, but expected to be closer to $\nu$ as $\epsilon$ becomes smaller.
+Vempala and Wibisono (2019) use the LSI assumption to control the difference between CLD and ULA along their path and prove the following result.
+Theorem C.1.
 Assume that $\nu$ satisfies the log-Sobolev inequality (13) with $\alpha>0$, that the Hessian of $V$ satisfies $LI\preceq\nabla^{2}V\left(h\right)\preceq LI$ for all $h$ and some $L<\infty$, and that $0<\epsilon\leq\alpha/\left(4L^{2}\right)$.
 Then, for $t\geq 0$
-$KL\left(\nu,\nu_{\epsilon,t}\right)\leq e^{-\alpha\epsilon t}KL\left(\nu,\nu_{0}\right)+\frac{8\epsilon dL^{2}}{\alpha}.$
+$\text{KL}\left(\nu_{\epsilon,t},\nu\right)\leq e^{-\alpha\epsilon t}\text{KL}\left(\nu_{0},\nu\right)+\frac{8\epsilon dL^{2}}{\alpha}.$
 The first exponential term is due to the mismatch of the initial distribution and $\nu$.
+Vempala and Wibisono (2019) show that $\nu_{0}$ may be chosen to make $\text{KL}\left(\nu_{0},\nu\right)$ of order $d$.
 The second term bounds the divergence between the limiting distribution $\nu_{\epsilon}\,\$and $\nu$.
-Similar results exist under different conditions on the potential $V$.
-Cheng et al.
-(2018) for example require $V$ to be strongly convex outside of a ball instead of the log-Sobolev inequality and gives bounds in terms of the $W_{1}$-Wasserstein metric.
+Similar results exist under different conditions on the potential $V$; Cheng et al.
+(2018) for example, require $V$ to be strongly convex outside of a ball instead of the log-Sobolev inequality and give bounds in terms of the $W_{1}$-Wasserstein metric.
 Raginsky et al.
 (2017) give bounds for $W_{2}$ under dissipativity assumptions.
-The next corollary adapts Theorem 4.1 to the situation studied in this paper.
-Corollary 4.2.
+We are not aware of similar bounds for the Rényi-infinity divergence.
+The next corollary adapts Theorem C.1 to the situation studied in this paper.
+Corollary C.2.
 For $\beta>0$ consider the Gibbs posterior $G_{\beta}$ corresponding to $\hat{L}\left(h\right)$, with centered Gaussian prior of width $\sigma$.
@@
 where $h_{0}\sim\nu_{0}$ and the $\xi_{t}\sim\mathcal{N}\left(0,I\right)$ are independent Gaussian random variables.
-Let $D\left(\beta\right)=KL\left(G_{\beta},\nu_{0}\right)$ and let $\nu_{\beta,\eta,t}$ be the distribution of $h_{t}$ after $t$ steps.
-Then,
-(i) $KL\left(G_{\beta},\nu_{\beta,\eta,t}\right)\leq e^{-\alpha\eta t/\beta}D\left(\beta\right)+\frac{8\eta d}{\beta\alpha}\left(\beta R+\frac{1}{\sigma^{2}}\right)^{2}.$
-(ii) $W_{2}\left(G_{\beta},\nu_{\beta,\eta,t}\right)\leq\frac{2}{\alpha}e^{-\alpha\eta t/\beta}D\left(\beta\right)+\frac{16\eta d}{\beta\alpha^{2}}\left(\beta R+\frac{1}{\sigma^{2}}\right)^{2}.$
-(iii) $d_{TV}\left(G_{\beta},\nu_{\beta,\eta,t}\right)\leq e^{-\alpha\eta t/\left(2\beta\right)}\sqrt{D\left(\beta\right)}+2\sqrt{\frac{\eta d}{\beta\alpha}}\left(\beta R+\frac{1}{\sigma^{2}}\right).$
+Let $D\left(\beta\right)=\text{KL}\left(\nu_{0},G_{\beta}\right)$ and let $\nu_{\beta,\eta,t}$ be the distribution of $h_{t}$ after $t$ steps.
+Then
+$\text{KL}(\nu_{\beta,\eta,t}\|G_{\beta})\leq e^{-\frac{\alpha\eta t}{\beta}}D(\beta)+\tfrac{8\eta d}{\beta\alpha}\big(\beta R\,{+}\tfrac{1}{\sigma^{2}}\big)^{2}$
 Proof.
-We make the identifications $V\left(h\right)=\beta\hat{L}\left(h\right)+\left\|h\right\|^{2}/\left(2\sigma^{2}\right)$, $\epsilon=\eta/\beta$ and $L=\beta R+\frac{1}{\sigma^{2}}$.
-Then, $\nu=G_{\beta}$ with Gaussian prior of width $\sigma$, ULA becomes (7) and (i) follows directly from Theorem 4.1, (ii) follows from Otto & Villani (2000) and the LSI assumption, and (iii) follows from Pinsker’s inequality (see e.g.
-Boucheron et al.
-(2013)).
+This follows directly from Theorem C.1 and the substitutions $V\left(h\right)=\beta\hat{L}\left(h\right)+\left\|h\right\|^{2}/\left(2\sigma^{2}\right)$, $\epsilon=\eta/\beta$ and $L=\beta R+\frac{1}{\sigma^{2}}$.
+Then $\nu=G_{\beta}$ with Gaussian prior of width $\sigma$ and ULA becomes (15).
 ∎
-4.2 Stability of the bounds
-We now show the stability of our bounds for approximation in total variation and $W_{2}$-Wasserstein metrics, under boundedness or Lipschitz conditions.
-Together with Corollary 4.2 this implies bounds for the algorithm defined in (7).
-We assume that there is a target approximation $\nu_{\beta}\left(\mathbf{x}\right)$ of $G_{\beta}\left(\mathbf{x}\right)$, for which we want to compute a high probability bound, either for the single draw version on $F\left(h,\mathbf{x}\right)$ as $\mathbf{x}\sim\mu^{n}$ and $h\sim\nu_{\beta}\left(\mathbf{x}\right)$, or, for the classical PAC-Bayesian version, on $\mathbb{E}_{h\sim\nu_{\beta}\left(\mathbf{x}\right)}\left[F\left(h,\mathbf{x}\right)\right]$ as $\mathbf{x}\sim\mu^{n}$.
-It is not surprising that the single-draw bound will require a much closer approximation of the Gibbs posterior.
-Since the bounding functional $\Gamma\left(h,\mathbf{x},\bm{\beta}\right)$ depends on the Gibbs posteriors $G_{\beta_{k}}\left(\mathbf{x}\right)$ for $k\in\{1,\dots,K-1\}$, we require corresponding approximations $\nu_{\beta_{k}}$ of $G_{\beta_{k}}\left(\mathbf{x}\right)$ to compute the bound.
-To streamline notation, we define
-$\Gamma_{\nu}\left(h,\mathbf{x},\bm{\beta}\right)=-\beta\hat{L}\left(h,\mathbf{x}\right)+\sum_{k=1}^{K}\left(\beta_{k}-\beta_{k-1}\right)\mathbb{E}_{g\sim\nu_{\beta_{k-1}}\left(\mathbf{x}\right)}\big[\hat{L}(g,\mathbf{x})\big]$
-for $0=\beta_{0}<\beta_{1}<\cdots<\beta_{K}=\beta$ and $\left(\nu_{0}\left(\mathbf{x}\right),\nu_{\beta_{1}}\left(\mathbf{x}\right),\cdots,\nu_{\beta_{k-1}}\left(\mathbf{x}\right)\right)\in\mathcal{P}\left(\mathcal{H}\right)^{K}$.
-Like $\Gamma\left(h,\mathbf{x},\mathbf{\beta}\right)$, the functional $\Gamma_{\nu}\left(h,\mathbf{x},\mathbf{\beta}\right)$ depends on the training data, but it can also be computed by repeated execution of the algorithm (7).
-The next theorem states the obtained bound in terms of the approximation errors in total variation.
-Theorem 4.3.
-Suppose that $\mathcal{H}=\mathbb{R}^{d}$ and that there are numbers $m,M<\infty$ such that for every $\mathbf{x}$ in $\mathcal{X}^{n}$ and $h\in\mathcal{H}$ we have $\left|\ell\left(h,\mathbf{x}\right)\right|\leq m$ and $\left|F\left(h,\mathbf{x}\right)\right|\leq M$.
-Let $0=\beta_{0}<\beta_{1}<\dots<\beta_{K}=\beta$ and $\nu_{\beta_{k}}\left(\mathbf{x}\right)\in\mathcal{P}\left(\mathcal{H}\right)$ be such that $d_{TV}\left(\nu_{\beta_{k}}\left(\mathbf{x}\right),G_{\beta_{k}}\left(\mathbf{x}\right)\right)=\epsilon_{\beta_{k}}$.
-Then,
-(i) with probability at least $1-\delta$ as $\mathbf{x}\sim\mu^{n}$ and $h\sim\nu_{\beta}\left(\mathbf{x}\right)$
-$F\left(h,\mathbf{x}\right)\leq\Gamma_{\nu}\left(h,\mathbf{x},\mathbf{\beta}\right)+\ln\mathbb{E}_{\mathbf{x}}\mathbb{E}_{h\sim\pi}\left[e^{F\left(h,\mathbf{x}\right)}\right]+\ln\frac{1}{\delta}+\ln\left(2e^{M+\beta m}\epsilon_{\beta}\right)+\sum_{k=1}^{K}\left(\beta_{k}-\beta_{k-1}\right)m\epsilon_{\beta_{k-1}}.$
-(ii) with probability at least $1-\delta$ as $\mathbf{x}\sim\mu^{n}$
-$\mathbb{E}_{h\sim\nu_{\beta}\left(\mathbf{x}\right)}\left[F\left(h,\mathbf{x}\right)\right]$ $\leq$ $\mathbb{E}_{h\sim\nu_{\beta}\left(\mathbf{x}\right)}\left[\Gamma_{\nu}\left(h,\mathbf{x},\bm{\beta}\right)\right]+\ln\mathbb{E}_{\mathbf{x}}\mathbb{E}_{h\sim\pi}\left[e^{F\left(h,\mathbf{x}\right)}\right]+\ln\frac{1}{\delta}$
-$+\left(M+\beta m\right)\epsilon_{\beta}+\sum_{k=1}^{K}\left(\beta_{k}-\beta_{k-1}\right)m\epsilon_{\beta_{k-1}}.$
-The proof, given in Section B.2, is similar to that of Theorem 3.2 and applies Markov’s inequality with $\nu_{\beta}$ instead of $G_{\beta}$.
-It then uses the fact that, if $f$ is a bounded measurable function, then $\left|\mathbb{E}_{\nu_{1}}\left[f\right]-\mathbb{E}_{\nu_{2}}\left[f\right]\right|\leq\left\|f\right\|_{\infty}d_{TV}\left(\nu_{1},\nu_{2}\right)$.
-For the single-draw version (i) this is applied to $\left(\mathbb{E}_{\nu_{\beta}}-\mathbb{E}_{G_{\beta}}\right)\left[e^{F\left(h,\mathbf{x}\right)+\beta\hat{L}\left(h,\mathbf{x}\right)+\ln Z_{\beta}\left(\mathbf{x}\right)}\right]$, which causes the exponential dependence on $\beta m$ and $M$.
-For (ii) we can apply this in the exponent to $\left(\mathbb{E}_{\nu_{\beta}}-\mathbb{E}_{G_{\beta}}\right)\big[F(h,\mathbf{x})+\beta\hat{L}(h,\mathbf{x})\big],$ and the logarithm makes the dependence linear.
-The rest of the proof is mechanical.
-The last terms in (i) and (ii) are the additional errors due to the approximations of the Gibbs posteriors.
-The worst term is clearly the first one in (i) due to the exponential dependence on the proxy function $F$, which is typically of order $n$ and on $\beta$, which is larger than $n$ in the regime in which we are interested.
-In practice, the requirement of such an approximation is prohibitive for the single-draw version.
-The bound in (ii) has more moderate approximation requirements.
-In this case, we can also give bounds in terms of the $W_{2}$-Wasserstein metric (as guaranteed by Corollary 4.2), if $F$ sand $\ell$ satisfy a Lipschitz condition instead of boundedness.
+C.2 Convergence of CLD under LSI
+The following is a straightforward adaptation of Theorem 1 and its proof in Vempala and Wibisono (2019). There is no claim to originality.
+Lemma C.3.
+Let the process $h_{\beta,t}$ on $\mathbb{R}^{d}$ be defined by the stochastic differential equation
+$dh_{\beta,t}\left(\mathbf{x}\right)=-\left(\nabla V\left(h_{\beta,t}\right)+\frac{h_{\beta,t}}{\beta\sigma^{2}}\right)dt+\sqrt{2/\beta}dB_{t},$ (16)
+and suppose that the measure with density $G_{\beta}\left(h\right):=\exp\left(-\left(\beta V+\frac{\left\|h\right\|^{2}}{2\sigma^{2}}\right)\right)$ satisfies an LSI with constant $\alpha$.
+Then, if $\nu_{\beta,t}$ is the distribution of $h_{\beta,t}$,
+$KL\left(\nu_{\beta,t},G_{\beta}\right)\leq e^{-2\alpha t/\beta}KL\left(\nu_{\beta,0},G_{\beta}\right).$
+Proof.
+Let $U_{\beta,t}=-\beta^{-1}\left(\ln\nu_{\beta,t}+\left\|h\right\|^{2}/2\sigma^{2}\right)$, so that $\nu_{\beta,t}=\exp\left(-\beta U_{\beta,t}-\frac{\left\|h\right\|^{2}}{2\sigma^{2}}\right)$ and
+$KL\left(\nu_{\beta,t},G_{\beta}\right)=\beta\int_{\mathbb{R}^{d}}\left(V-U_{\beta,t}\right)\nu_{\beta,t}d\lambda,$
+where $\lambda$ is Lebesgue measure on $\mathbb{R}^{d}$. The Fokker-Planck equation for (16) becomes
+$\frac{\partial\nu_{\beta,t}}{\partial t}$ $=$ $\nabla\cdot\left(\nu_{\beta,t}\nabla\left(V+\frac{\left\|h\right\|^{2}}{2\beta\sigma^{2}}\right)\right)+\beta^{-1}\Delta\nu_{\beta,t}$
+$=$ $\nabla\cdot\left(\nu_{\beta,t}\nabla\left(V-U_{\beta,t}\right)\right).$
+We have $0=\frac{d}{dt}\int\nu_{\beta,t}d\lambda=-\int\left(\frac{\partial}{\partial t}\beta U_{\beta,t}\right)\nu_{\beta,t}dt$, so with integration by parts
+$\frac{d}{dt}KL\left(\nu_{\beta,t},G_{\beta}\right)$ $=$ $\int_{\mathbb{R}^{d}}\beta\left(V-U_{\beta,t}\right)\left(\nabla\cdot\left(\nu_{\beta,s}\nabla\left(V-U_{\beta,t}\right)\right)\right)d\lambda$
+$=$ $-\beta\int_{\mathbb{R}^{d}}\left\langle\nabla\left(V-U_{\beta,t}\right),\nabla\left(V-U_{\beta,t}\right)\right\rangle\nu_{\beta,s}d\lambda$
+$=$ $-\beta^{-1}\int_{\mathbb{R}^{d}}\left\|\nabla\left(\beta V-\beta U_{\beta,t}\right)\right\|^{2}\nu_{\beta,t}d\lambda$
+$=$ $-\beta^{-1}J\left(\nu_{\beta,t},G_{\beta}\right)$
+$\leq$ $-2\alpha\beta^{-1}KL\left(\nu_{\beta,t},G_{\beta}\right),$
+where $J\left(\nu,\rho\right)=E_{\nu}\left\|\nabla\ln\frac{d\nu}{d\rho}\right\|^{2}$ is the relative Fisher information and the last inequality follows from the LSI.
+Integrating this inequality concludes the proof.
+∎
+C.3 The "Second law of thermodynamics"
+Restatement of Lemma 4.1
+Lemma C.4.
+Under the assumptions of Section 4.1, if $\nu$ is a stationary distribution of $\left\{h_{t}\right\}_{t\in I}$ and $s<t$ then $KL\left(\nu_{t},\nu\right)\leq KL\left(\nu_{s},\nu\right)$ and $R_{\infty}\left(\nu_{t},\nu\right)\leq R_{\infty}\left(\nu_{s},\nu\right)$, with equality in either case if and only if $\nu_{s}=\nu$.
+Proof.
+We identify measures with their densities with respect to some base distribution $\mu$, The transition kernel is denoted $K\left(h,g\right)=\mathbb{P}\left(h_{t-s}=h|h_{0}=g\right)=\mathbb{P}\left(h_{t}=h|h_{s}=g\right)$.
+$\nu_{t}\left(h\right)\ln\frac{\nu_{t}\left(h\right)}{\nu\left(h\right)}$ $=$ $\int K\left(h,g\right)\nu_{s}\left(g\right)d\mu\left(g\right)\ln\frac{\int K\left(h,g\right)\nu_{s}\left(g\right)d\mu\left(g\right)}{\int K\left(h,g\right)\nu\left(g\right)d\mu\left(g\right)}$ (17)
+$\leq$ $\int K\left(h,g\right)\nu_{s}\left(g\right)\ln\frac{\nu_{s}\left(g\right)}{\nu\left(g\right)}d\mu\left(g\right),$
+with equality if and only if $\nu_{s}=\nu$.
+The first identity is owed to the invariance of $\nu$.
+The inequality is the log-sum inequality (Cover, 1999), followed by cancellation of $K\left(h,g\right)$.
+Then by Fubini’s theorem
+$KL\left(\nu_{t},\nu\right)$ $=$ $\int\nu_{t}\left(h\right)\ln\frac{\nu_{t}\left(h\right)}{\nu\left(h\right)}d\mu\left(h\right)$
+$\leq$ $\int\left(\int K\left(h,g\right)d\mu\left(h\right)\right)\nu_{s}\left(g\right)\ln\frac{\nu_{s}\left(g\right)}{\nu\left(g\right)}d\mu\left(g\right)$
+$=$ $\int\nu_{s}\left(g\right)\ln\frac{\nu_{s}\left(g\right)}{\nu\left(g\right)}d\mu\left(g\right)=KL\left(\nu_{s},\nu\right),$
+which gives the first inequality. To prove the second inequality, divide (17) by $\nu_{t}\left(h\right)$ to get with Hölder’s inequality
+$\ln\frac{\nu_{t}\left(h\right)}{\nu\left(h\right)}$ $\leq$ $\frac{1}{\nu_{t}\left(h\right)}\int K\left(h,g\right)\nu_{s}\left(g\right)\ln\frac{\nu_{s}\left(g\right)}{\nu\left(g\right)}d\mu\left(g\right)$
+$\leq$ $\frac{1}{\nu_{t}\left(h\right)}\int K\left(h,g\right)\nu_{s}\left(g\right)d\mu\left(g\right)\left(\sup_{g}\ln\frac{\nu_{s}\left(g\right)}{\nu\left(g\right)}\right)$
+$=$ $R_{\infty}\left(\nu_{s},\nu\right).$
+Take the supremum in $h$ to get the second inequality. ∎
+C.4 Proofs for Section 4.2
+We assume that $\mathcal{H}=\mathbb{R}^{d}$ and prepare the proof of Lemma 4.4.
+The total variation distance is defined as $d_{TV}:(\rho,\nu)\in\mathcal{P}(\mathcal{H})\times\mathcal{P}(\mathcal{H})\mapsto\sup_{A\in\Omega}|\rho(A)-\nu(A)|$.
+If $f$ is a bounded measurable function, then
+$\left|\mathbb{E}_{\rho}\left[f\right]-\mathbb{E}_{\nu}\left[f\right]\right|\leq\left\|f\right\|_{\infty}d_{TV}\left(\rho,\nu\right).$
+By Pinsker’s inequality (see, e.g. Boucheron et al. (2013))
+$d_{TV}\left(\rho,\nu\right)\leq\sqrt{2KL\left(\rho,\nu\right)}.$
+The $W_{p}$-Wasserstein distance is $W_{p}(\rho,\nu)=(\inf_{W}\mathbb{E}_{(x,y)\sim W}[\|x-y\|^{p}])^{1/p}$ with the infimum being over all probability measures on $\mathcal{P}(\mathcal{H}\times\mathcal{H})$ with $\rho$ and $\nu$ as marginals.
 We will use the following fact: Since $W_{1}\leq W_{2}$ it follows from the Kantorovich-Rubinstein Theorem (Villani, 2009), that for any real Lipschitz function $f$ on $\mathcal{H}$ and probability measures $\nu_{1},\nu_{2}\in\mathcal{P}\left(\mathcal{H}\right)$
 $\left|\mathbb{E}_{h\sim\nu_{1}}\left[f\left(h\right)\right]-\mathbb{E}_{h\sim\nu_{2}}\left[f\left(h\right)\right]\right|\leq\left\|f\right\|_{\rm Lip}W_{1}\left(\nu_{1},\nu_{2}\right)\leq\left\|f\right\|_{\rm Lip}W_{2}\left(\nu_{1},\nu_{2}\right),$
-where $\left\|.\right\|_{\rm Lip}$ is the Lipschitz-seminorm. The following result is then immediate, with proof exactly as in (ii) of Theorem 4.3.
-Theorem 4.4.
-Assume the conditions of Theorem 4.3, except that instead of $\left|\ell\left(h,\mathbf{x}\right)\right|\leq m$ and $\left|F\left(h,\mathbf{x}\right)\right|\leq M$ we have $\left\|\ell\left(.,\mathbf{x}\right)\right\|_{\rm Lip}\leq m$ and $\left\|F\left(.,\mathbf{x}\right)\right\|_{\rm Lip}\leq M$ and that $W_{2}\left(\nu_{\beta_{k}}\left(\mathbf{x}\right),G_{\beta}\left(\mathbf{x}\right)\right)=\epsilon$.
-Then, with probability at least $1-\delta$ as $\mathbf{x}\sim\mu^{n}$
-$\mathbb{E}_{h\sim\nu_{\beta}\left(\mathbf{x}\right)}\left[F\left(h,\mathbf{x}\right)\right]$ $\leq$ $\mathbb{E}_{h\sim\nu_{\beta}\left(\mathbf{x}\right)}\left[\Gamma_{\nu}\left(h,\mathbf{x},\bm{\beta}\right)\right]+\ln\mathbb{E}_{\mathbf{x}}\mathbb{E}_{h\sim\pi}\left[e^{F\left(h,\mathbf{x}\right)}\right]+\ln\frac{1}{\delta}$
-$+\left(M+\beta m\right)\epsilon_{\beta}+\sum_{k=1}^{K}\left(\beta_{k}-\beta_{k-1}\right)m\epsilon_{\beta_{k-1}}.$
-5 Experiments
-The purpose of our experiments is to show that our method gives nontrivial bounds on the test error for real-world data, while correctly bounding the test error on impossible data, where the same algorithm also achieves a small training error.
-For this demonstration, we use binary classification tasks.
-The real-world data are either the MNIST dataset, subdivided into the two classes of characters 0-4 and 5-9, or the CIFAR-10 dataset to distinguish between animals and vehicles.
-For impossible data, we randomize the labels of the training data.
-Our experiments are computationally heavy, so we generally use small sample sizes, from 2000 to 8000 examples.
-The hypothesis space is the set of weight vectors for a neural network with ReLU activation functions constrained by a Gaussian prior distribution with $\sigma=5$.
-Neural network architectures are described in Section C.1.1 of the appendix.
-To approximately sample the weight vectors in the vicinity of the Gibbs posterior, we use ULA as in (7) or SGLD (Welling & Teh, 2011) with constant step size $\eta$.
-To ensure reproducibility, we provide the code and experimental results in an anonymous repository at https://anonymous.4open.science/r/Gibbs-Generalization-45F1.
-5.1 The loss function $\ell$
-Unless $\sigma$ is very small, which would prevent near-zero training errors for random labels, an unbounded loss function like binary cross-entropy entropy (BCE) would cause $\mathbb{E}_{h\sim G_{0}}\big[\hat{L}(h)\big]=\mathbb{E}_{h\sim\pi}\big[\hat{L}(h)\big]$ and therefore the first term in the summation of $\Gamma$ in equation 3 to be very large.
-For this reason we use a bounded loss function $\ell$, either bounded binary cross-entropy as described in Appendix D of Dziugaite & Roy (2018) or the savage loss (Masnadi-Shirazi & Vasconcelos, 2008).
-We compute bounds for the 0-1 loss, using the method described in Section 3.3.
-5.2 Approximating the ergodic mean
-As we know of no sufficient criterion for convergence, we terminate iterations at time $T$, when a very slow running mean $\mathbb{M}_{\rm stop}$ of the loss trajectory $\big(\hat{L}(h_{\beta_{k}t},\mathbf{x})\big)_{t=0}^{T}$ stops decreasing.
-A second running mean $\mathbb{M}_{\rm erg}$ is used as an approximation of the ergodic mean and thus of expectations in the invariant distribution.
-We thus replace all expectations $\mathbb{E}_{h\sim G_{\beta_{k}}}\big[\hat{L}(h,\mathbf{x})\big]$ occurring in the bounds by $\mathbb{M}_{\rm erg}\big[\big(\hat{L}(h_{\beta_{k}t},\mathbf{x})\big)_{t=0}^{T}\big]$.
-Both running means $\mathbb{M}_{\rm stop}$ and $\mathbb{M}_{\rm erg}$ are implemented as first-order, recursive lowpass filters described in Section C.1.3 of the appendix.
-5.3 Calibration
-The theoretical bounds of Corollary 4.2 in combination with Theorem 4.3 can only serve as guidance for the computation of practical bounds, because the quantities $R$ and $\alpha$ are impossible to estimate in practice.
-But even if we assume these to be in the order of unity, the bounds are too coarse to distinguish between different temperatures with realistic stepsizes.
-A simple calculation shows that $KL\left(G_{\beta},G_{2\beta}\right)\leq\beta\left(\mathbb{E}_{G_{\beta}}\left[\hat{L}\right]-\mathbb{E}_{G_{2\beta}}\left[\hat{L}\right]\right)\leq\beta\mathbb{E}_{G_{\beta}}\left[\hat{L}\right$] (Lemma B.5 in Appendix B.3).
-By Corollary 4.2 we should therefore have at least $8\eta dR^{2}/\alpha<\mathbb{E}_{G_{\beta}\left(\mathbf{x}\right)}\left[\hat{L}\right$] to distinguish between the expectations in the Gibbs posterior for $\beta$ and $2\beta$.
-The smallest neural network we use has $d=392,500$.
-If $\ell$ has values in $\left[0,1\right]$ then $\mathbb{E}_{G_{\beta}\left(\mathbf{x}\right)}\left[\hat{L}\right]\leq 1$, so even if $R$ and $\alpha$ are set to $1$, we would need stepsizes in the order of $10^{-7}$.
-Safe values of $\eta$, as suggested by the theoretical results in Section 4, are therefore impossible in practice, and the bound has to be adapted to a realistic choice of $\eta$.
-The first possibility is to simply ignore the un-estimable error terms in Theorem 4.3, and to compute the bounds from $\Gamma_{\nu}\left(h,\mathbf{x},\mathbf{\beta}\right)$ in place of $\Gamma\left(h,\mathbf{x},\mathbf{\beta}\right)$ as in Theorem 3.2.
-We show some of these uncalibrated bounds, which already make nontrivial predictions, in Section C.2.6.
-A second method uses a single calibration parameter, whose value is computed from the training data.
-We assume that the computed functional $\Gamma_{\nu}\left(h,\mathbf{x},\mathbf{\beta}\right)$ fails to estimate $\Gamma\left(h,\mathbf{x},\mathbf{\beta}\right)$ by a factor $r\left(\mathbf{x}\right)>0$, which we compute as
-$r\left(\mathbf{x}\right)=\min\left\{r:\forall k\in\left[K\right],\text{ }\kappa^{-1}\left(\mathbb{E}_{h\sim\nu_{\beta_{k}}\left(\mathbf{x}\right)}\left[\hat{L}_{01}\left(h,\mathbf{\tilde{x}}\right)\right],\frac{1}{n}\left(r\Gamma_{\nu}\left(h,\mathbf{\tilde{x}},\mathbf{\beta}_{1}^{k}\right)+\ln\frac{2\sqrt{n}}{\delta}\right)\right)\geq\frac{1}{2}\right\},$
-where $\mathbf{\tilde{x}}$ is the training set $\mathbf{x}$ with random labels and $\hat{L}_{01}$ the empirical 01-error.
-The calibration value $r$ is thus the smallest factor of $\Gamma_{\nu}$, for which we obtain a correct upper bound on the 01-error with random labels for all the $\beta_{k}$.
-By definition of $\Gamma_{\nu}$, replacing it by $r\Gamma_{\nu}$ is equivalent to a multiplicative shift of the temperature scale.
-It is a purely experimental finding, that our choice of $r$ leads to correct and surprisingly tight upper bounds on the test error of correctly labeled data in all cases we tried.
-We emphasize that our calibration procedure depends only on the training data.
-5.4 Results
-Figure 1: SGLD on MNIST and CIFAR-10 with 8000 training examples, MNIST above and CIFAR-10 below, random labels on the left, correct labels on the right.
-Both random and true labels are trained with the same algorithm and parameters on a fully connected ReLU network with two hidden layers of 1000 and 1500 units, respectively.
-The calibration factor for MNIST is 0.19, for CIFAR-10 0.22.
-Train error, test error and our bound for the Gibbs posterior average of the 0-1 loss are plotted against $\beta$.
-Several experiments confirm the validity of the proposed bounds.
-An example is shown in Figure 1, where a fully connected ReLU-network with two hidden layers of 1000 (respectively 1500) units each is trained with SGLD at inverse temperatures $\beta=$ 0, 500, 1000, 2000, 4000, 8000, 16000, 32000, and 64000.
-The train error for random labels is about 0.1 (or even less) at $\beta=64000$, where the bound is above 0.5.
-The test error for correct labels, however, is tightly bounded above.
-Notice that for MNIST, which has the tightest bounds, the training error for the true labels is rapidly decreasing from 0.5 to 0.17 at $\beta=500$ and to 0.1 at $\beta=1000$.
-The more moderate initial decrease for CIFAR-10 corresponds to the tendency to overfit on this more difficult dataset.
-This confirms the intuition, that good generalization at low temperatures is already announced in the high temperature regime.
-We generally found the uncalibrated bounds for ULA tighter than those for SGLD, consistent with the findings of Brosse et al.
-(2018).
-Experimental bounds for single draws from the posterior and various other experiments are reported in Section C.2.
-6 Conclusion
-Using the integral representation of the log-partition function, the Gibbs posterior admits the computation of upper bounds on the true error based on the training data and for any temperature.
-These bounds are stable under perturbation in the total-variation and Wasserstein metrics, and can be approximated by Langevin Monte Carlo (LMC) algorithms.
-However, for realistic experiments, the approximations obtained by these algorithms are coarse and require calibration, which leads to rather tight bounds in the interpolation regime of overparametrized neural networks.
-The fact that the calibrated bounds are very tight is, at this point, a purely experimental finding, requiring more theoretical investigation in future work.
-Acknowledgments
-This work was supported in part from the PNRR MUR Project PE000013 CUP J53C22003010006 "Future Artificial Intelligence Research (FAIR)", funded by the European Union – NextGenerationEU and by European Lighthouse of AI for Sustainability, Grant N.
-101120237.
-Appendix
-In this appendix, we summarize a glossary of notation, give additional theoretical results and missing proofs, and provide more information on the numerical experiments, as well as additional experimental results.
-Appendix A Table of notation
-Notation Brief description Section
-$\mathcal{X}$ space of data 2
-$\Sigma$ sigma algebra (events) on $\mathcal{X}$ 2
-$\mu$ probability of data 2
-$n$ sample size 1, 2
-$\mathbf{x}$ generic member $\left(x_{1},...,x_{n}\right)\in\mathcal{X}^{n}$, training sample 2
-$\mathcal{H}$ hypothesis space 2
-$\Omega$ sigma algebra (events) on $\mathcal{H}$ 2
-$\ell$ $\ell:\mathcal{H\times X\rightarrow}\left[0,\infty\right)$ loss function 2
-$f$ secondary loss function 3.3
-$\mathcal{P}\left(\mathcal{H}\right)$ probability measures on $\mathcal{H}$ 2
-$\pi$ nonnegative a-priori measure on $\mathcal{H}$ 2
-$\sigma$ width of Gaussian prior 4
-$L\left(h\right)$ $L\left(h\right)=\mathbb{E}_{x\sim\mu}\left[h\left(x\right)\right]$, expected loss of $h\in\mathcal{H}$ 2
-$\hat{L}\left(h,\mathbf{x}\right)$ $\hat{L}\left(h,\mathbf{x}\right)=\left(1/n\right)\sum_{i=1}^{n}\ell\left(h,x_{i}\right)$, empirical loss of $h\in\mathcal{H}$ 2
-$\beta$ inverse temperature 1, 2
-$Z_{\beta}\left(\mathbf{x}\right)$ partition function 2
-$G_{\beta,\pi}\left(\mathbf{x}\right)$ Gibbs posterior with energy $\hat{L}$ and prior $\pi$ 2
-$\mathbb{E}_{g\sim G_{\beta}\left(\mathbf{x}\right)}$ posterior expectation 2
-$\bm{\beta}$ increasing sequence $\left(\beta_{1}<...<\beta_{K}\right)$ of positive reals 3.1
-$\Gamma\left(h,\mathbf{x},\bm{\beta}\right)$ bounding functional 3.1
-$F\left(h,\mathbf{x}\right)$ placeholder for generalization gap 3.2
-$\kappa$ $kl\left(p,q\right)=p\ln\frac{p}{q}+\left(1-p\right)\ln\frac{1-p}{1-q}$, rel. entropy of Bernoulli variables 2
-$KL\left(\rho,\nu\right)$ $\int\left(\ln\frac{d\rho}{d\nu}\right)d\rho$, KL-divergence of $\rho,\nu\in\mathcal{P}\left(\mathcal{H}\right)$ 2, 4.1, 4.2
-$d_{TV}\left(\rho,\nu\right)$ total variation distance 2, 4.1, 4.2
-$W_{p}\left(\rho,\nu\right)$ $p$-Wasserstein distance 2, 4.1, 4.2
-$\Gamma_{\nu}\left(h,\mathbf{x},\bm{\beta}\right)$ LMC approximation of $\Gamma\left(h,\mathbf{x},\bm{\beta}\right)$ 4.2
-$\eta$ step size or learning rate 4.1
-$\nu_{\beta,\eta}$ invariant measure of LMC approximation of $G_{\beta}$ with step size $\eta$ 4.1
-$\nu_{\beta,\eta,t}$ LMC approximation of $G_{\beta}$ with step size $\eta$ at iteration $t$ 4.1
-$r({\mathbf{x}})$ calibration factor 5.3
-$\mathbf{\tilde{x}}$ randomly labeled data 5.3
-$\mathbb{M}_{\rm stop}$$\mathbb{M}_{\rm erg}$ filters for stopping and ergodic mean 5.2, C.1.3
-Appendix B Additional results and proofs
-B.1 Sub-gaussian losses
-The freedom in the choice of $F$ allows a number of bounds to be derived from Theorem 3.2.
-A centered real random variable $Y$ is called $\sigma$-sub-Gaussian if $\ln\mathbb{E}e^{\lambda Y-\mathbb{E}Y}\leq\lambda^{2}\sigma^{2}/2$ for all $\lambda\in\mathbb{R}$.
-Now, suppose that for some real function $f$ all the $x\in\mathcal{X\mapsto}f\left(h,x\right)$ are $\sigma$-sub-Gaussian as $x\sim\mu$.
-Let $\hat{L}(h,\mathbf{x})=\frac{1}{n}\sum_{i=1}^{n}f(h,x_{i})$ and $L(h)=E_{x}[f(h,\mathbf{x})]$.
-Then, $\mathbf{x}\in\mathcal{X}^{n}\mapsto\hat{L}\left(h,\mathbf{x}\right)$ as $\mathbf{x}\sim\mu^{n}$ is $\sigma/\sqrt{n}$-sub-Gaussian.
-It is tempting to set $F\left(h,\mathbf{x}\right)=\lambda\left(L\left(h\right)-\hat{L}\left(h,\mathbf{x}\right)\right)$ in Theorem 3.2, divide by $\lambda$ and then optimize over $\lambda$.
-Unfortunately, the last step is impossible, since the optimal $\lambda$ is data-dependent in its dependence on $\Gamma$ and ruins the exponential moment bound on $F$.
-A more careful argument establishes the following.
-Corollary B.1.
-Suppose that for all $h\in\mathcal{H}$ the random variables $x\in\mathcal{X\mapsto}f\left(h,x\right)$ as $x\sim\mu$ are $\sigma$-sub-Gaussian.
-For $\delta>0$ with probability at least $1-\delta$ as $\mathbf{x}\sim\mu^{n}$ and $h\sim G_{\beta,\pi}\left(\mathbf{x}\right)$ if $\Gamma\left(h,\mathbf{x},\bm{\beta}\right)\geq 1$, then,
-$E_{x}\left[f\left(h,x\right)\right]-\frac{1}{n}\sum_{i}f\left(h,x_{i}\right)\leq\sigma\sqrt{\frac{2[\Gamma\left(h,\mathbf{x},\mathbf{\beta}\right)\left(1+1/n\right)+\ln\left(\Gamma\left(h,\mathbf{x},\bm{\beta}\right)\left(n+1\right)/\delta\right)]}{n}}.$
-For the proof, we use the following auxiliary result.
-Lemma B.2.
-(Anthony & Bartlett, 1999, Lemma 15.6) Suppose $\Pr$ is a probability distribution and
-$\left\{E\left(\alpha_{1},\alpha_{2},\delta\right):0<\alpha_{1},\alpha_{2},\delta\leq 1\right\}$
-is a set of events, such that
-(i) For all $0<\alpha\leq 1$ and $0<\delta\leq 1$,
-$\Pr\left\{E\left(\alpha,\alpha,\delta\right)\right\}\leq\delta.$
-(ii) For all $0<\alpha_{1}\leq\alpha\leq\alpha_{2}\leq 1$ and $0<\delta_{1}\leq\delta\leq 1$
-$E\left(\alpha_{1},\alpha_{2},\delta_{1}\right)\subseteq E\left(\alpha,\alpha,\delta\right).$
-Then for $0<a,\delta<1$,
-$\Pr\bigcup_{\alpha\in\left(0,1\right]}E\left(\alpha a,\alpha,\delta\alpha\left(1-a\right)\right)\leq\delta.$
-We put this lemma in a more convenient form.
-Lemma B.3.
-Let $Y$ and $X\geq 0$ be real random variables, $\psi:\mathbb{R}\times\left(0,1\right)\rightarrow\mathbb{R}$ be increasing in the 1st argument and $\forall C>1$, $\delta\in\left(0,1\right)$,
-$\Pr\left\{X\leq C\wedge Y>\psi\left(C,\delta\right)\right\}<\delta.$
-Then, for every $\epsilon>0$
-$\Pr\left\{X\geq 1\wedge Y>\psi\left(X\left(1+\epsilon\right),\frac{\delta\epsilon}{X\left(1+\epsilon\right)}\right)\right\}.$
+where $\left\|.\right\|_{\rm Lip}$ is the Lipschitz-seminorm, $\left\|f\right\|_{Lip}=\inf\left\{s:f\left(h\right)-f\left(g\right)\leq s\left\|h-g\right\|\text{ for all }h,g\in\mathbb{R}^{d}\right\}$.
+If $\nu$ satisfies an LSI with constant $\alpha$ as in (13), then Otto and Villani (2000)
+$W_{p}(\rho,\nu)\leq\frac{2}{\alpha}KL(\rho,\nu).$
+Restatement of Lemma 4.4.
+Lemma C.5.
+Denote
+$\Delta:=\int_{0}^{\beta}\mathbb{E}_{h\sim G_{\gamma}\left(\mathbf{x}\right)}\left[\hat{L}\left(h,\mathbf{x}\right)\right]d\gamma-\Gamma(\nu_{0}^{K-1},\mathbf{x},\beta_{0}^{K})$
+(i) If $\mathbb{E}_{h\sim G_{\beta_{k}}\left(\mathbf{x}\right)}\left[\hat{L}\left(h,\mathbf{x}\right)\right]\leq\mathbb{E}_{h\sim\nu_{k}\left(\mathbf{x}\right)}\left[\hat{L}\left(h,\mathbf{x}\right)\right]$ for all $k$ and $\mathbf{x}$, then $\Delta\leq 0$.
+(ii) If $\ell\left(h,\mathbf{x}\right)$ is bounded in $h$ for all $\mathbf{x}$, $\left\|\ell\left(h,\mathbf{x}\right)\right\|\leq m$ then
+$\Delta\leq m\sum_{k=1}^{K}(\beta_{k}-\beta_{k-1})\sqrt{KL\left(\nu_{k-1}\left(\mathbf{x}\right),G_{\beta_{k-1}}\left(\mathbf{x}\right)\right)/2}.$
+(iii) If instead $\ell\left(h,\mathbf{x}\right)$ is $m$-Lipschitz in $h$ for all $\mathbf{x}$, $\ell\left(h,\mathbf{x}\right)-\ell\left(g,\mathbf{x}\right)\leq m\left\|h-g\right\|$ and $G_{\beta_{k}}\left(\mathbf{x}\right)$ satisfies an LSI with constant $\alpha$ for all $k$ and $\mathbf{x}$, then
+$\Delta\leq\frac{2m}{\alpha}\sum_{k=1}^{K}(\beta_{k}-\beta_{k-1}){KL\left(\nu_{k-1}\left(\mathbf{x}\right),G_{\beta_{k-1}}\left(\mathbf{x}\right)\right)}.$
 Proof.
-This follows from Lemma B.2 using the events
-$E\left(\alpha_{1},\alpha_{2},\delta\right)=\left\{X\leq\alpha_{2}^{-1}\wedge Y>f\left(\alpha_{1}^{-1},\delta\right)\right\}$
-and $a=1/\left(1+\epsilon\right)$. ∎
-Proof of Corollary B.1.
-Take $F=\lambda\left(L\left(h\right)-\hat{L}\left(h,\mathbf{x}\right)\right)$, so $\ln\mathbb{E}_{\mathbf{x}}\mathbb{E}_{g\sim\pi}\left[e^{F\left(g,\mathbf{x}\right)}\right]\leq\lambda^{2}\sigma^{2}/\left(2n\right)$.
-From Theorem 3.2 and the properties of sub-Gaussian variables we get with $\lambda=\sigma^{-1}\sqrt{2n\left(C+\ln\left(1/\delta\right)\right)}$ that
-$\Pr\left\{\Gamma\left(h,\mathbf{x},\bm{\beta}\right)\leq C\wedge L\left(h\right)-\hat{L}\left(h,\mathbf{x}\right)>\sigma\sqrt{\frac{2\left(C+\ln\left(1/\delta\right)\right)}{n}}\right\}$
-$=$ $\Pr\left\{\Gamma\left(h,\mathbf{x},\bm{\beta}\right)\leq C\wedge L\left(h\right)-\hat{L}\left(h,\mathbf{x}\right)>\frac{C+\ln\left(1/\delta\right)}{\lambda}+\frac{\lambda\sigma^{2}}{2n}\right\}$
-$=$ $\Pr\left\{\Gamma\left(h,\mathbf{x},\bm{\beta}\right)\leq C\wedge\lambda\left(L\left(h\right)-\hat{L}\left(h,\mathbf{x}\right)\right)>C+\lambda^{2}\sigma^{2}/\left(2n\right)+\ln\left(1/\delta\right)\right\}\leq\delta.$
-Substitution in Lemma B.3 with $\psi\left(C,\delta\right)=\sigma\sqrt{2\left(C+\ln\left(1/\delta\right)\right)/n}$, $X=\Gamma\left(h,\mathbf{x},\bm{\beta}\right)$, $Y=L\left(h\right)-\hat{L}\left(h,\mathbf{x}\right)$ and $\epsilon=1/n$ gives Corollary B.1.
-∎
-B.2 Proofs for Section 4.2
+By the last assertion of Lemma 3.1 the function $\beta\mapsto\mathbb{E}_{g\sim G_{\beta}(\mathbf{x})}\big[\hat{L}(g,\mathbf{x})\big]$ is non-increasing, so
+$\Delta$ $=$ $\int_{0}^{\beta}\mathbb{E}_{h\sim G_{\gamma}\left(\mathbf{x}\right)}\left[\hat{L}\left(h,\mathbf{x}\right)\right]d\gamma-\Gamma(\nu_{0}^{K-1},\mathbf{x},\beta_{0}^{K})$
+$\leq$ $\sum_{k=1}^{K}\left(\beta_{k}-\beta_{k-1}\right)\mathbb{E}_{h\sim G_{\beta_{k-1}}\left(\mathbf{x}\right)}\left[\hat{L}\left(h,\mathbf{x}\right)\right]-\Gamma(\nu_{0}^{K-1},\mathbf{x},\beta_{0}^{K})$
+$=$ $\sum_{k=1}^{K}\left(\beta_{k}-\beta_{k-1}\right)\left(\mathbb{E}_{h\sim G_{\beta_{k-1}}\left(\mathbf{x}\right)}\left[\hat{L}\left(h,\mathbf{x}\right)\right]-\mathbb{E}_{h\sim\nu_{k-1}\left(\mathbf{x}\right)}\left[\hat{L}\left(h,\mathbf{x}\right)\right]\right).$
+Now if $\mathbb{E}_{h\sim G_{\beta_{k-1}}\left(\mathbf{x}\right)}\left[\hat{L}\left(h,\mathbf{x}\right)\right]\leq\mathbb{E}_{h\sim\nu_{k-1}\left(\mathbf{x}\right)}\left[\hat{L}\left(h,\mathbf{x}\right)\right]$ then (i) is immediate.
+If $\ell$ is bounded by $m$ then $\left\|\hat{L}\left(.,\mathbf{x}\right)\right\|_{\infty}\leq m$ and
+$\left(\mathbb{E}_{h\sim G_{\beta_{k-1}}\left(\mathbf{x}\right)}-\mathbb{E}_{h\sim\nu_{k-1}\left(\mathbf{x}\right)}\right)\left[\hat{L}\left(h,\mathbf{x}\right)\right]\leq m\ d_{TV}\left(G_{\beta_{k-1}}\left(\mathbf{x}\right),\nu_{k-1}\left(\mathbf{x}\right)\right)\leq m\sqrt{KL\left(\nu_{k-1},G_{\beta_{k-1}}\right)/2},$
+by Pinsker’s inequality, which gives (ii). If $\ell$ is Lipschitz in the 1st argument with constant $m$, then so is $\hat{L}$, and
+$\left(\mathbb{E}_{h\sim G_{\beta_{k-1}}\left(\mathbf{x}\right)}-\mathbb{E}_{h\sim\nu_{k-1}\left(\mathbf{x}\right)}\right)\left[\hat{L}\left(h,\mathbf{x}\right)\right]\leq m\ W_{2}\left(G_{\beta_{k-1}}\left(\mathbf{x}\right),\nu_{k-1}\left(\mathbf{x}\right)\right)\leq 2mKL\left(\nu_{k-1},G_{\beta_{k-1}}\right)/\alpha,$
+which gives (ii). ∎
 Restatement of Theorem 4.5:
-Theorem B.4.
-Suppose that $\mathcal{H}=\mathbb{R}^{d}$ and that there are numbers $m,M<\infty$ such that for every $\mathbf{x}$ in $\mathcal{X}^{n}$ and $h\in\mathcal{H}$ we have $\left|\ell\left(h,\mathbf{x}\right)\right|\leq m$ and $\left|F\left(h,\mathbf{x}\right)\right|\leq M$.
-Let $0=\beta_{0}<\beta_{1}<\cdots<\beta_{K}=\beta$ and $\nu_{\beta_{k}}\left(\mathbf{x}\right)\in\mathcal{P}\left(\mathcal{H}\right)$ be such that $d_{TV}\left(\nu_{\beta_{k}}\left(\mathbf{x}\right),G_{\beta_{k}}\left(\mathbf{x}\right)\right)=\epsilon_{\beta_{k}}$.
-Then,
-(i) with probability at least $1-\delta$ as $\mathbf{x}\sim\mu^{n}$ and $h\sim\nu_{\beta}\left(\mathbf{x}\right)$
-$F\left(h,\mathbf{x}\right)$ $\leq$ $\Gamma_{\nu}\left(h,\mathbf{x},\mathbf{\beta}\right)+\ln\mathbb{E}_{\mathbf{x}}\mathbb{E}_{h\sim\pi}\left[e^{F\left(h,\mathbf{x}\right)}\right]+\ln\frac{1}{\delta}$
-$+\ln\left(2e^{M+\beta m}\epsilon_{\beta}\right)+\sum_{k=1}^{K}\left(\beta_{k}-\beta_{k-1}\right)m\epsilon_{\beta_{k-1}}.$
+Theorem C.6.
+Let $F:\mathcal{H}\times\mathcal{X}^{n}\rightarrow\mathbb{R}$ be some measurable function and $\beta_{0}^{K}$ and $\nu_{0}^{K-1}$ as in Definition 4.3.
+Let $\nu(\mathbf{x})$ be any data-dependent distribution on $\mathcal{H}$.
+Let $\Delta$ be bounded as in Lemma 4.4, depending on which of the conditions is fulfilled by $\ell$.
+Then
+(i) with probability at least $1-\delta$ as $\mathbf{x}\sim\mu^{n}$ and $h\sim\nu(\mathbf{x})$
+$F\left(h,\mathbf{x}\right)\hskip-4.26773pt$ $\leq\hskip-5.69046pt$ $-\beta\hat{L}\left(h,\mathbf{x}\right)+\Gamma(\nu_{0}^{K-1},\mathbf{x},\beta_{0}^{K})+\ln\mathbb{E}_{\mathbf{x}}\mathbb{E}_{h\sim\pi}\left[e^{F\left(h,\mathbf{x}\right)}\right]+\ln\frac{1}{\delta}$
+$+R_{\infty}\left(\nu\left(\mathbf{x}\right),G_{\beta}\left(\mathbf{x}\right)\right)+\Delta.$
+If $F$ and $\ell$ are bounded, then $R_{\infty}\left(\nu\left(\mathbf{x}\right),G_{\beta}\left(\mathbf{x}\right)\right)$ can be replaced by
+$\max\left\{0,\beta\left\|\ell\right\|_{\infty}+\left\|F\right\|_{\infty}+\ln\sqrt{2KL\left(\nu\left(\mathbf{x}\right),G_{\beta}\left(\mathbf{x}\right)\right)}\right\}.$
 (ii) with probability at least $1-\delta$ as $\mathbf{x}\sim\mu^{n}$
-$\mathbb{E}_{h\sim\nu_{\beta}\left(\mathbf{x}\right)}\left[F\left(h,\mathbf{x}\right)\right]$ $\leq$ $\mathbb{E}_{h\sim\nu_{\beta}\left(\mathbf{x}\right)}\left[\Gamma_{\nu}\left(h,\mathbf{x},\bm{\beta}\right)\right]+\ln\mathbb{E}_{\mathbf{x}}\mathbb{E}_{h\sim\pi}\left[e^{F\left(h,\mathbf{x}\right)}\right]+\ln\frac{1}{\delta}$
-$+\left(M+\beta m\right)\epsilon_{\beta}+\sum_{k=1}^{K}\left(\beta_{k}-\beta_{k-1}\right)m\epsilon_{\beta_{k-1}}.$
+$\mathbb{E}_{\nu\left(\mathbf{x}\right)}\!\left[F(h,\mathbf{x})\right]$ $\hskip-15.6491pt\leq\hskip-14.22636pt$ $\hskip-2.84544pt-\beta\mathbb{E}_{\nu\left(\mathbf{x}\right)}\big[\hat{L}\left(h,\mathbf{x}\right)\big]+\Gamma(\nu_{0}^{K-1},\mathbf{x},\beta_{0}^{K})+\ln\mathbb{E}_{\mathbf{x}}\mathbb{E}_{h\sim\pi}\left[e^{F\left(h,\mathbf{x}\right)}\right]+\ln\frac{1}{\delta}$
+$+KL\left(\nu\left(\mathbf{x}\right),G_{\beta}\left(\mathbf{x}\right)\right)+\Delta.$
 Proof.
-By the bound on $\ell$ we have
-$\Gamma\left(h,\mathbf{x},\bm{\beta}\right)\leq\Gamma_{\nu}\left(h,\mathbf{x},\bm{\beta}\right)+\sum_{k=1}^{K}\left(\beta_{k}-\beta_{k-1}\right)m\epsilon_{\beta_{k-1}}\text{.}$ (8)
-(i) From Markov’s inequality we have (in analogy to the proof of Theorem 3.2) with probability at least $1-\delta$ as $x\sim\mu^{n}$ and $h\sim\nu_{\beta}\left(\mathbf{x}\right)$, that
+Proof of (ii). By equation (6) and Lemma 4.4 we have
+$KL\left(\nu\left(\mathbf{x}\right),\pi\right)$ $\leq$ $KL\left(\nu\left(\mathbf{x}\right),G_{\beta}\left(\mathbf{x}\right)\right)-\beta\mathbb{E}_{h\sim\nu\left(\mathbf{x}\right)}\left[\hat{L}\left(h,\mathbf{x}\right)\right]+\int_{0}^{\beta}\mathbb{E}_{h\sim G_{\gamma}\left(\mathbf{x}\right)}d\gamma$
+$\leq$ $KL\left(\nu\left(\mathbf{x}\right),G_{\beta}\left(\mathbf{x}\right)\right)-\beta\mathbb{E}_{h\sim\nu\left(\mathbf{x}\right)}\left[\hat{L}\left(h,\mathbf{x}\right)\right]+\Gamma(\nu_{0}^{K-1},\mathbf{x},\beta_{0}^{K})+\Delta.$
+Substitution in (ii) of the PAC-Bayesian theorem proves (ii).
+Proof of (i) without amendment. Similarly, we have from Lemma 4.4
+$\ln\frac{d\nu\left(\mathbf{x}\right)}{d\pi}\left(h\right)$ $=$ $\ln\frac{d\nu\left(\mathbf{x}\right)}{dG_{\beta}\left(\mathbf{x}\right)}\left(h\right)+\ln\frac{dG_{\beta}\left(\mathbf{x}\right)}{d\pi}\left(h\right)$
+$\leq$ $R_{\infty}\left(\nu\left(\mathbf{x}\right),G_{\beta}\left(\mathbf{x}\right)\right)-\beta\hat{L}\left(h,\mathbf{x}\right)+\int_{0}^{\beta}\mathbb{E}_{h\sim G_{\gamma}\left(\mathbf{x}\right)}d\gamma$
+$=$ $R_{\infty}\left(\nu\left(\mathbf{x}\right),G_{\beta}\left(\mathbf{x}\right)\right)-\beta\hat{L}\left(h,\mathbf{x}\right)+\Gamma(\nu_{0}^{K-1},\mathbf{x},\beta_{0}^{K})+\Delta.$
+Then substitute in the PAC-Bayesian theorem (i).
+For the proof of the amendment to (i) we have to return to the proof of the PAC-Bayesian theorem.
+From Markov’s inequality we have (in analogy to the proof of Theorem 2.1) with probability at least $1-\delta$ as $x\sim\mu^{n}$ and $h\sim\nu\left(\mathbf{x}\right)$, that
 $F\left(h,\mathbf{x}\right)+\beta\hat{L}\left(h,\mathbf{x}\right)+\ln Z_{\beta}\left(\mathbf{x}\right)$
-$\leq$ $\ln\mathbb{E}_{\mathbf{x}}\mathbb{E}_{h\sim\nu_{\beta}\left(\mathbf{x}\right)}\left[e^{F\left(h,\mathbf{x}\right)+\beta\hat{L}\left(h,\mathbf{x}\right)+\ln Z_{\beta}\left(\mathbf{x}\right)}\right]+\ln\left(1/\delta\right)$
-$\leq$ $\ln\left(\mathbb{E}_{\mathbf{x}}\mathbb{E}_{h\sim G_{\beta}\left(\mathbf{x}\right)}\left[e^{F\left(h,\mathbf{x}\right)+\beta\hat{L}\left(h,\mathbf{x}\right)+\ln Z_{\beta}\left(\mathbf{x}\right)}\right]+e^{M+\beta m}\epsilon_{\beta}\right)+\ln\left(1/\delta\right)$
-$=$ $\ln\left(\mathbb{E}_{\mathbf{x}}\mathbb{E}_{h\sim\pi}\left[e^{F\left(h,\mathbf{x}\right)}\right]+e^{M+\beta m}\epsilon_{\beta}\right)+\ln\left(1/\delta\right)$
-$\leq$ $\ln\mathbb{E}_{\mathbf{x}}\mathbb{E}_{h\sim\pi}\left[e^{F\left(h,\mathbf{x}\right)}\right]+\ln\left(2e^{M+\beta m}\epsilon_{\beta}\right)+\ln\left(1/\delta\right).$
-In the second inequality we used $\ln Z_{\beta}\left(\mathbf{x}\right)\leq 0$ and in the last line we used for $a,b\geq 1$ that $\ln\left(a+b\right)\leq\ln\max\left\{a,b\right\}+\ln 2\leq\ln a+\ln b+\ln 2=\ln a+\ln 2b$.
-Subtract $\beta\hat{L}\left(h,\mathbf{x}\right)+\ln Z_{\beta}\left(\mathbf{x}\right)$, use (4) and (8).
-(ii) Again with Markov’s inequality, with probability at least $1-\delta$ as $x\sim\mu^{n}$
-$\mathbb{E}_{h\sim\nu_{\beta}\left(\mathbf{x}\right)}\left[F\left(h,\mathbf{x}\right)+\beta\hat{L}\left(h,\mathbf{x}\right)+\ln Z_{\beta}\left(\mathbf{x}\right)\right]$
-$\leq$ $\ln\mathbb{E}_{\mathbf{x}\sim\mu^{n}}\left[e^{\mathbb{E}_{h\sim\nu_{\beta}\left(\mathbf{x}\right)}\left[F\left(h,\mathbf{x}\right)+\beta\hat{L}\left(h,\mathbf{x}\right)+\ln Z_{\beta}\left(\mathbf{x}\right)\right]}\right]$
-$\leq$ $\ln\mathbb{E}_{\mathbf{x}\sim\mu^{n}}\left[e^{\mathbb{E}_{h\sim G_{\beta}\left(\mathbf{x}\right)}\left[F\left(h,\mathbf{x}\right)+\beta\hat{L}\left(h,\mathbf{x}\right)+\ln Z_{\beta}\left(\mathbf{x}\right)\right]+\left(M+\beta m\right)\epsilon_{\beta}}\right].$
-In the second inequality we used the fact that $\ln Z_{\beta}\left(\mathbf{x}\right)\leq 0$ and the bounds on $F$ and $\ell$. Then, Jensen’s inequality bounds the last line as
-$\ln\mathbb{E}_{\mathbf{x}\sim\mu^{n}}\mathbb{E}_{h\sim G_{\beta}\left(\mathbf{x}\right)}\left[e^{F\left(h,\mathbf{x}\right)+\beta\hat{L}\left(h,\mathbf{x}\right)+\ln Z_{\beta}\left(\mathbf{x}\right)}\right]e^{\left(M+\beta m\right)\epsilon_{\beta}}$
-$=$ $\ln\mathbb{E}_{\mathbf{x}\sim\mu^{n}}\mathbb{E}_{h\sim\pi}\left[e^{F\left(h,\mathbf{x}\right)}\right]+\left(M+\beta m\right)\epsilon_{\beta}.$
-Again subtract $\beta\hat{L}\left(h,\mathbf{x}\right)+\ln Z_{\beta}\left(\mathbf{x}\right)$, use (4) and (8). ∎
-B.3 Miscellaneous lemmata
-Lemma B.5.
+$\leq$ $\ln\mathbb{E}_{\mathbf{x}}\mathbb{E}_{h\sim\nu\left(\mathbf{x}\right)}\left[e^{F\left(h,\mathbf{x}\right)+\beta\hat{L}\left(h,\mathbf{x}\right)+\ln Z_{\beta}\left(\mathbf{x}\right)}\right]+\ln\left(1/\delta\right)$
+$\leq$ $\ln\left(\mathbb{E}_{\mathbf{x}}\mathbb{E}_{h\sim G_{\beta}\left(\mathbf{x}\right)}\left[e^{F\left(h,\mathbf{x}\right)+\beta\hat{L}\left(h,\mathbf{x}\right)+\ln Z_{\beta}\left(\mathbf{x}\right)}\right]+e^{\beta\left\|\ell\right\|_{\infty}+\left\|F\right\|_{\infty}}d_{TV}\left(\nu\left(\mathbf{x}\right),G_{\beta}\left(\mathbf{x}\right)\right)\right)+\ln\left(1/\delta\right)$
+$=$ $\ln\left(\max\left\{\mathbb{E}_{\mathbf{x}}\mathbb{E}_{h\sim\pi}\left[e^{F\left(h,\mathbf{x}\right)}\right],1\right\}+\max\left\{e^{\beta\left\|\ell\right\|_{\infty}+\left\|F\right\|_{\infty}}d_{TV}\left(\nu\left(\mathbf{x}\right),G_{\beta}\left(\mathbf{x}\right)\right),1\right\}\right)+\ln\left(1/\delta\right)$
+$\leq$ $\max\left\{\ln\mathbb{E}_{\mathbf{x}}\mathbb{E}_{h\sim\pi}\left[e^{F\left(h,\mathbf{x}\right)}\right],0\right\}+\max\left\{\beta\left\|\ell\right\|_{\infty}+\left\|F\right\|_{\infty}+\ln\left(2d_{TV}\left(\nu\left(\mathbf{x}\right),G_{\beta}\left(\mathbf{x}\right)\right)\right),0\right\}$
+$\ \ \ +\ln\left(1/\delta\right).$
+In the second inequality we used $\ln Z_{\beta}\left(\mathbf{x}\right)\leq 0$ and the property of the total variation metric.
+In the next inequality we used Lemma 3.1, and in the last line we used for $a,b\geq 1$ that $\ln\left(a+b\right)\leq\ln\max\left\{a,b\right\}+\ln 2\leq\ln a+\ln b+\ln 2$.
+Subtract $\beta\hat{L}\left(h,\mathbf{x}\right)+\ln Z_{\beta}\left(\mathbf{x}\right)$, use Lemma 3.1, Lemma 4.4 and Pinsker’s inequality to bound the total variation distance.
+C.5 Miscellaneous Lemmata
+Lemma C.7.
 For $0<\beta<\infty$
-$\max\left\{KL\left(G_{\beta},G_{2\beta}\right),KL\left(G_{2\beta},G_{2}\right)\right\}\leq\beta\left(\mathbb{E}_{h\sim G_{\beta}}\left[\hat{L}\left(h\right)\right]-\mathbb{E}_{h\sim G_{2\beta}}\left[\hat{L}\left(h\right)\right]\right)$
+$\max\left\{KL\left(G_{\beta},G_{2\beta}\right),KL\left(G_{2\beta},G_{2\beta}\right)\right\}\leq\beta\left(\mathbb{E}_{h\sim G_{\beta}}\left[\hat{L}\left(h\right)\right]-\mathbb{E}_{h\sim G_{2\beta}}\left[\hat{L}\left(h\right)\right]\right)$
 Proof.
-Using Lemma 3.1
 $KL\left(G_{\beta},G_{2\beta}\right)$ $=$ $\mathbb{E}_{h\sim G_{\beta}}\left[-\beta\hat{L}\left(h\right)-\ln Z_{\beta}+2\beta\hat{L}\left(h\right)+\ln Z_{2\beta}\right]$
 $=$ $\mathbb{E}_{h\sim G_{\beta}}\left[\beta\hat{L}\left(h\right)\right]-\int_{\beta}^{2\beta}\mathbb{E}_{h\sim G_{\gamma}}\left[\hat{L}\left(h\right)\right]d\gamma$
-$\leq$ $\beta\left(\mathbb{E}_{h\sim G_{\beta}}\left[\hat{L}\left(h\right)\right]-\mathbb{E}_{h\sim G_{2\beta}}\left[\hat{L}\left(h\right)\right]\right).$
-Similarly
-$KL\left(G_{2\beta},G_{2}\right)$ $=$ $\mathbb{E}_{h\sim G_{2\beta}}\left[-2\beta\hat{L}\left(h\right)-\ln Z_{2\beta}+\beta\hat{L}\left(h\right)+\ln Z_{\beta}\right]$
+$\leq$ $\beta\left(\mathbb{E}_{h\sim G_{\beta}}\left[\hat{L}\left(h\right)\right]-\mathbb{E}_{h\sim G_{2\beta}}\left[\hat{L}\left(h\right)\right]\right)$
+Similarly,
+$KL\left(G_{2\beta},G_{\beta}\right)$ $=$ $\mathbb{E}_{h\sim G_{2\beta}}\left[-2\beta\hat{L}\left(h\right)-\ln Z_{2\beta}+\beta\hat{L}\left(h\right)+\ln Z_{\beta}\right]$
 $=$ $-\mathbb{E}_{h\sim G_{\beta}}\left[\beta\hat{L}\left(h\right)\right]+\int_{\beta}^{2\beta}\mathbb{E}_{h\sim G_{\gamma}}\left[\hat{L}\left(h\right)\right]d\gamma$
-$\leq$ $\beta\left(\mathbb{E}_{h\sim G_{\beta}}\left[\hat{L}\left(h\right)\right]-\mathbb{E}_{h\sim G_{2\beta}}\left[\hat{L}\left(h\right)\right]\right).$
-Appendix C Experimental details and additional results
-C.1 Experimental details
+$\leq$ $\beta\left(\mathbb{E}_{h\sim G_{\beta}}\left[\hat{L}\left(h\right)\right]-\mathbb{E}_{h\sim G_{2\beta}}\left[\hat{L}\left(h\right)\right]\right)$
+C.6 The Calibration Factor
+Let $\bar{A}=-\beta\mathbb{E}_{h\sim\nu_{\beta}}\left[\hat{L}\left(h,\mathbf{\bar{x}}\right)\right]+\Gamma\left(\nu_{0}^{K-1},\mathbf{\bar{x}},\beta_{0}^{K}\right)$ be the area estimate obtained for the random labels (corrected by $\beta\mathbb{E}_{h\sim\nu_{\beta}}\left[\hat{L}\left(h,\mathbf{\bar{x}}\right)\right]$, which should play little role for large $\beta>n$).
+Then
+$r\left(\mathbf{x}\right)=\min\left\{r:\forall k\in\left[K\right],\text{ }\kappa^{-1}\left(\mathbb{E}_{h\sim\nu_{\beta_{k}}\left(\mathbf{x}\right)}\left[\hat{L}_{01}\left(h,\mathbf{\bar{x}}\right)\right],\frac{1}{n}\left(r\bar{A}+\ln\frac{2\sqrt{n}}{\delta}\right)\right)\geq\frac{1}{2}\right\},$
+where $\mathbf{\bar{x}}$ is the training set $\mathbf{x}$ with random labels and $\hat{L}_{01}$ the empirical 01-error.
+The calibration value $r$ is thus the smallest factor of $\bar{A}$, for which we obtain a correct upper bound on the 01-error with random labels for all the $\beta_{k}$.
+We emphasize that the calibration procedure depends only on the training data.
+Appendix D Experimental Details and Additional Results
+D.1 Experimental Details
 All the codes to reproduce the results are provided through this https://anonymous.4open.science/r/Gibbs-Generalization-45F1.
-For all the experiments we use the isotropic Gaussian prior with $\mu=0$ and $\sigma=5$.
-This induces an L2-regularization term in the energy function that is stated in the proof of Corollary 4.2.
+For all the experiments we use an isotropic Gaussian prior with $\mu=0$, for bounded loss with $\sigma=5$ and for unbounded loss with $\sigma=0.1$.
+This induces an L2-regularization term in the energy function that is stated in the proof of Corollary C.2.
 The confidence parameter $\delta$ appearing in our bounds is set to $0.01$ for all experiments
@@
 For both datasets, MNIST and CIFAR-10, we use neural networks with ReLU activation functions.
-C.1.1 Network architecture
+D.1.1 Networks Architecture
 The fully connected networks consist of one, two, three hidden layers, each containing a constant number of units.
 Besides that we are using LeNet-5 architecture for MNIST and VGG16 architecture for CIFAR-10 to achieve low test error.
-For loss function $\ell$, we are using either bounded binary cross-entropy (BBCE) as described in Appendix D of Dziugaite & Roy (2018) or the savage loss (Masnadi-Shirazi & Vasconcelos, 2008).
+For loss function $\ell$, we are mostly using bounded loss function such as bounded binary cross-entropy (BBCE) as described in Appendix D of Dziugaite and Roy (2018) or the Savage loss (Masnadi-Shirazi and Vasconcelos, 2008).
+As unbounded loss function we tried binary cross-entropy (BCE) (Section D.2.6), but with a smaller value of $\sigma$, so as to avoid excessive training errors for small values of $\beta$.
 The LeNet-5 network follows a systematic pattern of alternating convolutional and pooling layers, followed by fully connected layers (LeCun et al., 2002).
@@
 Throughout the network, ReLU activation functions replace the original tanh activations, which improves gradient flow and training performance in modern implementations.
-VGG-16 is a widely used deep convolutional neural network architecture known for its simplicity and strong performance in image classification tasks (Simonyan & Zisserman, 2014).
+VGG-16 is a widely used deep convolutional neural network architecture known for its simplicity and strong performance in image classification tasks (Simonyan and Zisserman, 2014).
 The architecture follows a consistent design using only $3\times 3$ convolutional filters and $2\times 2$ max pooling operations throughout the network.
@@
 Each block is followed by a $2\times 2$ max pooling layer for spatial downsampling.
-All convolutional layers employ $3\times 3$ kernels with padding to preserve spatial dimensions, and ReLU activation functions introduce non-linearity.The convolutional feature extractor is followed by a classifier head consisting of three fully connected layers: two hidden layers with 1024 neurons each, using ReLU activation, and a final output layer with 1 neurons for binary classification.
+All convolutional layers employ $3\times 3$ kernels with padding to preserve spatial dimensions, and ReLU activation functions introduce non-linearity.The convolutional feature extractor is followed by a classifier head consisting of three fully connected layers: two hidden layers with 1024 neurons each, using ReLU activation, and a final output layer with 1 neuron for binary classification.
 We also removed dropout to ensure that SGLD minimizes the defined energy function without any additional terms.
@@
 For evaluating our models, we are using all 10,000 test examples for both datasets.
-C.1.2 Minibatches
-When using SGLD, we adopt minibatches of size proportional to $\sqrt{n}$. Thus, for $n=2000$ the mini-batch size is 50, and for $n=8000$ it is 100.
-C.1.3 Moving average filters
+D.1.2 Minibatches
+When using SGLD, we adopt minibatches of size proportional to $\sqrt{n}$. Thus, for $n=2000$ the minibatch size is 50, and for $n=8000$ it is 100.
+D.1.3 Moving Average Filters
 As we explained in Section 5.3, we are using a running mean $\mathbb{M}(x_{1},\cdots,x_{t})$ of $\hat{L}(h_{j},\mathbf{x})$ from $j=1,\cdots,t$ both as a criterion to stop the experiment and an estimation for $\mathbb{E}_{h\sim G_{\beta_{k}}}\left[\hat{L}\left(h,\mathbf{x}\right)\right]$.
@@
 $\mathbb{M}_{t}$ $=\alpha\hat{L}(h_{t},\mathbf{x})+(1-\alpha)\,\mathbb{M}_{t-1},$
-with $\mathbb{M}_{0}=0$ and small $\alpha$.
+with $\mathbb{M}_{0}=1$ and small $\alpha$.
 We use the first (symmetric) form in the experiments with ULA, and the second (standard exponential moving average) form with SGLD for convenience.
@@
 As $\alpha\to 0$ and $t\to\infty$, the quantity $\mathbb{M}_{t}$ converges to the ergodic mean.
-C.2 Experimental results
-C.2.1 Illustration of Bound Computation
-In this section, we demonstrate again the figure in the main body in more details.
-The figure 2 illustrates how our bounds are computed.
-The sequence of mean training losses in $\ell$ is used to compute for each $\beta$ the functional $\Gamma$ and the "KL-Bound", which corresponds to the right hand side of the inequalities in Corollary 3.3.
+D.2 Experimental Results
+D.2.1 Illustration of Bound Computation
+In this section, we discuss figure 4 in the main body in more details.
+Figure 5 illustrates how our bounds are computed.
+The sequence of mean training losses in $\ell$ is used to compute for each $\beta$ the functional $\Gamma$ and the "KL-Bound", which corresponds to the right-hand side of the inequalities in Corollary 11.
 Our bound on the test loss is then computed by applying the function $\kappa^{-1}$ to the empirical 0-1 error and to this kl-bound.
-The graph of "KL(Train, Test)" corresponds to the left hand side in Corollary 3.3.
+The graph of "KL(Train, Test)" corresponds to the left-hand side in Corollary 11.
 It is remarkable that the close fit of the upper bound on the random labels is achieved by the adjustment of a single calibration parameter.
 Figure 5: A more detailed version of Figure 4 to illustrate how the bounds are computed.
-C.2.2 Single-draws
+D.2.2 Single-draws
 For the setting described in Section 5.5, we also present the bounds for the single-draw case in Figure 6.
@@
 However, as visible in the plots, the results exhibit fluctuations and irregularities caused by stochastic effects, which make them less reliable.
-Figure 3: SGLD on MNIST and CIFAR-10 with 2000 training examples using BBCE loss function.
+Figure 6: SGLD on MNIST and CIFAR-10 with 8000 training examples using BBCE loss function.
 The first row corresponds to MNIST and the second row to CIFAR-10.
@@
 Both random and true labels are trained with exactly the same algorithm and parameters on a fully connected ReLU network with two hidden layers of 1000 (respectively 1500) units.
-The calibration factor for MNIST is 0.19, for CIFAR-10 0.22.
+The calibration factor for MNIST is 0.77, for CIFAR-10 0.89.
 Train error, test error and our bound for a single-draw of the 0-1 loss are plotted against $\beta$.
-C.2.3 Different Architectures
+D.2.3 Different Architectures
 In this section, we evaluate the performance of different models and architectures on both MNIST and CIFAR-10, demonstrating that our bound can be used to guide model selection.
@@
 Furthermore, we employ the LeNet-5 architecture for MNIST and VGG-16 for CIFAR-10 to achieve high test accuracy.
-Detailed descriptions of these architectures are provided in Section C.1.1.
+Detailed descriptions of these architectures are provided in Section D.1.1.
 Figure 7 demonstrates the robustness of our bound across different models.
@@
 The bounds at both low and high temperatures reliably reflect test error performance at low temperature.
-C.2.4 ULA
+D.2.4 ULA
 We have also conducted experiments using ULA for both datasets.
@@
 Train error, test error and our bound for the Gibbs posterior average of the 0-1 loss are plotted against $\beta$.
-C.2.5 Savage Loss function
+D.2.5 Savage Loss Function
 We additionally performed experiments using the Savage loss to verify the robustness of our results across different loss functions.
@@
 Train error, test error and our bound for the Gibbs posterior average of the 0-1 loss are plotted against $\beta$.
-C.2.6 Uncalibrated Bounds
-We show uncalibrated bounds for the MNIST dataset with the BBCE and Savage loss functions under the experimental conditions of Sections C.2.4 and C.2.5.
-The bounds are somewhat looser than the calibrated ones, but still far from trivial.
-As in all other scenarios the test errors are upper bounded correctly.
-Figure 7: ULA on MNIST with 2000 training examples using BBCE and savage loss functions.
-The first row corresponds to MNIST with BBCE and the second row to savage.
+D.2.6 Unbounded Loss Function
+In this section, we use the binary cross-entropy loss to compute the $\Gamma$ functional.
+Since binary cross-entropy is unbounded, the loss can become very large at high temperatures.
+To avoid this issue, we set the standard deviation of the Gaussian prior to 0.1 in this section.
+The following plot shows the results under the same setup as Section D.2.5, except that we use binary cross-entropy instead of the Savage loss.
+Figure 10: ULA on MNIST with 2000 training examples using binary cross entropy loss function.
 Random labels are shown on the left, correct labels on the right.
 Both random and true labels are trained with the same algorithm and parameters on a fully connected ReLU network with one hidden layers of 500 units.
-Train error, test error and our bound for the Gibbs posterior average of the 0-1 loss are plotted against $\beta$ without any calibration.
-C.2.7 Real-world use cases
+The calibration factor is 0.34.
+Train error, test error and our bound for the Gibbs posterior average of the 0-1 loss are plotted against $\beta$.
+D.2.7 Real-World Use Cases
 We further evaluated Stochastic Gradient Descent (SGD) to examine the practical relevance of our bounds in real-world interpolation regimes.
@@
 2HL (W=1000) 3HL (W=500) LeNet-5
-Test Error with SGD 0.0364 0.0363 0.0308
-Test Error at $\beta=64k$ 0.0498 0.0549 0.0317
-Test Bound at $\beta=64k$ 0.0860 0.1314 0.0375
+Test Error, SGD 0.0364 0.0363 0.0308
+Test Error, SGLD ($\beta=64k$) 0.0498 0.0549 0.0317
+Test Bound, SGLD ($\beta=64k$) 0.0860 0.1314 0.0375
 (a) MNIST, 8k training examples (true labels).
 2HL (W=1500) 3HL (W=1000) VGG-16
-Test Error with SGD 0.1423 0.1415 0.0933
-Test Error at $\beta=64k$ 0.1719 0.1782 0.0903
-Test Bound at $\beta=64k$ 0.2266 0.2807 0.2030
+Test Error, SGD 0.1423 0.1415 0.0933
+Test Error, SGLD ($\beta=64k$) 0.1719 0.1782 0.0903
+Test Bound, SGLD ($\beta=64k$) 0.2266 0.2807 0.2030
 (b) CIFAR-10, 8k training examples (true labels).